### v1

# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

Abstract
The quantum expectation value and the stationary noise spectral density for a Fabri-Pérot gravitational-wave detector with a DC readout scheme are discussed in detail only through the quantum electrodynamics of lasers and the Heisenberg equations of mirrors’ motion. We demonstrate that the initial conditions of the mirrors’ motion concentrate around the fundamental frequency of the pendulum and are not related to the frequency range of our interest. Although in the ideal case, there is a consensus that the shot noise for the laser in the high-frequency range in the signal-referred noise spectral density decreases if the injected laser power is increased, our obtained noise spectral density shows that the shot noise does not decrease even if the injected laser power is increased. This is due to the leakage of the classical radiation pressure forces from the classical carrier field to the output port, and the classical carrier field is used as the reference in the DC readout scheme. Since the classical radiation pressure forces affect the mirror motion as a classical constant force, this classical constant force changes the equilibrium point of the pendulum of mirrors’ motion. To recover the ideal case, we must consider adjusting the interferometer’s tuning point to place the mirrors in equilibrium points. We investigate the case where the equilibrium tuning is incomplete and show that the behavior of the above shot noise is due to this incompleteness. We also discuss the maximal deviation of the mirror displacements from the equilibrium point in the incomplete tuning to recover the near ideal case.
gravitational-wave detector; quantum noise; DC readout scheme
I Introduction
A decade ago, the first detection of the gravitational-wave signal GW150914 from a black hole-black hole binary was achieved. This event marked the beginning of gravitational-wave astronomy and multi-messenger astronomy including gravitational-wave detection [1, 2]. We are now at the stage where we can directly measure gravitational waves, and we can conduct scientific research through these gravitational-wave events. Many events, mainly from black hole-black hole binaries, have already been detected. We can also expect that one of the future directions of gravitational-wave astronomy is the development as a “precise science.” This will involve detailed studies of source science from the astrophysical point of view, the tests of general relativity, and the developments of the global network of gravitational-wave detectors [2, 3, 4, 5]. In addition to the current network of ground-based detectors, as future ground-based gravitational-wave detectors, the projects of Einstein Telescope [6] and Cosmic Explorer [7] are in progress, aiming to achieve more sensitive detections.
To advance the field of gravitational-wave detections as a more precise science, it is essential to improve detector sensitivity. We note that theoretical developments in detector science are also crucial for enhancing our basic understanding of these gravitational-wave detectors. The current gravitational-wave detectors are limited by the shot noise, which is one of the quantum noises in the interferometer. In interferometric gravitational detectors, it is considered that there are two kinds of quantum noises. One is the shot noise of the laser, which constrains the sensitivity in the high-frequency range of the gravitational-wave detectors. On the other hand, the quantum radiation pressure forces are considered as a constraint on the sensitivity in the low-frequency region. In the common understanding in the gravitational-wave detection community, if the injected power is increased, the shot noise is decreased, while the quantum radiation pressure forces are increased [8]. The envelop of this trade-off relation in the signal-referred noise spectral density is regarded as the “standard quantum limit.” It is also a common understanding in the gravitational-wave detection community that this standard quantum limit is also estimated from the Heisenberg uncertainty relation of the noncommutativity of the quantum positions and the momenta of test masses [9].
Furthermore, in the current research on the gravitational-wave detection, there are some reports which state that the current LIGO gravitational-wave detectors already violate the above “standard quantum limit” [10, 11]. Even from a theoretical perspective, this presents an opportune moment to revisit and refine the theoretical arguments, aiming for greater accuracy than those presented in previous works in Refs. [8, 9].
On the other hand, a mathematically rigorous quantum measurement theory in quantum mechanics has also been developed (Ref. [12] and the references therein), which introduces new error-disturbance relations, distinct from the Heisenberg uncertainty relation. These new error-disturbance relations have already been confirmed through experiments [14, 15, 16]. One of the motivations of this development was the detection of gravitational-waves [13]. However, the actual application of this mathematical theory to the gravitational-wave detectors requires its extension to the quantum field theories, because the quantum noise in gravitational-wave detectors is analyzed through the quantum field theories of lasers [8]. Moreover, in the quantum measurement theory, it is essential to specify the final measured operator in the quantum measurement process due to the von Neumann chain problem [17]. In interferometric gravitational-wave detectors, the directly measured quantum operator is identified within the “readout scheme” of the detectors. While current gravitational-wave detectors utilize a feedback control system, and the final measured data of gravitational-wave detectors consist of the electric currents from this feedback control, the readout scheme in gravitational-wave detectors is the optical system that determines the finally measured quantum operator in the optical fields detected at the photodetectors, which receive signals from the main interferometer. Therefore, research into this readout scheme is crucial for advancing mathematical quantum measurement theory and its application to gravitational-wave detections.
Due to the issue of the von Neumann chain, this paper adopts the perspective that the quantum properties in the laser interferometers are preserved until the photons are detected. At the photo-detectors, the quantum information of photons is transformed into an electric current. The processes that occur within the photodetectors are complex and vary depending on the type of device used. As a result, we speculate that quantum decoherence takes place during this photodetection process, and the resulting electric current, generated from the statistical results of the detection, can be regarded as classical currents. If this perspective is incorrect and the quantum nature is retained even in the feedback electric current, we would need to explore the concept of a quantum feedback control system [18]. This consideration is beyond the current scope of this paper.
Current gravitational-wave detectors utilize a “DC readout scheme,” which is explained in Sec. III of this paper. Our first step is to investigate the DC-readout scheme from a quantum theoretical perspective. We previously discussed the mathematically rigorous quantum description of the balanced homodyne detection as a readout scheme in gravitational-wave detectors [19], since the balanced homodyne detection is planned for installation in future gravitational-wave detectors. In this paper, we will consider the DC readout scheme as a readout scheme for gravitational-wave detectors. Our investigation in this paper is a natural extension of the quantum theoretical arguments of a balanced homodyne detection [19]. The key aspect of our argument lies in the specification of the quantum operator that is finally observed. We assume that the finally observed quantum operator in the photodetectors is Glauber’s photon number [20], as discussed in Sec. III. It is often stated that the finally observed quantum operator is the power operator of the laser. However, even when we utilize the laser’s power operator, the conclusions we arrive at for the DC readout scheme are consistent with those related to Glauber’s photon number. Generally speaking, if we employ a different quantum operator, such as the mode-by-mode number operator, we may draw different conclusions [21]. This variation depends on the photodetection device, as mentioned above. Therefore, the specification of our finally measured quantum operator is essential to our arguments.
After discussing the general arguments related to the DC readout scheme, we will examine the input-output relation for a Fabri-Pérot gravitational-wave detector. A specific input-output relation is essential for a detailed analysis of the DC readout scheme. We aim to derive this input-output relation of a Fabri-Pérot gravitational-wave detector, starting without specifying the small motions of the mirrors. These small motions of mirrors encapsulate information about external forces, which include gravitational-wave signals and the radiation pressure forces from the laser. In this paper, we describe the small motions of mirrors using the Heisenberg equation in quantum mechanics for a forced harmonic oscillator. While previous literature has often considered the mirrors’ small motions as free except for the gravitational-wave signal and the radiation pressure forces, we have introduced the fundamental frequency $\omega_{p}$ of the vibration isolation system of the pendulum supporting the mirrors.
After solving the Heisenberg equations for the motion of the mirrors, we find that the initial conditions for the forced harmonic oscillator concentrate at the frequency $\omega_{p}$, which is outside the frequency range of our interest. This indicates that the mirrors’ initial conditions, which have information about the noncommutativity of the position and the momentum in quantum mechanics, do not relate to the frequency range relevant to gravitational-wave detectors.
We find that the classical radiation pressure force also affects the mirror, which is finite due to the introduced fundamental frequency $\omega_{p}$ of pendulums. This classical radiation pressure manifests as a classical carrier field that leaks to the output port of the interferometer. This leakage influences the noise estimation in the DC readout scheme. We discuss the effects of this leakage of classical radiation pressure forces in the stationary noise spectral densities. Notably, when we consider the effects of this leakage, there is a situation where the shot noise contribution to the signal-referred noise spectral density in the high-frequency range does not decrease, even when the incident laser power is increased. To avoid the effects of the leakage of the classical radiation pressure forces, we have to adjust the tuning point of the Fabry-Pérot interferometer very carefully. We also estimate the maximum deviation from the ideal mirror position which realizes the near ideal case.
Figure 1: The setup for the Fabry-Pérot interferometer includes notations for photoelectric quadratures. We define the classical distances $L$, $l_{x}$, and $l_{y}$, along with the small quantum displacements represented by $\hat{X}_{XEM}$, $\hat{X}_{XITM}$, $\hat{X}_{YEM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ for the end mirrors (EMs) and the intermediate mirrors (ITMs). Additionally, the quadratures for the laser are denoted by $\hat{d}$ and $\hat{a}$ as the input quadratures to the interferometer, while $\hat{b}$ represents the output quadrature from the interferometer. The quadratures $\hat{c}_{x}$ and $\hat{c}_{y}$ are for the laser, which is separated by the beam splitter (BS), and $\hat{c}_{x}^{\prime}$ and $\hat{c}_{y}^{\prime}$ denote the quadratures that returned from the $x$- and $y$-cavities to the BS. The notation of the quadratures of the laser between the ITMs and EMS is illustrated in Fig. 2.
We need to emphasize that we do not discuss the power recycling, the signal recycling, or the squeezed state input techniques [8, 9, 22, 23, 24] in this paper. Our focus is solely on re-evaluating a simple Fabri-Pérot gravitational-wave detector. Therefore, within this paper, we cannot discuss the recent findings regarding the “violation of the standard quantum limit” in LIGO gravitational-wave detectors [10, 11]. However, we expect that the arguments presented in this paper can be extended to include techniques such as power recycling, signal recycling, or squeezed state input techniques. In this regard, we are confident that the arguments in this paper are meaningful.
The organization of this paper is as follows. In Sec. II, we summarize the basic notation that is used in this paper. In Sec. III, we develop the general arguments of the DC readout scheme. In Sec. IV, we derive the input-output relation without any specification of the small mirror displacements. In Sec. V, we consider the Heisenberg equations of motion to determine the small mirror displacements and their solution. In Sec. VI, we derive the final input-output relation by the specifications of the small mirror displacements through the Heisenberg equation of motion. We also discuss the comparison with the noise spectral densities in Ref. [8] also given in this section. In Sec. VII, we discuss the expectation value and the stationary noise spectral density in the DC readout from the results in Secs. III and VI. In this section, the resulting noise spectral density shows that the shot noise in the high-frequency region does not decrease even if the injected laser power is increased. In Sec. VIII, we consider changing the tuning point of the interferometer to recover the ideal noise spectral density in Ref. [8].
Some notations which we use within this paper are illustrated in Figs. 1 and 2. In this paper, some numerical values appear in the main text to provide an estimation of the order of magnitude for certain variables, but these values are merely estimation measures and have nothing to do with a specific gravitational-wave detector.
II Preliminary
II.1 Electric field notation
As in the usual quantum electrodynamics, the one-dimensional electric field operator $\hat{E}_{a}(t-z)$ at time $t$ and the length $z$ to the propagation direction in interferometers is described by
$\hat{E}_{a}(t-z)=\hat{E}_{a}^{(+)}(t-z)+\hat{E}_{a}^{(-)}(t-z),$ (1)
$\hat{E}_{a}^{(-)}(t-z)=\left[\hat{E}_{a}^{(+)}(t-z)\right]^{\dagger},$ (2)
$\hat{E}_{a}^{(+)}(t-z)=\int_{0}^{\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\hat{a}(\omega)e^{-i\omega(t-z)},$ (3)
where $\hat{a}(\omega)$ is the photon annihilation operator associated with the electric field $\hat{E}_{a}(t-z)$, which satisfies the commutation relation
$\left[\hat{a}(\omega),\hat{a}^{\dagger}(\omega^{\prime})\right]=2\pi\delta(\omega-\omega^{\prime}),$ (4)
$\left[\hat{a}(\omega),\hat{a}(\omega^{\prime})\right]=\left[\hat{a}^{\dagger}(\omega),\hat{a}^{\dagger}(\omega^{\prime})\right]=0.$ (5)
$\cal A$} is the cross-sectional area of the optical beam. To discuss the input-output relation of the interferometer, based on one-photon formulation, it is convenient to introduce the operator $\hat{A}(\omega)$ defined by
$\hat{A}(\omega):=\hat{a}(\omega)\Theta(\omega)+\hat{a}^{\dagger}(-\omega)\Theta(-\omega)$ (6)
so that the electric field (1) is represented as
$\hat{E}_{a}(t)=\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\hat{A}(\omega)e^{-i\omega t},$ (7)
where $\Theta(\omega)$ is the Heaviside step function
[…]
Due to the property of the Dirac $\delta$-function $\int_{-\infty}^{+\infty}dte^{+i(\omega^{\prime}-\omega)t}=2\pi\delta(\omega^{\prime}-\omega)$, the inverse relation of Eq. (7) is given by
$\hat{A}(\omega)=\sqrt{\frac{{\cal A}c}{2\pi\hbar|\omega|}}\int_{-\infty}^{+\infty}dte^{+i\omega t}\hat{E}_{a}(t).$ (11)
Therefore, the operator $\hat{A}(\omega)$ includes complete information of the electric field operator $\hat{E}_{a}(t)$ and is convenient to derive the input-output relation of simple interferometers.
From the commutation relations of the quadrature operator $\hat{a}(\omega)$ and $\hat{a}^{\dagger}(\omega)$ defined in Eq. (4) and (5), the commutation relations of electric fields $\hat{E}^{(\pm)}(t)$ are given by
$\left[\hat{E}^{(+)}_{a}(t),\hat{E}^{(+)}_{a}(t^{\prime})\right]$ $=$ $\left[\hat{E}^{(-)}_{a}(t),\hat{E}^{(-)}_{a}(t^{\prime})\right]=0,$ (12)
$\left[\hat{E}^{(+)}_{a}(t),\hat{E}^{(-)}_{a}(t^{\prime})\right]$ $=$ $\frac{2\pi\hbar}{{\cal A}c}\int_{0}^{+\infty}\frac{d\omega}{2\pi}\omega e^{-i\omega(t-t^{\prime})}$ (13)
$=:$ $\frac{2\pi\hbar}{{\cal A}c}\Delta_{a}(t-t^{\prime}).$
The subscription “aa” of the function $\Delta_{a}(t-t^{\prime})$ indicates that this is the vacuum fluctuation originating from the electric field $\hat{E}_{a}$ with the quadrature $\hat{a}(\omega)$.
We note that the function $\Delta_{a}(t-t^{\prime})$ has an ultraviolet divergence. However, in the actual measurements of the time sequence of the variables, the time in a measurement is discrete with a finite time bin. This time bin gives the maximum frequency $\omega_{\max}$, which becomes the natural ultraviolet cut-off of the frequency in the obtained data. Incidentally, in the actual measurements of the time sequence of the variables, the whole measurement time is also finite. It gives the minimum frequency $\omega_{\min}$ which corresponds to a natural infrared cut-off in frequency. Therefore, we may regard that the integration range over $\omega$ in the definition of the function $\Delta_{a}(t-t^{\prime})$ in Eq. (13) is $\omega_{\min},\omega_{\max}]$ instead of $0,+\infty]$. For this reason, throughout this paper, we do not regard the divergence in the definition of the function $\Delta_{a}(t-t^{\prime})$ as a serious one. We apply similar arguments when we evaluate the averaged laser power $I_{0}$ in Sec. IV.5.
To discuss the quantum properties of the laser, we have to specify the quantum state of the electric field, which is expressed by Eqs. (1)–(3). One of the quantum states of the electric fields which considered within this paper is the vacuum state $|0\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
$\hat{a}(\omega)|0\rangle:=0\quad{}^{\forall}\omega>0.$ (14)
On the other hand, we also consider the coherent state $|\alpha\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
$\hat{a}(\omega)|\alpha\rangle_{a}:=\alpha(\omega)|\alpha\rangle_{a},$ (15)
where $\alpha(\omega)$ is a complex function of $\omega$ which has the dimension $\mbox{Hz}]^{-1/2}$. As well-known, the coherent state $|\alpha\rangle_{a}$ and the vacuum state $|0\rangle_{a}$ are related through the displacement operator $D_{a}[\alpha]$ as [8]
$|\alpha\rangle_{a}$ $=$ $D_{a}[\alpha]|0\rangle_{a},$ (16)
$D_{a}[\alpha]$ $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)\hat{a}^{\dagger}(\omega)-\alpha^{*}(\omega)\hat{a}(\omega)\right)\right].$
Here, we note that the subscription “aa” in the state $|0\rangle_{a}$ and $|\alpha\rangle_{a}$ indicate that these states are associated with the electric field operator $\hat{E}_{a}(t)$, with the quadrature $\hat{a}(\omega)$.
II.2 Multi-mode number and power operators
In this paper, we examine the models of photodetection in which the photocurrent is proportional to Glauber’s multi-mode photon number $\hat{N}_{b}(t)$ defined by
$\hat{N}_{b}(t)$
$:=$ $\frac{\kappa_{n}c}{2\pi\hbar}{\cal A}\hat{E}^{(-)}_{b}(t)\hat{E}^{(+)}_{b}(t)$
$=$ $\int_{0}^{\infty}\frac{d\omega_{1}}{2\pi}\int_{0}^{\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\hat{b}^{\dagger}(\omega_{1})\hat{b}(\omega_{2})e^{+i(\omega_{1}-\omega_{2})t},$
where $\hat{E^{(\pm)}}_{b}(t)$ are the positive and negative frequency parts of the output optical electric field $\hat{E}_{b}(t)$ of the laser and $\kappa_{n}$ is the phenomenological coefficients of Glauber’s photon number $\hat{N}_{b}(t)$ and the photocurrent which includes the quantum efficiency. We note that $\kappa_{n}$ has the dimension of [time].
The reasons why we regard the Glauber photon number (LABEL:eq:Glauber-photon-number) as the direct observable in the photodetection were extensively discussed in Ref. [19]. Although the number operator for the single-mode photon is defined by $\hat{n}(\omega):=\hat{a}^{\dagger}(\omega)\hat{a}(\omega)$, the superposition of the electric field operator is possible within the field equations. In contrast, the superposition of $\hat{n}(\omega)$ is not possible within the field equations. A natural extension of the number operator to the multi-mode electric field is the above Glauber photon number (LABEL:eq:Glauber-photon-number).
On the other hand, in the gravitational-wave detection community, it is commonly regarded as the probability of the excitation of the photocurrent is proportional to the power $\hat{P}_{b}(t)\propto(\hat{E}_{b}(t))^{2}/(4\pi)$ of the output optical field $\hat{E}_{b}(t)$. However, even if we consider the model of photodetection in which the photocurrent is proportional to the power operator $\hat{P}_{b}$, we can reach the same conclusion as in Sec. III in the case of the DC readout. For this reason, within this paper, we mainly examine the models of photodetection in which the photocurrents are proportional to Glauber’s multi-mode photon number (LABEL:eq:Glauber-photon-number), for simplicity. We also check the model in which the photocurrent is proportional to the power $\hat{P}_{b}(t)$ of the optical field $\hat{E}_{b}$ in Appendix A. These models yield the same results within this paper.
III General arguments for the DC-readout scheme
In this paper, we consider the situation where the photodetector measures the output field $\hat{E}_{b}(t)$ whose quadrature $\hat{b}(\omega)$. This quadrature $\hat{b}(\omega)$ is
$\hat{b}(\omega)=:\langle\hat{b}(\omega)\rangle+\hat{b}_{n}(\omega),$ (19)
where $\langle\hat{b}(\omega)\rangle$ is the expectation value of the quadrature $\hat{b}(\omega)$ associated with the quantum state $|\mbox{in}\rangle$ injected to the interferometer and $\hat{b}_{n}(\omega)$ is the quantum operator which expresses the quantum noise of the quadrature $\hat{b}(\omega)$ which satisfies $\langle\hat{b}_{n}(\omega)\rangle=0$. Here, we express the $\langle\hat{b}(\omega)\rangle$ as
$\langle\hat{b}(\omega)\rangle={{\mathfrak{A}}}(\omega)+{{\mathfrak{B}}}2\pi\delta(\omega-\omega_{0}),$ (20)
where ${{\mathfrak{A}}}(\omega)$ is a classical complex function which includes gravitational-wave signals as shown in Sec. VI. $\mathfrak{B}$}} is a complex number which corresponds to the amplitude of the classical carrier field with the central frequency $\omega_{0}$.
Since we assume that the photodetector measures the quantum operator $\hat{N}_{b}(t)$ defined by Eq. (LABEL:eq:Glauber-photon-number), the Fourier transformation of the observed data is given by
$\hat{{{\mathcal{N}}}}_{b}(\omega)$ $:=$ $\int_{-\infty}^{+\infty}dt\frac{\kappa_{n}c}{2\pi\hbar}{\cal A}\hat{E}^{(-)}_{b}(t)\hat{E}^{(+)}_{b}(t)e^{+i\omega t}$
$=$ $\kappa_{n}\int_{0}^{\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega_{1}+\omega)|}\hat{b}^{\dagger}(\omega_{1})\hat{b}(\omega_{1}+\omega).$
Substituting Eq. (19) and (20) into Eq. (LABEL:eq:Fourier-Galuber-photon-number-expression), we obtain
$\hat{{{\mathcal{N}}}}_{b}(\omega)$ $=$ $\kappa_{n}\omega_{0}|{{\mathfrak{B}}}|^{2}2\pi\delta(\omega)$ (23)
$+\kappa_{n}\left({{\mathfrak{B}}}\sqrt{|\omega_{0}(\omega_{0}-\omega)|}{{\mathfrak{A}}}^{*}(\omega_{0}-\omega)\right$.
$\quad\quad\quad\left.+{{\mathfrak{B}}}^{*}\sqrt{|\omega_{0}(\omega_{0}+\omega)|}{{\mathfrak{A}}}(\omega_{0}+\omega)\right)$
$+\kappa_{n}\left({{\mathfrak{B}}}\sqrt{|\omega_{0}(\omega_{0}-\omega)|}\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\right$.
$\quad\quad\quad\left.+{{\mathfrak{B}}}^{*}\sqrt{|\omega_{0}(\omega_{0}+\omega)|}\hat{b}_{n}(\omega_{0}+\omega)\right)$
$+O\left(|{{\mathfrak{B}}}|^{0}\right).$
From this, the expectation value of the operator $\hat{{{\mathcal{N}}}}_{b}(\omega)$ is given by
$\langle\hat{{{\mathcal{N}}}}_{b}(\omega)\rangle$ $:=$ $\langle\mbox{in}|\hat{{{\mathcal{N}}}}_{b}(\omega)|\mbox{in}\rangle$ (24)
$=$ $\kappa_{n}\omega_{0}|{{\mathfrak{B}}}|^{2}2\pi\delta(\omega)$
$+\kappa_{n}\left({{\mathfrak{B}}}\sqrt{|\omega_{0}(\omega_{0}-\omega)|}{{\mathfrak{A}}}^{*}(\omega_{0}-\omega)\right$.
$\quad\quad\quad\left.+{{\mathfrak{B}}}^{*}\sqrt{|\omega_{0}(\omega_{0}+\omega)|}{{\mathfrak{A}}}(\omega_{0}+\omega)\right)$
$+O\left(|{{\mathfrak{B}}}|^{0}\right).$
As commented above, it will be shown in Sec. VI that ${{\mathfrak{A}}}(\omega)$ includes gravitational-wave signals. On the other hand, in the expectation value (24), the term of order $|{{\mathfrak{B}}}|^{2}$ is the amplitude of the classical carrier field of the laser, which is predictable. If the amplitude $|{{\mathfrak{B}}}|$ of the classical carrier field is sufficiently large, the leading term in Eq. (24) is of order $|{{\mathfrak{B}}}|^{2}$, and the second leading order is the term of $|{{\mathfrak{B}}}|^{1}$ which includes gravitational-wave signals ${{\mathfrak{A}}}(\omega)$. The remaining terms are not interesting in the DC readout scheme. Since the leading term of order $|{{\mathfrak{B}}}|^{2}$ is classical, predictable, and measurable, we can subtract the term of order $|{{\mathfrak{B}}}|^{2}$ to extract the sub-leading term of $|{{\mathfrak{B}}}|^{1}$ which includes the gravitational-wave signal ${{\mathfrak{A}}}(\omega)$. Thus, we can measure the signal term of order $|{{\mathfrak{B}}}|^{1}$ by the subtraction of the leading term of order $|{{\mathfrak{B}}}|^{2}$ from the expectation value (24). From this consideration, we may define the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\omega)$ for the signal ${{\mathfrak{A}}}(\omega)$ as
$\hat{s}_{{{\mathcal{N}}}_{b}}(\omega):=\frac{1}{\kappa_{n}}\hat{{{\mathcal{N}}}}_{b}(\omega)-\omega_{0}|{{\mathfrak{B}}}|^{2}2\pi\delta(\omega)$ (25)
so that its expectation value is given by
$\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\omega)\rangle$ $=$ ${{\mathfrak{B}}}\sqrt{|\omega_{0}(\omega_{0}-\omega)|}{{\mathfrak{A}}}^{*}(\omega_{0}-\omega)$ (26)
$+{{\mathfrak{B}}}^{*}\sqrt{|\omega_{0}(\omega_{0}+\omega)|}{{\mathfrak{A}}}(\omega_{0}+\omega)$
$+O\left(\left|{{\mathfrak{B}}}\right|^{0}\right).$
Here, we ignore the remaining term $O\left(\left|{{\mathfrak{B}}}\right|^{0}\right)$. The time-domain version of this signal operator is given by
$\hat{s}_{N_{b}}(t):=\frac{1}{\kappa_{n}}\hat{N}_{b}(t)-\omega_{0}|{{\mathfrak{B}}}|^{2}.$ (27)
From the signal operator $\hat{s}_{N_{b}}(t)$ defined by Eq. (27), we can define the noise operator for this measurement scheme as
$\hat{s}_{Nn}(t):=\hat{s}_{N_{b}}(t)-\langle\hat{s}_{N_{b}}(t)\rangle.$ (28)
Through Eqs. (23) and (27), the noise operator $\hat{s}_{Nn}(t)$ defined by Eq. (25) is given by
$\hat{s}_{Nn}(t)$ (29)
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\left\{+{{\mathfrak{B}}}\sqrt{|(\omega_{0}-\omega)\omega_{0}|}\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\right$.
$\quad\quad\quad\quad\quad\left.+{{\mathfrak{B}}}^{*}\sqrt{|\omega_{0}(\omega_{0}+\omega)|}\hat{b}_{n}(\omega_{0}+\omega)\right\}e^{-i\omega t}$
$+O\left(|{{\mathfrak{B}}}|^{0}\right).$
As the noise estimation, we consider the time-averaged (stationary) noise correlation function $C_{(av)s_{Nn}}(\tau)$ defined by
$C_{(av)s_{Nn}}(\tau)$
$:=$ $\lim_{T\rightarrow}\frac{1}{T}\int_{-T/2}^{T/2}dt\frac{1}{2}\langle\mbox{in}|\hat{s}_{Nn}(t+\tau)\hat{s}_{Nn}(t)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\hat{s}_{Nn}(t)\hat{s}_{Nn}(t+\tau)|\mbox{in}\rangle.$
From this definition of the time-averaged noise correlation function (LABEL:eq:stationary-noise-correlation-function-def), we define the noise spectral density $S_{s_{Nn}}(\omega)$ as the Fourier transformation of $C_{(av)s_{Nn}}(\tau)$ as
$S_{s_{Nn}}(\omega):=\int_{-\infty}^{+\infty}d\tau C_{(av)s_{Nn}}(\tau)e^{+i\omega\tau}.$ (31)
Substituting the expression (29) of the noise operator $\hat{s}_{Nn}(t)$ into Eq. (LABEL:eq:stationary-noise-correlation-function-def), we reach the expression of the stationary noise-spectral density $S_{s_{Nn}}(\omega)$ as
$S_{s_{Nn}}(\omega)$ $=$ $\frac{1}{2}\omega_{0}\int_{0}^{\infty}\frac{d\omega_{1}}{2\pi}\left[{{\mathfrak{B}}}^{2}\sqrt{|(\omega_{0}-\omega)\omega_{1}|}f(\omega_{0}-\omega_{1}+\omega)\left\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{1})+\hat{b}_{n}^{\dagger}(\omega_{1})\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\right\rangle\right$. (32)
$\quad\quad\quad\quad\quad\quad\left.+|{{\mathfrak{B}}}|^{2}\sqrt{|(\omega_{0}+\omega)\omega_{1}|}f(\omega_{0}-\omega_{1}+\omega)\left\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}^{\dagger}(\omega_{1})+\hat{b}_{n}^{\dagger}(\omega_{1})\hat{b}_{n}(\omega_{0}+\omega)\right\rangle\right$.
$\quad\quad\quad\quad\quad\quad\left.+|{{\mathfrak{B}}}|^{2}\sqrt{|(\omega_{0}-\omega)\omega_{1}|}f(\omega_{0}-\omega_{1}-\omega)\left\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}(\omega_{1})+\hat{b}_{n}(\omega_{1})\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\right\rangle\right$.
$\quad\quad\quad\quad\quad\quad\left.+({{\mathfrak{B}}}^{*})^{2}\sqrt{|(\omega_{0}+\omega)\omega_{1}|}f(\omega_{0}-\omega_{1}-\omega)\left\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}(\omega_{1})+\hat{b}_{n}(\omega_{1})\hat{b}_{n}(\omega_{0}+\omega)\right\rangle\right].$
Here, we defined the one-point support function $f(a)$ of $a\in{{\mathbb{R}}}$ by
$f(a)$ $:=$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{-T/2}^{T/2}dte^{-iat}$ (33)
$=$ $\left\{\begin{array}[]{ccccc}1&&\mbox{for}&&a=0,\\ 0&&\mbox{for}&&a\neq 0,\end{array}\right$. (36)
as discussed in Ref. [19]. We also note that for a finite function $g(a)$,
$\int_{-\infty}^{+\infty}\delta(a)g(a)f(a)da=g(0),$ (37)
$\int_{-\infty}^{+\infty}\delta(b)g(b)f(a)db=f(a)g(0).$ (38)
From the input-output relation which will be derived in Sec. VI, we can confirm the following expectation values:
$\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{0}-\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}),$ (39)
$\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}^{\dagger}(\omega_{0}-\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}),$ (40)
$\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}(\omega_{0}+\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}),$ (41)
$\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}(\omega_{0}+\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}).$ (42)
Through these expectation values (39)–(42) and the properties of the one-point support function $f(a)$ summarized in Eqs. (33)–(38), we may write
$2\pi\delta(\omega-\omega^{\prime})S_{Nn}(\omega)$ $=$ $\omega_{0}|{{\mathfrak{B}}}|^{2}\left[e^{+2i\Theta}\sqrt{|(\omega_{0}-\omega)(\omega_{0}+\omega^{\prime})|}\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{0}+\omega^{\prime})\rangle\right$. (43)
$\quad\quad\quad\quad\left.+\frac{1}{2}\sqrt{|(\omega_{0}+\omega)(\omega_{0}+\omega^{\prime})|}\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}^{\dagger}(\omega_{0}+\omega^{\prime})+\hat{b}_{n}^{\dagger}(\omega_{0}+\omega^{\prime})\hat{b}_{n}(\omega_{0}+\omega)\rangle\right$.
$\quad\quad\quad\quad\left.+\frac{1}{2}\sqrt{|(\omega_{0}-\omega)(\omega_{0}-\omega^{\prime})|}\langle\hat{b}_{n}(\omega_{0}-\omega^{\prime})\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)+\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}(\omega_{0}-\omega^{\prime})\rangle\right$.
$\quad\quad\quad\quad\left.+e^{-2i\Theta}\sqrt{|(\omega_{0}+\omega)(\omega_{0}-\omega^{\prime})|}\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}(\omega_{0}-\omega^{\prime})\rangle\right]$
$+O\left(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0}\right),$
where we defined ${{\mathfrak{B}}}=:|{{\mathfrak{B}}}|e^{i\Theta}$. Here, we note that the expression (43) is described by the quadrature with the frequency $\omega_{0}\pm\omega$. This is the motivation of the sideband picture, which describes the quantum fluctuations around the central frequency $\omega_{0}$ of the incident laser. We follow the historical notation in which $\Omega$ denotes the sideband frequency. Then, we define the upper- and lower-sideband quadrature $\hat{b}_{n\pm}(\Omega)$ by
$\hat{b}_{n\pm}(\Omega):=\hat{b}_{n}(\omega_{0}\pm\Omega).$ (44)
Furthermore, the amplitude quadrature $\hat{b}_{n1}(\Omega)$ and the phase quadrature $\hat{b}_{n2}(\Omega)$ are often used, which are defined by
$\hat{b}_{n+}(\Omega)=:\frac{1}{\sqrt{2}}\left(\hat{b}_{n1}(\Omega)+i\hat{b}_{n2}(\Omega)\right),$ (45)
$\hat{b}_{n-}(\Omega)=:\frac{1}{\sqrt{2}}\left(\hat{b}_{n1}^{\dagger}(\Omega)+i\hat{b}_{n2}^{\dagger}(\Omega)\right).$ (46)
Moreover, we also introduce the operator $\hat{b}_{n\Theta}(\Omega)$ by
$\hat{b}_{n\Theta}(\Omega)=\cos\Theta\hat{b}_{n1}(\Omega)+\sin\Theta\hat{b}_{n2}(\Omega).$ (47)
Through the operator $\hat{b}_{n\Theta}(\Omega)$, the noise spectral density $S_{s_{Nn}}(\Omega)$ defined by Eq. (43) is given by
$2\pi\delta(\Omega-\Omega^{\prime})S_{s_{Nn}}(\Omega)$ (48)
$=$ $\omega_{0}|{{\mathfrak{B}}}|^{2}\left\langle\hat{b}_{n\Theta}(\Omega)\hat{b}_{n\Theta}^{\dagger}(\Omega^{\prime})+\hat{b}_{n\Theta}^{\dagger}(\Omega^{\prime})\hat{b}_{n\Theta}(\Omega)\right\rangle$
$+O\left(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0}\right).$
Here, we note that
$\left[\hat{b}_{n\Theta}(\Omega),\hat{b}_{n\Theta}^{\dagger}(\Omega)\right]=0.$ (49)
The formula for the stationary noise-spectral density $S_{s_{Nn}}(\Omega)$ is the general result under the premise of the DC-readout scheme. To derive the result (48), the sufficiently large amplitude $|{{\mathfrak{B}}}|$ of the classical carrier is essential. Due to this large $|{{\mathfrak{B}}}|$, we can neglect the residual term $O(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0})$. This is the main difference of the balanced homodyne detection discussed in Ref. [19].
IV Input-Output relation of the Fabri-Pérot interferometer
In this section, we derive the input-output relation of the Fabri-Pérot interferometer, which is used in the setup of gravitational-wave detectors. The Fabri-Pérot interferometer consists of the beam splitter (BS), intermediate masses (ITMs), and the end mirrors (EMs). The injected laser to the Fabri-Pérot interferometer is separated into the directions of the $x$- and the $y$-arms. The laser is amplified between the ITM and EM. In gravitational-wave detectors, BS, ITMs, and EMs are suspended through the vibration-isolation system to measure the relative positions of these mirrors precisely. In these relative positions, gravitational-wave signals are included.
IV.1 Mirror Displacements
To describe the relative positions of BS, ITMs, and EMs, we introduce the proper reference frame [25] whose origin is BS. As depicted in Fig. 1, we denote the coordinate values in the proper reference frame of ITM (XITM) and EM (XEM) along $x$-arm by $l_{x}+\hat{X}_{XITM}$ and $l_{x}+L+\hat{X}_{XEM}$, respectively. Here, $l_{x}$ is the distance between BS and XITM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals. $L$ is the distance between XITM and XEM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals. We regard $l_{x}$ and $L$ as classical distances. In addition to the $l_{x}$ and $L$, we introduce the quantum displacement $\hat{X}_{XITM}$ and $\hat{X}_{EM}$, which are induced by the injected laser and other forces, including the gravitational-wave signal. In this paper, we regard $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ as quantum operators that describe the quantum mirror positions. In Sec. V, the operators $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ are determined as the solution to the Heisenberg equation of motion.
Similarly, we also introduce the coordinate values in the proper reference frame of ITM (YITM) and EM (YEM) along $y$-arm $l_{y}+\hat{X}_{YITM}$ and $l_{y}+L+\hat{X}_{YEM}$, respectively. We also regard $\hat{X}_{YITM}$ and $\hat{X}_{YEM}$ as describing quantum displacements, which are induced by the injected laser and other forces, including the gravitational-wave signal. As in the case of the operators $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$, in Sec. V, the operators $\hat{X}_{YITM}$ and $\hat{X}_{YEM}$ are determined as the solution to the Heisenberg equation of motion.
Based on the above setup, we consider the electric field of the laser through quantum electrodynamics.
IV.2 Beam Splitter Junction
First, we consider the junction conditions for optical quadratures at BS. Following the notation depicted in Fig. 1, the final output electric field operator $\hat{E}_{b}(t)$ is given by
$\hat{E}_{b}(t)=\frac{1}{\sqrt{2}}\left[\hat{E}_{c_{y}^{\prime}}(t)-\hat{E}_{c_{x}^{\prime}}(t)\right],$ (50)
where $\hat{E}_{c_{y}^{\prime}}(t)$ and $\hat{E}_{c_{x}^{\prime}}(t)$ are electric field operators injected from the $y$-arm and $x$-arm to BS, respectively. Here, we defined
$\hat{B}(\omega)$ $:=$ $\hat{b}(\omega)\Theta(\omega)+\hat{b}^{\dagger}(-\omega)\Theta(-\omega),$ (51)
$\hat{C}_{x}^{\prime}(\omega)$ $:=$ $\hat{c}_{x}^{\prime}(\omega)\Theta(\omega)+\hat{c}_{x}^{{}^{\prime}\dagger}(-\omega)\Theta(-\omega),$ (52)
$\hat{C}_{y}^{\prime}(\omega)$ $:=$ $\hat{c}_{y}^{\prime}(\omega)\Theta(\omega)+\hat{c}_{y}^{{}^{\prime}\dagger}(-\omega)\Theta(-\omega)$ (53)
as in Eq. (6). In terms of the operators $\hat{B}(\omega)$, $\hat{C}_{x}^{\prime}(\omega)$, and $\hat{C}_{y}^{\prime}(\omega)$, the relation (50) is given by
$\hat{B}(\omega)=\frac{1}{\sqrt{2}}\left(\hat{C}_{y}^{\prime}(\omega)-\hat{C}_{x}^{\prime}(\omega)\right).$ (54)
Similarly, the electric-field operators $\hat{E}_{c_{y}}(t)$ and $\hat{E}_{c_{x}}(t)$, that propagated from BS to each arm, are also given by the input field operators $\hat{E}_{d}(t)$ and $\hat{E}_{a}(t)$, which are injected from the light source and photodetectors to BS, respectively, as follows:
$\hat{E}_{c_{x}}(t)$ $=$ $\frac{1}{\sqrt{2}}\left(\hat{E}_{d}(t)-\hat{E}_{a}(t)\right),$ (55)
$\hat{E}_{c_{y}}(t)$ $=$ $\frac{1}{\sqrt{2}}\left(\hat{E}_{d}(t)+\hat{E}_{a}(t)\right).$ (56)
In terms of the quadrature as in Eq. (6), these relations yield
$\hat{C}_{x}(\omega)$ $=$ $\frac{1}{\sqrt{2}}\left(\hat{D}(\omega)-\hat{A}(\omega)\right),$ (57)
$\hat{C}_{y}(\omega)$ $=$ $\frac{1}{\sqrt{2}}\left(\hat{D}(\omega)+\hat{A}(\omega)\right),$ (58)
where we defined the operators
$\hat{C}_{x}(\omega)$ $:=$ $\hat{c}_{x}(\omega)\Theta(\omega)+\hat{c}_{x}^{\dagger}(-\omega)\Theta(-\omega),$ (59)
$\hat{C}_{y}(\omega)$ $:=$ $\hat{c}_{y}(\omega)\Theta(\omega)+\hat{c}_{y}^{\dagger}(-\omega)\Theta(-\omega),$ (60)
$\hat{D}(\omega)$ $:=$ $\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ (61)
as in Eq. (6). The notations of the quadratures are also depicted in Fig. 1.
IV.3 Arm Propagation
Now, we consider the propagation effects along the $x$- and $y$-arms, respectively.
IV.3.1 Propagation between BS and ITMs
From the propagation effect from XITM (YITM) to BS, in the notation depicted in Fig. 1, we have
$\hat{E}_{c^{\prime}_{x}}(t)$ $=$ $\hat{E}_{f^{\prime}_{x}}\left[t-\frac{l_{x}+\hat{X}_{XITM}(t-l_{x}/c)}{c}\right]$ (62)
$=$ $\hat{E}_{f^{\prime}_{x}}\left[t-\tau_{x}^{\prime}-\frac{\hat{X}_{XITM}(t-\tau_{x}^{\prime})}{c}\right],$
$\hat{E}_{c^{\prime}_{y}}(t)$ $=$ $\hat{E}_{f^{\prime}_{y}}\left[t-\tau_{y}^{\prime}-\frac{\hat{X}_{YITM}(t-\tau_{y}^{\prime})}{c}\right]$ (63)
as the output electric field operators $\hat{E}_{c^{\prime}_{x}}(t)$ and $\hat{E}_{c^{\prime}_{y}}(t)$ from each arm. Here, we defined $l_{x}=:c\tau_{x}^{\prime}$ and $l_{y}=:c\tau_{y}^{\prime}$. In the Fourier space, we Eq. (62) is given by
$\hat{E}_{c^{\prime}_{x}}(t)$ $=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{2\pi\hbar|\omega_{1}|}{{\cal A}c}}\hat{C}_{x}^{\prime}(\omega_{1})e^{-i\omega_{1}t}$ (64)
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{\frac{2\pi\hbar|\omega_{2}|}{{\cal A}c}}\hat{F}_{x}^{\prime}(\omega_{2})$
$\quad\times\exp\left[-i\omega_{2}\left(t-\tau_{x}^{\prime}-\frac{\hat{X}_{XITM}(t-\tau_{x}^{\prime})}{c}\right)\right]$
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{\frac{2\pi\hbar|\omega_{2}|}{{\cal A}c}}\hat{F}_{x}^{\prime}(\omega_{2})e^{-i\omega_{2}t}e^{+i\omega_{2}\tau_{x}^{\prime}}$
$\quad\times\left(1+i\frac{\omega_{2}}{c}\int_{-\infty}^{+\infty}\frac{d\omega_{3}}{2\pi}\hat{Z}_{XITM}(\omega_{3})e^{-i\omega_{3}(t-\tau_{x}^{\prime})}\right)$
$+O\left(\left(\hat{X}_{XITM}\right)^{2}\right),$
where we used
$\hat{X}_{XITM}$ $=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{3}}{2\pi}\hat{Z}_{XITM}(\omega_{3})e^{-i\omega_{3}t},$ (65)
$\hat{X}_{YITM}$ $=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{3}}{2\pi}\hat{Z}_{YITM}(\omega_{3})e^{-i\omega_{3}t}.$ (66)
Operating $\int_{-\infty}^{+\infty}dte^{+i\omega t}$ to Eq. (64), we obtain
$\hat{C}_{x}^{\prime}(\omega)$ (67)
$=$ $e^{+i\omega\tau_{x}^{\prime}}\hat{F}_{x}^{\prime}(\omega)$
$+ie^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{\frac{|\omega_{2}|}{|\omega|}}\frac{\omega_{2}}{c}\hat{F}_{x}^{\prime}(\omega_{2})\hat{Z}_{XITM}(\omega-\omega_{2})$
$+O\left(\left(\hat{X}_{XITM}\right)^{2}\right),$
for $x$-arm. Similarly, for $y$-arm, we obtain
$\hat{C}_{y}^{\prime}(\omega)$ (68)
$=$ $e^{+i\omega\tau_{y}^{\prime}}\hat{F}_{y}^{\prime}(\omega)$
$+ie^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{\frac{|\omega_{2}|}{|\omega|}}\frac{\omega_{2}}{c}\hat{F}_{y}^{\prime}(\omega_{2})\hat{Z}_{YITM}(\omega-\omega_{2})$
$+O\left(\left(\hat{X}_{YITM}\right)^{2}\right).$
Furthermore, the propagation effects of the laser yield the electric field in each arm by
$\hat{E}_{f_{x}}(t)$ $=$ $\hat{E}_{c_{x}}\left[t-\tau_{x}^{\prime}-\frac{\hat{X}_{XITM}(t)}{c}\right],$ (69)
$\hat{E}_{f_{y}}(t)$ $=$ $\hat{E}_{c_{y}}\left[t-\tau_{y}^{\prime}-\frac{\hat{X}_{YITM}(t)}{c}\right].$ (70)
We consider the Fourier transformation of Eqs. (69) and (70). As in the case of Eqs. (67) and (68), we consider the Fourier expansion as Eq. (7). Then, we obtain
$\hat{F}_{x}(\omega)$ (71)
$=$ $e^{+i\omega\tau_{x}^{\prime}}\hat{C}_{x}(\omega)$
$+i\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}e^{+i\omega_{1}\tau_{x}^{\prime}}\hat{C}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})$
$+O\left(\hat{X}_{XITM}^{2}\right).$
Similarly, we obtain
$\hat{F}_{y}(\omega)$ (72)
$=$ $e^{+i\omega\tau_{y}^{\prime}}\hat{C}_{y}(\omega)$
$+i\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}e^{+i\omega_{1}\tau_{y}^{\prime}}\hat{C}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})$
$+O\left(\hat{X}_{YITM}^{2}\right).$
IV.3.2 Junction at ITMs and cavity propagation
Next, we consider the arm propagation of the laser along each arm. We denote the notation of the electric field of the electric field as depicted in Fig. 2.
Figure 2: Arm propagation in the Fabry-Pérot interferometer and photoelectric quadratures notations. The quadratures $\hat{f}$ and $\hat{f}^{\prime}$ are those of the laser incident from the BS to ITM and the laser from ITM to BS, respectively. The $\hat{g}$ and $\hat{g}^{\prime}$ are the quadratures for the lasers from ITM to the Fabri-Pérot cavity and from the Fabri-Pérot cavity to ITM, respectively. The quadratures $\hat{j}$ and $\hat{j}^{\prime}$ are quadratures for lasers that reach the EM and are reflected by the EM, respectively.
We also denote the power reflection and transmission coefficients by $R$ and $T$, respectively. The amplitude reflection and transmission coefficients are chosen to be real, with signs $\{-\sqrt{R},+\sqrt{T}\}$, $\{-\sqrt{\tilde{R}},+\sqrt{\tilde{T}}\}$ for the light that impinges on a mirror from the outside of the cavity at ITMs and EMs, respectively, and $\{+\sqrt{R},+\sqrt{T}\}$, $\{+\sqrt{\tilde{R}},+\sqrt{\tilde{T}}\}$ for light that impinges from the inside of the cavity at ITMs and EMs, respectively. These satisfy the condition
$R+T=\tilde{R}+\tilde{T}=1.$ (73)
Then, the junction conditions for the electric field operators $\hat{E}_{f_{x,y}}$, $\hat{E}_{f^{\prime}_{x,y}}$, $\hat{E}_{g_{x,y}}$, and $\hat{E}_{g^{\prime}_{x,y}}$, are given by
$\hat{E}_{g_{x,y}}(t)=\sqrt{T}\hat{E}_{f_{x,y}}(t)+\sqrt{R}\hat{E}_{g^{\prime}_{x,y}}(t),$ (74)
$\hat{E}_{f^{\prime}_{x,y}}(t)=-\sqrt{R}\hat{E}_{f_{x,y}}(t)+\sqrt{T}\hat{E}_{g^{\prime}_{x,y}}(t).$ (75)
Next, we consider the propagation of the laser within the ITMs and EMs. In this paper, we assume that the EMs have perfect reflection, i.e.,
$\tilde{R}=1,\quad\tilde{T}=0.$ (76)
and we obtain
$\hat{E}_{j^{\prime}_{x,y}}(t)=\hat{E}_{j_{x,y}}(t).$ (77)
If we have to consider the loss model of the imperfection of EM as in Ref. [8], we have to change Eq. (77).
Furthermore, the derivations of the electric field relations between $\hat{E}_{g_{x,y}}$ , $\hat{E}_{j_{x,y}}$, $\hat{E}_{j^{\prime}_{x,y}}$, and $\hat{E}_{g^{\prime}_{x,y}}$ are delicate. However, the physical effects are just free propagation of the laser. This is because the time-dependence of the relative positions $\hat{X}_{x}(t):=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}(t):=\hat{X}_{YEM}-\hat{X}_{YITM}$. We have to be careful to treat this time dependence. However, we apply the perfect reflection of EMs and the relation of $\hat{E}_{g_{x,y}}$ and $\hat{E}_{g^{\prime}_{x,y}}$ should be identical to the Michelson interferometer from the setup depicted in Fig. 2. Then, we obtain
$\hat{E}_{g^{\prime}_{x}}(t)$ $=$ $\hat{E}_{g_{x}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)\right)\right],$ (78)
$\hat{E}_{g^{\prime}_{y}}(t)$ $=$ $\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right],$ (79)
where $\tau=L/c$. Furthermore, through Eqs. (74), (75), (78), and (79), we obtain the relations
$\hat{E}_{f^{\prime}_{x}}(t)$ $=$ $-\sqrt{1-T}\hat{E}_{f_{x}}(t)$ (80)
$+\sqrt{T}\hat{E}_{g_{x}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)\right)\right],$
$\hat{E}_{g_{x}}(t)$ $=$ $\sqrt{T}\hat{E}_{f_{x}}(t)$
$+\sqrt{1-T}\hat{E}_{g_{x}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)\right)\right],$
and
$\hat{E}_{f^{\prime}_{y}}(t)$ $=$ $-\sqrt{1-T}\hat{E}_{f_{y}}(t)$ (82)
$+\sqrt{T}\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right],$
$\hat{E}_{g_{y}}(t)$ $=$ $\sqrt{T}\hat{E}_{f_{y}}(t)$
$+\sqrt{1-T}\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right].$
Here, we note that Eq. (LABEL:eq:xarm-prop-g-and-f-g) (resp. (LABEL:eq:yarm-prop-g-and-f-g)) gives the relation between the electric fields $\hat{E}_{f_{x}}$ and $\hat{E}_{g_{x}}$ (resp. $\hat{E}_{f_{y}}$ and $\hat{E}_{g_{y}}$). If we can obtain the inverse relation of Eq. (LABEL:eq:xarm-prop-g-and-f-g) (resp. (LABEL:eq:yarm-prop-g-and-f-g)), we can obtain the relation between the electric fields $\hat{E}_{f^{\prime}_{x}}$ and $\hat{E}_{f_{x}}$ (resp. $\hat{E}_{f^{\prime}_{y}}$ and $\hat{E}_{f_{y}}$) through the substitution of this inverse relation of (LABEL:eq:xarm-prop-g-and-f-g) (resp. (LABEL:eq:yarm-prop-g-and-f-g)) into Eq. (80) (resp. (82)).
To obtain the expressions of $\hat{E}_{g_{x,y}}$ in terms of $\hat{E}_{f_{x,y}}$, the treatment of the Fourier transformation is convenient. First, we consider the Fourier transformation of Eq. (80). Introducing the Fourier transformation of the displacement variable $\hat{X}_{x,y}(t)$ as
$\hat{X}_{x,y}(t)$ $=:$ $\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\hat{Z}_{x,y}(\omega_{2})e^{-i\omega_{2}t},$ (84)
we only considered the effects of the linear level of $\hat{X}_{x}$ in Eq. (80). Furthermore, through the inverse Fourier transformation, we obtain
$\hat{F}^{\prime}_{x}(\omega)$ $=$ $-\sqrt{1-T}\hat{F}_{x}(\omega)+\sqrt{T}\hat{G}_{x}(\omega)e^{+2i\omega\tau}$ (85)
$+i\frac{2}{c}\sqrt{T}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}e^{+i(\omega_{1}+\omega)\tau}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}$
$\quad\quad\times\hat{G}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{x}\right)^{2}\right).$
Similarly, we consider the Fourier transformation of Eq. (LABEL:eq:xarm-prop-g-and-f-g):
$\hat{G}_{x}(\omega)$ $=$ $\sqrt{T}\hat{F}_{x}(\omega)+\sqrt{1-T}\hat{G}_{x}(\omega)e^{+2i\omega\tau}$ (86)
$+i\frac{2}{c}\sqrt{1-T}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}e^{+i(\omega_{1}+\omega)\tau}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}$
$\quad\quad\times\hat{G}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{x}\right)^{2}\right),$
where we used (84). Then, we obtain
$\hat{G}_{x}(\omega)$ $=$ $\sqrt{T}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\hat{F}_{x}(\omega)$ (87)
$+i\frac{2}{c}\sqrt{1-T}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}e^{+i(\omega_{1}+\omega)\tau}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\hat{G}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{x}\right)^{2}\right).$
The substitution of $\hat{G}_{x}(\omega_{1})$ in the left-hand side of Eq. (87) into $\hat{G}_{x}(\omega_{1})$ in the right-hand side of Eq. (87) yields
$\hat{G}_{x}(\omega)$ (88)
$=$ $\sqrt{T}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\hat{F}_{x}(\omega)$
$+i\frac{2}{c}\sqrt{T(1-T)}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}e^{+i(\omega_{1}+\omega)\tau}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{F}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{x}\right)^{2}\right).$
Substituting Eq. (88) into Eq. (85), we obtain
$\hat{F}^{\prime}_{x}(\omega)$ (89)
$=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}\hat{F}_{x}(\omega)$
$+i\frac{2}{c}Te^{+2i\omega\tau}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}e^{-i(\omega-\omega_{1})\tau}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{F}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{x}\right)^{2}\right).$
Furthermore, the substitution of Eq. (89) into Eq. (67) yields
$\hat{C}_{x}^{\prime}(\omega)$ $=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega\tau_{x}^{\prime}}\hat{F}_{x}(\omega)$ (90)
$+i\frac{2}{c}Te^{+2i\omega\tau}e^{+i\omega\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}$
$\quad\quad\quad\times\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}e^{-i(\omega-\omega_{1})\tau}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{F}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{2}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}\hat{F}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{XITM}\right)^{2},\left(\hat{X}_{x}\right)^{2},\hat{X}_{x}\hat{X}_{XITM}\right).$
For $y$-arm quadrature relation we replace $x\rightarrow y$ and $XITM\rightarrow YITM$ as
$\hat{C}_{y}^{\prime}(\omega)$ $=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega\tau_{y}^{\prime}}\hat{F}_{y}(\omega)$ (91)
$+i\frac{2}{c}Te^{+2i\omega\tau}e^{+i\omega\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}$
$\quad\quad\quad\times\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}e^{-i(\omega-\omega_{1})\tau}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{F}_{y}(\omega_{1})\hat{Z}_{y}(\omega-\omega_{1})$
$+e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}\hat{F}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{YITM}\right)^{2},\left(\hat{X}_{y}\right)^{2},\hat{X}_{y}\hat{X}_{YITM}\right).$
Moreover, the substitution Eq. (71) into Eq. (90) yields
$\hat{C}_{x}^{\prime}(\omega)$ $=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+2i\omega\tau_{x}^{\prime}}\hat{C}_{x}(\omega)$ (92)
$+ie^{+2i\omega\tau_{x}^{\prime}}e^{+2i\omega\tau}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}\hat{C}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})$
$+i\frac{2}{c}Te^{+i\omega\tau}e^{+i\omega\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}$
$\quad\quad\quad\times\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}e^{+i\omega_{1})\tau}e^{+i\omega_{1}\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{C}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})$
$+e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\times e^{+2i\omega_{1}\tau}e^{+i\omega_{1}\tau_{x}^{\prime}}\hat{C}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{XITM}\right)^{2},\left(\hat{X}_{x}\right)^{2},\hat{X}_{x}\hat{X}_{XITM}\right).$
Similarly, for $C_{y}^{\prime}(\omega)$, we have
$\hat{C}_{y}^{\prime}(\omega)$ $=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+2i\omega\tau_{y}^{\prime}}\hat{C}_{y}(\omega)$ (93)
$+ie^{+2i\omega\tau_{y}^{\prime}}e^{+2i\omega\tau}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}\hat{C}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})$
$+i\frac{2}{c}Te^{+i\omega\tau}e^{+i\omega\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}$
$\quad\quad\times\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}e^{+i\omega_{1}\tau}e^{+i\omega_{1}\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\hat{C}_{y}(\omega_{1})\hat{Z}_{y}(\omega-\omega_{1})$
$+e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\times e^{+2i\omega_{1}\tau}e^{+i\omega_{1}\tau_{y}^{\prime}}\hat{C}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})$
$+O\left(\left(\hat{X}_{YITM}\right)^{2},\left(\hat{X}_{y}\right)^{2},\hat{X}_{y}\hat{X}_{YITM}\right).$
IV.4 Input-output relation with mirrors’ motion
From Eqs. (54), (57), (58), and the arm propagation relations (92) and (93), we obtain
$\!\!\!\!\!\!\!\hat{B}(\omega)$ (94)
$=$ $\frac{1}{\sqrt{2}}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[e^{+i\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})}\hat{C}_{y}(\omega)-e^{-i\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})}\hat{C}_{x}(\omega)\right]$
$+i\frac{1}{\sqrt{2}}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}$
$\quad\quad\quad\times\left[e^{+i\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})}\hat{C}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})-e^{-i\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})}\hat{C}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})\right]$
$+i\frac{1}{\sqrt{2}}\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+i(\omega+\omega_{1})\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\quad\times\left[e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}}\hat{C}_{y}(\omega_{1})\hat{Z}_{y}(\omega-\omega_{1})-e^{-i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}}\hat{C}_{x}(\omega_{1})\hat{Z}_{x}(\omega-\omega_{1})\right]$
$+\frac{1}{\sqrt{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\quad\times\left[e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}}\hat{C}_{y}(\omega_{1})\hat{Z}_{YITM}(\omega-\omega_{1})-e^{-i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}}\hat{C}_{x}(\omega_{1})\hat{Z}_{XITM}(\omega-\omega_{1})\right]$
$+O\left(\left(\hat{X}\right)^{2}\right).$
The last integral in Eq. (94) which includes $\hat{Z}_{YITM}$ and $\hat{Z}_{XITM}$ are the retarded effects during the propagation of the laser from the intermediate masses to the beam splitter due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. The third integral in Eq. (94) which includes $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{XEM}-\hat{X}_{XITM}$, is also the retarded effect during the propagation of the laser in the Fabri-Pérot cavities. The second integral in Eq. (94) are the retarded effect during the propagation from the beam splitter to the intermediate masses due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. The first line is the direct shot noise of this interferometer. The last line in Eq. (94) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2})$.
Substituting Eqs. (57) and (58) into Eq. (94), we obtain the output field quadrature $\hat{B}(\omega)$. Furthermore, for our convention, we define $\hat{Z}_{com,diff}(\omega)$ by
$\hat{Z}_{x}(\omega)$ $=:$ $\hat{Z}_{com}(\omega)+\hat{Z}_{diff}(\omega),$ (95)
$\hat{Z}_{y}(\omega)$ $=:$ $\hat{Z}_{com}(\omega)-\hat{Z}_{diff}(\omega),$ (96)
where $\hat{Z}_{com}(\omega)$ represents the $x$-arm and $y$-arm common motion of the relative motion between the end-mirror and the intermediate mass and $\hat{Z}_{dif}(\omega)$ represents the $x$-arm and $y$-arm differential motion of the relative motion the end-mirror and intermediate mass. Furthermore, we also define $\hat{Z}_{comITM,diffITM}(\omega)$ by
$\hat{Z}_{XITM}(\omega)$ $=:$ $\hat{Z}_{comITM}(\omega)+\hat{Z}_{diffITM}(\omega),$ (97)
$\hat{Z}_{YITM}(\omega)$ $=:$ $\hat{Z}_{comITM}(\omega)-\hat{Z}_{diffITM}(\omega).$ (98)
Substituting Eqs. (95), (96), (97), and (98) into Eq. (94), we obtain
$\!\!\!\!\!\!\!\hat{B}(\omega)$ (99)
$=$ $\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}$
$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}(\omega)+\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{A}(\omega)\right]$
$+ie^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\frac{\omega_{1}}{c}$
$\quad\quad\quad\times\left[+\left(i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}(\omega_{1})+\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{A}(\omega_{1})\right)\hat{Z}_{comITM}(\omega-\omega_{1})\right$.
$\quad\quad\quad\quad\quad\left.-\left(\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}(\omega_{1})+i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{A}(\omega_{1})\right)\hat{Z}_{diffITM}(\omega-\omega_{1})\right]$
$+i\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+i(\omega+\omega_{1})\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\quad\times\left[+\left(i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{D}(\omega_{1})+\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{A}(\omega_{1})\right)\hat{Z}_{com}(\omega-\omega_{1})\right$.
$\quad\quad\quad\quad\quad\left.-\left(\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{D}(\omega_{1})+i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{A}(\omega_{1})\right)\hat{Z}_{diff}(\omega-\omega_{1})\right]$
$+\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}i\frac{\omega_{1}}{c}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\quad\times\left[\left(i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{D}(\omega_{1})+\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{A}(\omega_{1})\right)\hat{Z}_{comITM}(\omega-\omega_{1})\right$.
$\quad\quad\quad\quad\quad\left.-\left(\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{D}(\omega_{1})+i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\hat{A}(\omega_{1})\right)\hat{Z}_{diffITM}(\omega-\omega_{1})\right]$
$+O\left(\left(\hat{X}\right)^{2}\right).$
IV.5 Coherent state of the optical fields
Here, Eq. (99) implies that the output operator $\hat{B}$ is given by the operators $\hat{A}$, $\hat{D}$, $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$. Later, we see that the displacements $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$ are given by $\hat{A}$ and $\hat{D}$ together with the gravitational-wave signal through the equations of motions for end-mirrors and the intermediate mirrors. Therefore, to discuss the information from the output operator $\hat{B}$, we have to specify the quantum states associated with the operators $\hat{A}$ and $\hat{D}$. The state associated with the operator $\hat{A}$ is the state of the electric field that is injected from the anti-symmetric port. At this anti-symmetric port, the photo-detector is located as in Fig. 1. On the other hand, the state associated with the operator $\hat{D}$ is the state of the electric field that is injected from the symmetric port. At this symmetric port, the light source exists as depicted in Fig. 1. The total state of photon in the output port $\hat{B}$, i.e., $\hat{b}$, is determined by the specification of the states associated with the operators $\hat{D}$ and $\hat{A}$, i.e., the annihilation and creation operators $(\hat{d},\hat{d}^{\dagger})$ and $(\hat{a},\hat{a}^{\dagger})$, respectively.
Within this paper, we assume that there is no entanglement in the states associated with the operators $\hat{d}$ and $\hat{a}$. Furthermore, we assume that the state associated with the operator $\hat{d}$ is a coherent state with the complex amplitude $\alpha(\omega)$ and the state associated with the operator $\hat{a}$ is the vacuum state. Then, the total state $|\mbox{in}\rangle$ of the photon is given by the direct product of the photon states of each frequency as
$|\mbox{in}\rangle$ $=$ $\prod_{\omega}|\alpha(\omega)\rangle_{d}\otimes|0\rangle_{a}=\prod_{\omega}D_{d}[\alpha(\omega)]|0\rangle_{d}\otimes|0\rangle_{a}$ (100)
$=:$ $D_{d}|0\rangle_{d}\otimes|0\rangle_{a},$
$D_{d}$ $:=$ $\prod_{\omega}D_{d}(\alpha(\omega))$ (101)
$=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)d^{\dagger}(\omega)-\alpha^{*}(\omega)d(\omega)\right)\right].$
In the Heisenberg picture, the operator $\hat{d}$ is replaced as
$D_{d}^{\dagger}\hat{d}(\omega)D_{d}=\hat{d}(\omega)+\alpha(\omega),$ (102)
$D_{d}^{\dagger}\hat{d}^{\dagger}(\omega)D_{d}=\hat{d}^{\dagger}(\omega)+\alpha^{*}(\omega),$ (103)
through the displacement operator $D_{d}$ defined by Eq. (101). Then, the operator $\hat{D}(\omega)$ defined by Eq. (61) is transformed as
$D_{d}^{\dagger}\hat{D}(\omega)D_{d}=\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega),$ (104)
where
$\hat{D}_{c}(\omega):=\alpha(\omega)\Theta(\omega)+\alpha^{*}(-\omega)\Theta(-\omega),$ (105)
$\hat{D}_{v}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega).$ (106)
To evaluate the output signal expectation value and its fluctuations from the input-output relation (99) under the coherent state of the quadrature $\hat{D}(\omega)$, the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ is useful instead of $\hat{B}(\omega)$, because we treat operators in the Heisenberg picture. Furthermore, we regard that $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$ are small correction due to the radiation-pressure noise and gravitational-wave signals. Since $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$ also describe small fluctuations, we neglect the quadratic terms $\hat{D}_{v}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ and $\hat{A}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ in the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$.
Operating $D_{d}^{\dagger}$ and $D_{d}$ to Eq. (99), substituting Eqs. (104) into this $D_{d}^{\dagger}$-$D_{d}$-operated version of Eq. (99), and applying the above approximation, then, we obtain the input-output relation as
$\!\!\!\!\!\!D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ (107)
$=$ $i\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}_{c}(\omega)$
$+\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}$
$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}_{v}(\omega)+\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{A}(\omega)\right]$
$+\frac{i}{c}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\omega_{1}\hat{D}_{c}(\omega_{1})$
$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
$+i\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+i(\omega+\omega_{1})\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}\hat{D}_{c}(\omega_{1})$
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\omega-\omega_{1})D_{d}\right]$
$+\frac{i}{c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}\hat{D}_{c}(\omega_{1})$
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
[…]
The input-output relation (107) is the most general input-output relation within our consideration. The first term in the first line of Eq. (107) is the leakage of the classical carrier field due to the phase offset $\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})$. The terms in the second line are the vacuum fluctuations which correspond to the shot noise in the conventional input-output relations in Refs. [8]. The terms in the third- and fourth-lines are the response of the intermediate mirror motion which includes the radiation pressure noise through the motions of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$. The terms in the fifth-, sixth-lines are the response of the relative motion of the end mirrors and the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$. Finally, the term in the seventh- and eighth-lines are response of the intermediate mirror motion which includes the radiation pressure noise through the motion of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$. These come from the retarded effects which was explained after Eq. (94). The last line in Eq. (107) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2}$, $\hat{D}_{v}\hat{X}_{x}$, $\hat{D}_{v}\hat{X}_{y}$, $\hat{D}_{v}\hat{X}_{XITM}$, $\hat{D}_{v}\hat{X}_{YITM}$, $\hat{A}\hat{X}_{x}$, $\hat{A}\hat{X}_{y}$, $\hat{A}\hat{X}_{XITM}$, $\hat{A}\hat{X}_{YITM})$. The input-output relation (107) is one of the main results of this paper.
To conduct the actual evaluation of the input-output relation (107), we have to evaluate $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in some way. In the case of gravitational-wave detectors, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ are evaluated through the equations of motions for the end-mirrors as discussed in the next section.
IV.5.1 Monochromatic coherent amplitude
Before the evaluations of the mirror displacements, we consider the coherent state of the incident laser. In the conventional gravitational-wave detectors, the state of the optical beam from the light source is in the single-mode coherent state with the complex amplitude
$\alpha(\omega)=2\pi N\delta(\omega-\omega_{0}).$ (108)
We note that $\alpha(\omega)$ is real. The corresponding electric field with the amplitude (108) of $\alpha(\omega)$ is the continuous monochromatic carrier field with the frequency $\omega_{0}$. The electric field $\hat{E}_{d}$ is given through Eq. (7) as
$\hat{E}_{d}(t)=\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\hat{D}(\omega)e^{-i\omega t}.$ (109)
Then, we have obtain
$D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\left(\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega)\right)e^{-i\omega t}.$
Here, we note that Eq. (105) yields that
$\hat{D}_{c}(\omega)$ $:=$ $\alpha(\omega)\Theta(\omega)+\alpha^{*}(-\omega)\Theta(-\omega)$
$=$ $2\pi N\left\{\delta(\omega-\omega_{0})\Theta(\omega)+\delta(\omega+\omega_{0})\Theta(-\omega)\right\}.$
The quantum expectation value $I$ of the power of the electric field $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$ is given by
$I=N^{2}\hbar\omega_{0}(1+\cos(2\omega_{0}t))+\frac{\hbar}{2}\int_{0}^{+\infty}\frac{d\omega}{2\pi}\omega.$ (112)
The first term in Eq. (112) is the contribution from the classical carrier field $\alpha(\omega)$, and the second term comes from the vacuum fluctuations of the electric field operator $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$. The second term diverges as well-known. We neglect it when we estimate the classical power. However, if we take into account the cut-off in the frequency range due to the time bin and maximal observation time as noted in Sec. II.1, the second term is estimated as a negligible term. To evaluate the averaged classical power $I_{0}$ by neglecting the second term and by the time-average of Eq. (112) as
$I_{0}$ $=$ $\lim_{T\rightarrow\infty}\frac{1}{T}\int_{-T}^{T}dtN^{2}\hbar\omega_{0}(1+\cos(2\omega_{0}t))$ (113)
$=$ $N^{2}\hbar\omega_{0}.$
Then, we denote the factor $N$ in the function $\alpha(\omega)$ is given by
$N=\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}.$ (114)
IV.5.2 Final Input-output relation with mirrors’ motion
Now, we consider the input-output relation (107) with the monochromatic condition (IV.5.1). Furthermore, we consider the sideband picture $\omega=\omega_{0}\pm\Omega$ of the input-output relation under this monochromatic condition (IV.5.1). In this sideband picture, the terms include $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, or $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ appears in the expression of the input-output relation. Here, we neglect these terms because their behavior as
$\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}f(2\omega_{0}+\Omega)e^{-i(\Omega t}$
$=$ $\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}f(2\omega_{0}+\Omega)e^{-i(2\omega_{0}+\Omega-2\omega_{0})t}$
$=$ $e^{+2i\omega_{0}t}\int_{-\infty}^{+\infty}\frac{d\omega^{\prime}}{2\pi}f(\omega^{\prime})e^{-i\omega^{\prime}t},$
$\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}f(2\omega_{0}-\Omega)e^{-i(\Omega t}$
$=$ $\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}f(2\omega_{0}-\Omega)e^{+i(2\omega_{0}-\Omega+2\omega_{0})t}$
$=$ $e^{+2i\omega_{0}t}\int_{-\infty}^{+\infty}\frac{d\omega^{\prime}}{2\pi}f(\omega^{\prime})e^{+i\omega^{\prime}t}.$ (116)
These show that the terms of $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ have the behavior of the rapid oscillation with the frequency $2\omega_{0}$ in the time domain. Since we concentrate only on the behavior of the sideband frequency $\Omega$, these rapid oscillation terms are outside of the frequency range of interest. For these reasons, we ignore the terms $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ as an approximation [26]. Applying this approximation, the input-output relation (107) is given by
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{D}$ (117)
$=$ $N\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+2i\omega_{0}\tau}e^{+i\omega_{0}(\tau_{y}^{\prime}+\tau_{x}^{\prime})}i\sin(\omega_{0}(\tau_{y}^{\prime}-\tau_{x}^{\prime}))2\pi\delta(\Omega)$
$+\left[1-\sqrt{1-T}e^{-2i(\omega_{0}\pm\Omega)\tau}\right]\left[1-\sqrt{1-T}e^{+2i(\omega_{0}\pm\Omega)\tau}\right]^{-1}e^{+2i(\omega_{0}\pm\Omega)\tau}e^{+i(\omega_{0}\pm\Omega)(\tau_{y}^{\prime}+\tau_{x}^{\prime})}$
$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}_{v}(\omega_{0}\pm\Omega)+\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{A}(\omega_{0}\pm\Omega)\right]$
$+Nie^{+2i(\omega_{0}\pm\Omega)\tau}e^{+i(\omega_{0}\pm\Omega)(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i(\omega_{0}\pm\Omega)\tau}\right]\left[1-\sqrt{1-T}e^{+2i(\omega_{0}\pm\Omega)\tau}\right]^{-1}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\frac{\omega_{0}}{c}$
$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
$+Ni\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i(\omega_{0}\pm\Omega)\tau}\right]^{-1}\omega_{0}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+i(2\omega_{0}\pm\Omega)\tau}e^{+i(2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
$+Ni\frac{\omega_{0}}{c}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+2i\omega_{0}\tau}e^{+i(2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"}.$
To complete the evaluation of this input-output relation (117), we have to specify the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators. These are determined through the Heisenberg equation of motion in the next section.
V Equations for mirrors’ motions and their solutions
The purpose of this section is the evaluation of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators. To discuss this evaluation within quantum mechanics, we first discuss a quantum mechanical model of a forced harmonic oscillator in Sec. V.1. In Sec. V.2, we discuss the equations of mirrors’ motion in the proper reference frame [25] whose center is BS of the interferometer. In Sec. V.3, we evaluate the radiation pressure forces that act on each mirror. In Sec. V.4, we derived the explicit form of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators based on the arguments in Sec. V.1. The results of this section are used to evaluate the input-output relation of the Fabri-Pérot interferometer in Sec. VI.
V.1 Quantum Mechanical model for mirror motions and its solutions
In this subsection, we consider the quantum mechanical Hamiltonian $\hat{H}$ for a forced harmonic oscillator with the position operator $\hat{X}(t)$ and the momentum operator $\hat{P}(t)$ with external quantum force $\hat{F}(t)$:
$\hat{H}=\frac{\hat{P}^{2}}{2\mu}+\frac{1}{2}\mu m_{p}^{2}\hat{X}^{2}-\hat{F}(t)\hat{X}$. (118)
The position operator $\hat{X}$ corresponds to the mirror displacement operators $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. $\omega_{p}$ corresponds to the pendulum fundamental frequency of mirrors. The operators $\hat{X}(t)$ and $\hat{P}(t)$ satisfy the canonical commutation relation
$\left[\hat{X}(t),\hat{P}(t)\right]=i\hbar.$ (119)
We assume that
$\left[\hat{X},\hat{F}(t)\right]=\left[\hat{P},\hat{F}(t)\right]=0.$ (120)
From the Hamiltonian (118), Heisenberg’s equations of motion are given by
$\frac{d}{dt}\hat{X}$ $=$ $\frac{1}{i\hbar}\left[\hat{X},\hat{H}\right]=\frac{1}{\mu}\hat{P}$, (121)
$\frac{d}{dt}\hat{P}$ $=$ $\frac{1}{i\hbar}\left[\hat{P},\hat{H}\right]=-\mu\omega_{p}^{2}\hat{X}+\hat{F}(t).$ (122)
When the external force $\hat{F}(t)=0$, Eqs. (121) and (122) are the Heisenberg equations of motion for the simple harmonic oscillator. Then the solution to this equation is given by [27]
$\hat{X}_{\hat{F}=0}(t)$ $=$ $\hat{X}(0)\cos\omega_{p}t+\left[\frac{\hat{P}(0)}{\mu\omega_{p}}\right]\sin\omega_{p}t,$ (123)
$\hat{P}_{\hat{F}=0}(t)$ $=$ $-\mu\omega_{p}\hat{X}(0)\sin\omega_{p}t+\hat{P}(0)\cos\omega_{p}t.$ (124)
Substituting Eq. (121) into Eq. (122), the Heisenberg equation is given by
$\mu\frac{d^{2}}{dt^{2}}\hat{X}+\mu\omega_{p}^{2}\hat{X}-\hat{F}(t)=0.$ (125)
To solve this equation, we use the usual method to solve the second-order differential equation, i.e., we consider the solution
$\hat{X}(t)=\hat{{{\mathcal{A}}}}(t)\cos\omega_{p}t+\hat{{{\mathcal{B}}}}(t)\sin\omega_{p}t$ (126)
with the condition
$\left(\frac{d}{dt}\hat{{{\mathcal{A}}}}(t)\right)\cos\omega_{p}t+\left(\frac{d}{dt}\hat{{{\mathcal{B}}}}(t)\right)\sin\omega_{p}t=0.$ (127)
From Eq. (125) with Eqs. (126) and (127), we obtain
$-\left(\frac{d}{dt}\hat{{{\mathcal{A}}}}(t)\right)\sin\omega_{p}t+\left(\frac{d}{dt}\hat{{{\mathcal{B}}}}(t)\right)\cos\omega_{p}t=\frac{1}{\mu\omega_{p}}\hat{F}(t).$
From Eqs. (127) and (128), we obtain the solution operators $\hat{{{\mathcal{A}}}}(t)$ and $\hat{{{\mathcal{B}}}}(t)$. Then, we reach the solution $\hat{X}(t)$ to Eq. (125)
$\hat{X}(t)$ $=$ $\hat{X}(-\infty)\cos\omega_{p}t+\frac{1}{\mu\omega_{p}}\hat{P}(-\infty)\sin\omega_{p}t$
$+\frac{1}{\mu\omega_{p}}\int_{-\infty}^{+\infty}dt^{\prime}\Theta(t-t^{\prime})\hat{F}(t^{\prime})\sin(\omega_{p}(t-t^{\prime})).$
[…]
$\left[\hat{X}(-\infty),\hat{P}(-\infty)\right]=i\hbar.$ (130)
Here, we consider the Fourier transformation of the solution (LABEL:eq:Heisenberg-eqs-mirrors-3-sol). The Heaviside step function $\Theta(t-t^{\prime})$ is a distribution. Therefore, we have to be careful about the order of the limit when we consider the Fourier transformation of the Heaviside step function, which is given by
$\int_{-\infty}^{+\infty}dx\Theta(x)\phi(x)$ (131)
$=$ $\lim_{\epsilon\rightarrow 0}\int_{-\infty}^{+\infty}dx\int_{-\infty}^{+\infty}\frac{dk}{2\pi i}\frac{e^{+i(k-i\epsilon)x}}{k-i\epsilon}\phi(x)$
for any smooth function $\phi(x)$. Taking care of the order of the limit of the step function, through the Fourier transformations
$\hat{F}(t)$ $=:$ $\int_{-\infty}^{+\infty}\frac{d\omega_{F}}{2\pi}{{\mathcal{F}}}(\omega_{F})e^{-i\omega_{F}t},$ (132)
$\hat{X}(t)$ $=:$ $\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\hat{Z}(\omega)e^{-i\omega t},$ (133)
we obtain
$\hat{Z}(\omega)$ $=$ $\frac{1}{2}\left(\hat{X}(-\infty)-\frac{i\hat{P}(-\infty)}{\mu\omega_{p}}\right)2\pi\delta(\omega_{p}+\omega)$ (134)
$+\frac{1}{2}\left(\hat{X}(-\infty)+\frac{i\hat{P}(-\infty)}{\mu\omega_{p}}\right)2\pi\delta(\omega_{p}-\omega)$
$-\frac{1}{\mu(\omega^{2}-\omega_{p}^{2})}\hat{{{\mathcal{F}}}}(\omega).$
[…]
Through the canonical commutation relations (119) and (130), we can derive the consistency relation of the solution (LABEL:eq:Heisenberg-eqs-mirrors-3-sol) as the commutation relation of $\hat{{{\mathcal{F}}}}(\omega)/\mu$ as follows:
$\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\frac{\omega-\omega_{1}}{\left(\omega_{1}^{2}-\omega_{p}^{2}\right)\left((\omega-\omega_{1})^{2}-\omega_{p}^{2}\right)}$
$\quad\quad\times\left[\frac{1}{\mu}\hat{{{\mathcal{F}}}}(\omega_{1}),\frac{1}{\mu}\hat{{{\mathcal{F}}}}(\omega-\omega_{1})\right]=0.$ (135)
This consistency relation is checked in Appendix C through the explicit form of $\hat{{{\mathcal{F}}}}(\omega)/\mu$ obtained in Sec. V.3.
V.2 Equations of motion for each mirror
Now, we consider the equation of motion for EMs and ITMs to evaluate the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$, $D_{d}\hat{Z}_{com}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in the input-output relation (117). We assume the masses of ITMs are equal to $m_{ITM}$ and those of EMs are equal to $m_{EM}$. Since we apply the proper reference frame whose center is BS, the displacement of BS vanishes, which means that we have to control the external forces to BS vanish. Furthermore, we only consider the motions of mirrors are restricted so that XITM and XEM move only along the $x$-direction and YITM and YEM move only along the $y$-direction and $x$-direction and $y$-direction is orthogonal. In the proper reference frame [25], the effects of gravitational waves are represented by the components $R_{txtx}$, $R_{tyty}$ of the Riemann curvature. Moreover, we denote the radiation pressure force to mirrors, XITM, YITM, XEM, and YEM from the laser in the interferometer by $\hat{F}_{rpXITM}$, $\hat{F}_{rpYITM}$, $\hat{F}_{rpXEM}$, $\hat{F}_{rpYEM}$, respectively. Then, the equations of motion for $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{XEM}$, and $\hat{X}_{YEM}$, we obtain
$\frac{d^{2}}{dt^{2}}\hat{X}_{XITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XITM}+\frac{1}{m_{ITM}}\hat{F}_{rpXITM}$ (136)
$-R_{txtx}l_{x},$
$\frac{d^{2}}{dt^{2}}\hat{X}_{YITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{YITM}+\frac{1}{m_{ITM}}\hat{F}_{rpYITM}$ (137)
$-R_{tyty}l_{y},$
$\frac{d^{2}}{dt^{2}}\hat{X}_{XEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XEM}+\frac{1}{m_{EM}}\hat{F}_{rpXEM}$ (138)
$-R_{txtx}(L+l_{x}),$
$\frac{d^{2}}{dt^{2}}\hat{X}_{YEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{YEM}+\frac{1}{m_{EM}}\hat{F}_{rpYEM}$ (139)
$-R_{tyty}(L+l_{y}).$
Here, we assumed that $\omega_{p}$ is the fundamental frequency of the mirror pendulums of the vibration isolation system of XITM, YITM, XEM, and YEM. We also applied the approximation $L,l_{x},l_{y}$ $\gg$ $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{XEM}$, $\hat{X}_{YEM}$.
As the components of the Riemann tensor, we only consider the plus modes of gravitational waves associated with the $x$- and $y$-directions of the interferometer. In this case, the components of $R_{txtx}$ and $R_{tyty}$ the Riemann tensor are given by
$R_{txtx}=-\frac{1}{2}\frac{\partial^{2}}{\partial t^{2}}h_{+}=-R_{tyty},\quad h:=h_{+}.$ (140)
Furthermore, we ignore the gravitational-wave term in Eqs. (136) and (137) due to the approximation $L\gg l_{x},l_{y}$. Then, the equations (136)–(139) of motion are given by
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{XITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XITM}+\frac{1}{m_{ITM}}\hat{F}_{rpXITM},$ (141)
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{YITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{YITM}+\frac{1}{m_{ITM}}\hat{F}_{rpYITM},$ (142)
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{XEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XEM}+\frac{1}{m_{EM}}\hat{F}_{rpXEM}$ (143)
$+\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L),$
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{YEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{YEM}+\frac{1}{m_{EM}}\hat{F}_{rpYEM}$ (144)
$-\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L).$
From Eqs. (141)–(144), the equations for the relative displacements $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{YEM}-\hat{X}_{YITM}$ are given by
$\frac{d^{2}}{dt^{2}}\hat{X}_{x}+\omega_{p}^{2}\hat{X}_{x}$ $=$ $\frac{1}{m_{EM}}\hat{F}_{rpXEM}-\frac{1}{m_{ITM}}\hat{F}_{rpXITM}$ (145)
$+\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L),$
$\frac{d^{2}}{dt^{2}}\hat{X}_{y}+\omega_{p}^{2}\hat{X}_{y}$ $=$ $\frac{1}{m_{EM}}\hat{F}_{rpYEM}-\frac{1}{m_{ITM}}\hat{F}_{rpYITM}$ (146)
$-\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L).$
The equations (141), (142), (145), and (146) of motion for $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, and the relative motion $\hat{X}_{x}$ and $\hat{X}_{y}$ have the same form as the Heisenberg equation (125). Then, the solution to Eqs. (141), (142), (145), and (146) should have the form of Eq. (LABEL:eq:Heisenberg-eqs-mirrors-3-sol). Furthermore, their Fourier transformations are given in the form (134). The Fourier transformations of the operators $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ are expressed by Eqs. (65), (66), and (84). Moreover, we introduce the Fourier transformations of the radiation pressure forces by the Fourier transformations ${{\mathscr{F}}}_{rpXITM}(\omega)$, ${{\mathscr{F}}}_{rpYITM}(\omega)$, ${{\mathscr{F}}}_{rpx}(\omega)$, and ${{\mathscr{F}}}_{rpy}(\omega)$ of the radiation pressure forces, which are defined by
${{\mathscr{F}}}_{rpXITM}(\omega):=\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}\hat{F}_{rpXITM}(t),$
${{\mathscr{F}}}_{rpYITM}(\omega):=\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}\hat{F}_{rpYITM}(t),$
${{\mathscr{F}}}_{rpx}(\omega)$ $:=$ $\int_{-\infty}^{+\infty}dte^{+i\omega t}\left[\frac{1}{m_{EM}}\hat{F}_{rpXEM}(t)\right$.
$\quad\quad\quad\quad\quad\quad\quad\left.-\frac{1}{m_{ITM}}\hat{F}_{rpXITM}(t)\right],$
${{\mathscr{F}}}_{rpy}(\omega)$ $:=$ $\int_{-\infty}^{+\infty}dte^{+i\omega t}\left[\frac{1}{m_{EM}}\hat{F}_{rpYEM}(t)\right$.
$\quad\quad\quad\quad\quad\quad\quad\left.-\frac{1}{m_{ITM}}\hat{F}_{rpYITM}(t)\right].$
Since we express the mirrors’ displacement in the input-output relation (117) by their common modes and differential modes as $\hat{Z}_{com}(\pm\Omega)$, $\hat{Z}_{diff}(\pm\Omega)$, $\hat{Z}_{comITM}(\pm\Omega)$, and $\hat{Z}_{diffITM}(\pm\Omega)$, we use the definitions (95)–(98) for these variables.
To describe the Fourier transformations of Eqs. (141), (142), (145), and (146), we have to introduce the Fourier transformation $H(\omega)$ of the gravitational wave signal $h(t)$ as
$h(t,z)=:\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}e^{-i\omega_{2}t}H(\omega,z),$ (151)
where $z$ is the propagation direction of the laser in the interferometer.
[…]
$\hat{Z}_{comITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpYITM}(\omega)+{{\mathscr{F}}}_{rpXITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
$\hat{Z}_{difITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpXITM}(\omega)-{{\mathscr{F}}}_{rpYITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
$\hat{Z}_{com}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpy}(\omega)+{{\mathscr{F}}}_{rpx}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$ (154)
$\hat{Z}_{dif}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpx}(\omega)-{{\mathscr{F}}}_{rpy}(\omega)}{2(\omega^{2}-\omega_{p}^{2})}$ (155)
$+\frac{\omega^{2}}{2(\omega^{2}-\omega_{p}^{2})}H(\omega,L)L.$
In the input-output relation (117), $\hat{Z}_{dif}(\omega)$, $\hat{Z}_{com}(\omega)$, $\hat{Z}_{comITM}(\omega)$, and $\hat{Z}_{difITM}(\omega)$ do not appear in this expression, but $D_{d}^{\dagger}\hat{Z}_{dif}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{difITM}(\omega)D_{d}$ appear in Eq. (117). Through Eqs. (154) and (155), $\hat{Z}_{dif}(\omega)$ and $\hat{Z}_{com}(\omega)$ are linearly related to ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$. Furthermore, Eqs. (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx) and (LABEL:eq:Eq-of-motion-XYITM-dif-Fourier-approx) yields that $\hat{Z}_{comITM}(\omega)$ and $\hat{Z}_{difITM}(\omega)$ linearly depends on ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$. Therefore, we have to evaluate $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$ to the direct evaluation of the input-output relation (117). These evaluations are discussed in the next subsection.
V.3 Evaluation of Radiation pressure forces
Now, we evaluate the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$. The details of the evaluation of these radiation pressure forces are given in Appendix B. In this section, we only explain the outline of Appendix B.
As described in Ref. [8], the radiation pressure forces on the mirrors are determined by the power operator ${\cal A}\hat{E}^{2}(t)/4\pi$, where $\hat{E}(t)$ is the electric field operator that touches the mirror. Since we assume the perfect reflection at EMs, the radiation pressure force on EMs are given by
$2\times\frac{{\cal A}}{4\pi}\hat{E}_{j_{x,y}}^{2}(t)$ (156)
as depicted in Fig. 2. From the propagation effect from the ITMs to EMs, the radiation pressure forces to XEM and YEM are given by
$\hat{F}_{rpXEM}(t)$ $=$ $\frac{2{\cal A}}{4\pi}\hat{E}_{g_{x}}^{2}\left[t-\tau-\frac{1}{c}\hat{X}_{x}(t-\tau)\right],$ (157)
$\hat{F}_{rpYEM}(t)$ $=$ $\frac{2{\cal A}}{4\pi}\hat{E}_{g_{y}}^{2}\left[t-\tau-\frac{1}{c}\hat{X}_{y}(t-\tau)\right],$ (158)
respectively. We evaluate the radiation pressure forces (157) and (158) through the propagation effects (69) and (70) from BS to ITMs for the electric field operator, the junction condition (74) for the electric field operator at the ITMs, the propagation effects (78) and (79) between ITMs and EMs, and the junction conditions (55) and (56) for the electric field operator at BS, and their Fourier transformations. Then, we obtain the expression of the Fourier transformations of the radiation pressure forces $\hat{F}_{rpXEM}(t)$ and $\hat{F}_{rpYEM}(t)$ as Eqs. (326) and (327), respectively.
On the other hand, the radiation pressure forces on ITMs are more complicated than those on EMs. As depicted in Fig. 2, four electric fields contribute to the radiation pressure forces to ITMs, which are $\hat{E}_{f_{x}}(t)$, $\hat{E}_{f_{x}^{\prime}}(t)$, $\hat{E}_{g_{x}}(t)$, and $\hat{E}_{f_{x}^{\prime}}(t)$. Taking into account the directions of the forces from these electric fields to ITMs, the radiation pressure force to XITM is given by
$\hat{F}_{rpXITM}(t)$ $=$ $\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{x}}(t)\right)^{2}+\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{x}^{\prime}}(t)\right)^{2}$
$-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{x}}(t)\right)^{2}-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{x}^{\prime}}(t)\right)^{2}.$
Similarly, the radiation pressure force on the YITM is given by
$\hat{F}_{rpYITM}(t)$ $=$ $\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{y}}(t)\right)^{2}+\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{y}^{\prime}}(t)\right)^{2}$
$-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{y}}(t)\right)^{2}-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{y}^{\prime}}(t)\right)^{2}.$
In Appendix B, we evaluate these radiation pressure forces (157)–(LABEL:eq:total-radiation-pressure-force-YITM) within the first-order of the displacements $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$ through the Fourier transformations (65), (66), and (84). In this evaluation, we use the junction conditions (55) and (56) for the electric field operators at BS, the propagation effects (69) and (70) between BS and ITMs, the junction conditions (74) at ITMs, the propagation effects (78) and (79), and their Fourier transformations. Then, we reach the expression of the Fourier transformations of the radiation pressure forces $\hat{F}_{rpXITM}(t)$ and $\hat{F}_{rpYITM}(t)$ as Eqs. (329) and (331), respectively.
To consider the situation where the state for the incident electric field from the light source in the coherent state with the complex amplitude $\alpha(\omega)$, we apply the displacement operator $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right, which are defined by Eq. (101), to the obtained radiation pressure forces (326), (327), (329), and (331). Using (104), we separate the incident electric field quadrature $\hat{D}(\omega)$ from the light source into its classical part and the part of quantum fluctuations. Then, we evaluate the classical part, the linear-order contributions of the quantum fluctuations from $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$, and the linear-order contributions from the mirror displacements $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. After this evaluation, we consider the situation where the incident electric field from the light source is in the monochromatic coherent state with the central frequency $\omega_{0}$ through the choice of the complex amplitude $\alpha(\omega)$ for the coherent state by Eq. (108).
Thus, we reach the expressions of the Fourier transformation of the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ as Eqs. (338), (339), (340), and (341), respectively.
Here, we consider the tuning condition of the Fabri-Pérot cavity by
$\omega_{0}\tau=2n\pi,\quad n\in{{\mathbb{N}}}.$ (161)
Due to this tuning condition, the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ given by Eqs. (338), (339), (340), and (341), respectively, are expressed as
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (162)
$=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
$-2N\sqrt{1-T}\frac{\hbar}{m_{ITM}c}e^{+i\Omega\tau}e^{+i\Omega\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)$
$\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)\right],$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (163)
$=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
$-2N\sqrt{1-T}\frac{\hbar}{m_{ITM}c}e^{+i\Omega\tau_{y}^{\prime}}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)$
$\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)\right],$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (164)
$=$ $+\frac{N^{2}\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[1-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
$+\frac{N\hbar}{c}e^{+i\Omega\tau_{x}^{\prime}}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\cos(\Omega\tau)\left[1-\sqrt{1-T}\right]\right]$
$\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)\right],$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$ (165)
$=$ $+N^{2}\frac{\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[1-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
$+N\frac{\hbar}{c}e^{+i\Omega\tau}e^{+i\Omega\tau_{y}^{\prime}}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\cos(\Omega\tau)\left[1-\sqrt{1-T}\right]\right]$
$\quad\quad\times\left[\sqrt{|(\Omega-\omega_{0})\omega_{0}|}\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|(\Omega+\omega_{0})\omega_{0}|}\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)\right],$
respectively. Here, we note that Eqs. (162)–(165) includes the terms which proportional to $\delta(\Omega)$. These terms represent the classical constant force in time. Due to the existence of these terms, we have to develop further discussion about the tuning condition (161) in Sec. VIII. Before we reach the discussions in Sec. VIII, we keep the tuning condition (161) to observe the results caused by these classical constant forces.
V.4 Solutions to the Heisenberg equations for mirrors’ motion
Now, we can write down the solution to the Heisenberg equations (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx)–(155). Substituting the expression of the radiation pressure forces (162)–(165) into Eqs. (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx)–(155), we can obtain the explicit solution to the Heisenberg equations as follows:
$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ (166)
$=$ $+\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$
$-\frac{N\hbar}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}e^{+i\Omega\left(\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}\right)}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]$
$\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\cos\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega-\omega_{0})+i\sin\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{A}(\Omega-\omega_{0})\right)\right$.
$\quad\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\cos\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})+i\sin\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right)\right],$
$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ (167)
$=$ $-\frac{N\hbar}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}e^{+i\Omega\left(\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}\right)}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]$
$\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega-\omega_{0})-\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega-\omega_{0})\right]\right$.
$\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})-\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right]\right]$
$+\frac{\Omega^{2}}{2(\Omega^{2}-\omega_{p}^{2})}H(\Omega,L+l)L,$
$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ (168)
$=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
$+\frac{2N\sqrt{1-T}\hbar}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\cos(\Omega\tau)\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}e^{+i\Omega\left(\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}\right)}$
$\quad\times\left[+\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega-\omega_{0})-i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega-\omega_{0})\right]\right$.
$\quad\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})-i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right]\right],$
$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ (169)
$=$ $-\frac{2N\sqrt{1-T}\hbar}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)e^{+i\Omega\left(\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}\right)}$
$\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[-i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega-\omega_{0})+\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega-\omega_{0})\right]\right$.
$\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[-i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})+\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right]\right].$
Here, we evaluate the order of the phase $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2$. In Sec. VI, we choose the phase offset of the interferometer so that $\omega_{0}(\tau_{y}^{\prime}-\tau_{x}^{\prime})$ $\sim$ $O(1)$. Then, the order of $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2$ is given by
$\Omega\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}$ $\sim$ $10^{-10}\left(\frac{\Omega}{2\pi\times 10^{4}\mbox{Hz}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right).$
This indicates that we may regard that $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2$ is zero. Furthermore, we choose $(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2$ so that
$\Omega\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}$ $\sim$ $2\times 10^{-3}\left(\frac{\Omega}{2\pi\times 10^{4}\mbox{Hz}}\right)\left(\frac{(l_{x}+l_{y})/2}{10\mbox{m}}\right).$
We may also regard that $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2$ is zero.
Then, we apply the approximation $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2\sim 0$ and $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2\sim 0$. In addition, we apply the approximation $\Omega\ll\omega_{0}$. Through these approximations, (166)–(169) to the Heisenberg equations for mirrors’ motion [28]
$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ $=$ $\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$ (172)
$-\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]\left[\hat{D}_{v}(\Omega-\omega_{0})+\hat{D}_{v}(\Omega+\omega_{0})\right],$
$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ $=$ $\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$ (173)
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right]$
$+\frac{\Omega^{2}}{2(\Omega^{2}-\omega_{p}^{2})}H(\Omega,L+l)L,$
$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
$+\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\cos(\Omega\tau)\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\hat{D}_{v}(\Omega-\omega_{0})+\hat{D}_{v}(\Omega+\omega_{0})\right],$
$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ $=$ $-\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right].$ (175)
We have to note that the classical constant terms in Eq. (172) and (174) diverge if we choose $\omega_{p}=0$, i.e., if we assume the mirrors are in completely free motion. This is the reason why we introduced the nonvanishing pendulum fundamental frequency $\omega_{p}$ in each equation of motion for mirrors. This introduction of the nonvanishing pendulum fundamental frequency $\omega_{p}$ is a natural consequence from the viewpoint of the actual ground-based gravitational-wave detectors. From a mathematical point of view, this introduction of $\omega_{p}$ is a kind of regularization of the divergence. However, we can regard the introduction of this regularization as reasonable for the actual ground-based gravitational-wave detectors. In this sense, we have to emphasize that this introduction $\omega_{p}$ has a physical meaning more than a mere mathematical regularization.
[…]
VI Final input-output relation for a Fabri-Pérot Graviational-Wave Detector
Now, we return to the evaluation of the input-output relation (117). First, we introduce the offset phase $\theta$ by
$\theta:=\omega_{0}(\tau_{y}^{\prime}-\tau_{x}^{\prime}).$ (176)
We also apply the approximation $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2\sim 0$ and $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2\sim 0$ as estimated in Eqs. (LABEL:eq:Omegatauy-taux-half) and (LABEL:eq:Omegatauy+taux-half). Furthermore, we consider the tuning condition of the Fabri-Pérot cavity by Eq. (161). In addition, we employ the tuning condition of the lengths between the BS and ITMs by
$\omega_{0}\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}=2m\pi,\quad m\in{{\mathbb{N}}}.$ (177)
Moreover, we apply the approximation $\Omega\ll\omega_{0}$.
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (178)
$=$ $+iN\sin\theta 2\pi\delta(\Omega)$
$+\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}e^{\pm 2i\Omega\tau}\left[i\sin\theta\hat{D}_{v}(\omega_{0}\pm\Omega)+\cos\theta\hat{A}(\omega_{0}\pm\Omega)\right]$
$+i\frac{N\omega_{0}}{c}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
$-i\frac{2NT\omega_{0}}{c}e^{\pm i\Omega\tau}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}+\cos\theta D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
$+\frac{iN\omega_{0}}{c}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right].$
Substituting the radiation pressure forces (172)–(175) into Eq. (178), and applying the approximation $L\gg l_{x,y}$, we reached the input-output relation
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (179)
$=$ $+iN\sin\theta\left[1+i\frac{2N^{2}\hbar\omega_{0}^{2}}{c^{2}\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-4}\left[\frac{T^{2}}{m_{EM}}+\frac{4(1-T)\left[1-\sqrt{1-T}\right]^{2}}{m_{ITM}}\right]\right]2\pi\delta(\Omega)$
$+\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}e^{\pm 2i\Omega\tau}\left[i\sin\theta\hat{D}_{v}(\omega_{0}\pm\Omega)+\cos\theta\hat{A}(\omega_{0}\pm\Omega)\right]$
$-2i\frac{N^{2}\hbar\omega_{0}^{2}}{c^{2}\Omega^{2}}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-2}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-2}$
$\quad\quad\quad\times\left[\frac{T^{2}}{m_{EM}}+2\sqrt{1-T}\cos(\Omega\tau)\left[1-\sqrt{1-T}\right]\left[T-\left[1-\sqrt{1-T}\right]^{2}\cos(\Omega\tau)\right]\frac{1}{m_{ITM}}\right]$
$\quad\quad\quad\times\left[i\sin\theta\left(\hat{D}_{v}(-(\omega_{0}\mp\Omega))+\hat{D}_{v}(\omega_{0}\pm\Omega)\right)+\cos\theta\left(\hat{A}(-(\omega_{0}\mp\Omega))+\hat{A}(\omega_{0}\pm\Omega)\right)\right]$
$-ie^{\pm i\Omega\tau}\cos\theta\frac{NTL\omega_{0}}{c}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}H(\pm\Omega,L).$
Here, from the definition of the quadrature (6) and the situation $\omega_{0}\gg\Omega$, we note that the operator $\hat{B}(\omega_{0}\pm\Omega)$, $\hat{D}_{v}(\omega_{0}\pm\Omega)$, $\hat{D}_{v}(-(\omega_{0}\mp\Omega))$, $\hat{A}(\omega_{0}\pm\Omega)$, and $\hat{A}(-(\omega_{0}\mp\Omega))$ are given by
$\hat{B}(\omega_{0}\pm\Omega)$ $=$ $\hat{b}(\omega_{0}\pm\Omega)=:\hat{b}_{\pm}(\Omega),$ (180)
$\hat{D}_{v}(\omega_{0}\pm\Omega)$ $=$ $\hat{d}(\omega_{0}\pm\Omega)=:\hat{d}_{\pm}(\Omega),$ (181)
$\hat{D}_{v}(-(\omega_{0}\mp\Omega))$ $=$ $\hat{d}^{\dagger}(\omega_{0}\mp\Omega)=:\hat{d}_{\mp}^{\dagger}(\Omega),$ (182)
$\hat{A}(\omega_{0}\mp\Omega)$ $=$ $\hat{a}(\omega_{0}\mp\Omega)=:\hat{a}_{\pm}(\Omega),$ (183)
$\hat{A}(-(\omega_{0}\mp\Omega))$ $=$ $\hat{a}^{\dagger}(\omega_{0}\mp\Omega)=:\hat{a}_{\mp}^{\dagger}(\Omega).$ (184)
Furthermore, using Eq. (114), we define the variable $\mathfrak{R}$}}, the phase $\beta$, $\kappa$, and $h_{SQL}$ by
${{\mathfrak{R}}}$ $:=$ $\frac{2I_{0}\omega_{0}T^{2}}{c^{2}\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-4}$
$\quad\times\left[\frac{1}{m_{EM}}+\frac{1}{m_{ITM}}\frac{4(1-T)\left[1-\sqrt{1-T}\right]^{2}}{T^{2}}\right],$
$e^{\pm 2i\beta}$ $:=$ $\frac{e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]}{\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]}.$ (186)
$\kappa$ $:=$ $\frac{4I_{0}T^{2}\omega_{0}}{c^{2}\Omega^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[2-T-2\sqrt{1-T}\cos(2\Omega\tau)\right]^{-1}$ (187)
$\quad\quad\quad\times\left[\frac{1}{m_{EM}}+\frac{2\sqrt{1-T}\left[1-\sqrt{1-T}\right]\cos(\Omega\tau)}{T}\left[1-\frac{\left[1-\sqrt{1-T}\right]^{2}}{T}\cos(\Omega\tau)\right]\frac{1}{m_{ITM}}\right],$
$h_{SQL}$ $:=$ $\sqrt{\frac{4\hbar}{\Omega^{2}L^{2}}}\left[\frac{1}{m_{EM}}+\frac{2\sqrt{1-T}\left[1-\sqrt{1-T}\right]\cos(\Omega\tau)}{T}\left[1-\frac{\left[1-\sqrt{1-T}\right]^{2}}{T}\cos(\Omega\tau)\right]\frac{1}{m_{ITM}}\right]^{1/2}.$ (188)
Through the quadratures (180)–(184), variables $\mathfrak{R}$}}, $\beta$, $\kappa$, and $h_{SQL}$, the input-output relation (179) is given by
$D_{d}^{\dagger}\hat{b}_{\pm}(\Omega)D_{d}$ $=$ $+i\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\sin\theta\left[1+i{{\mathfrak{R}}}\right]2\pi\delta(\Omega)$ (189)
$+e^{\pm 2i\beta}\left[i\sin\theta\hat{d}_{\pm}(\Omega)+\cos\theta\hat{a}_{\pm}(\Omega)\right]$
$+e^{\pm 2i\beta}\frac{\kappa}{2}\left[\sin\theta\left(\hat{d}^{\dagger}_{\mp}(\Omega)+\hat{d}_{\pm}(\Omega)\right)-i\cos\theta\left(\hat{a}^{\dagger}_{\mp}(\Omega)+\hat{a}_{\pm}(\Omega)\right)\right]$
$-ie^{\pm i\beta}\sqrt{\kappa}\cos\theta\frac{H(\pm\Omega,L)}{h_{SQL}}.$
Here, we consider the representation of the input-output relation (189) in terms of the two-photon formulation. In the two-photon formulation, we introduce the amplitude operators $\hat{a}_{1}$, $\hat{b}_{1}$, $\hat{d}_{1}$ and the phase operators $\hat{a}_{2}$, $\hat{b}_{2}$, $\hat{d}_{2}$ as follows:
$\hat{a}_{1}$ $:=$ $\frac{1}{\sqrt{2}}(\hat{a}_{+}+\hat{a}_{-}^{\dagger}),\quad\hat{a}_{2}:=\frac{1}{\sqrt{2}i}(\hat{a}_{+}-\hat{a}_{-}^{\dagger}),$ (190)
$\hat{b}_{1}$ $:=$ $\frac{1}{\sqrt{2}}(\hat{b}_{+}+\hat{b}_{-}^{\dagger}),\quad\hat{b}_{2}:=\frac{1}{\sqrt{2}i}(\hat{b}_{+}-\hat{b}_{-}^{\dagger}),$ (191)
$\hat{d}_{1}$ $:=$ $\frac{1}{\sqrt{2}}(\hat{d}_{+}+\hat{d}_{-}^{\dagger}),\quad\hat{d}_{2}:=\frac{1}{\sqrt{2}i}(\hat{d}_{+}-\hat{d}_{-}^{\dagger}).$ (192)
From these definitions, we obtain
$D_{d}^{\dagger}\hat{b}_{1}(\Omega)D_{d}$ $=$ $-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}{{\mathfrak{R}}}\sin\theta 2\pi\delta(\Omega)$ (193)
$+e^{+2i\beta}\left[-\sin\theta\hat{d}_{2}(\Omega)+\cos\theta\hat{a}_{1}(\Omega)\right]$
$+e^{+2i\beta}\kappa\sin\theta\hat{d}_{1}(\Omega).$
$D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}$ $=$ $\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)$ (194)
$+e^{+2i\beta}\left[\sin\theta\hat{d}_{1}(\Omega)+\cos\theta\hat{a}_{2}(\Omega)\right]$
$-e^{+2i\beta}\cos\theta\kappa\hat{a}_{1}(\Omega)$
$-e^{+i\beta}\sqrt{2\kappa}\cos\theta\frac{H(\Omega,L)}{h_{SQL}}.$
In Ref. [8], $D_{d}^{\dagger}\hat{b}_{2}D_{d}$ is regarded as the signal operator in gravitational-wave detectors. The first line of Eq. (194) is the contribution from the classical coherent light. The second line of Eq. (194) corresponds to the shot noise due to the errors of the photon number estimation in the coherent state of the laser. The third line of Eq. (194) is the contribution from the radiation pressure noise. Then, the final line of Eq. (194) corresponds to the gravitational-wave signal cooperating with the response function $e^{+i\beta}\sqrt{2\kappa}\cos\theta/h_{SQL}$.
In Ref. [8], the first line of the classical coherent light in Eqs. (193) and (194) are regarded as trivial, and are neglected. Furthermore, in Ref. [8], the quantum noise against the signal $H(\Omega,L)$ of gravitational waves is determined by the quantum operator $\hat{h}_{n}$, which is given by
$\hat{h}_{n}$ $:=$ $e^{-i\beta}\frac{h_{SQL}}{\sqrt{2\kappa}\cos\theta}$ (195)
$\quad\times\left[\left(D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)\right)\right$.
$\quad\quad\quad\left.+e^{+i\beta}\sqrt{2\kappa}\cos\theta\frac{H(\Omega,L)}{h_{SQL}}\right]$
$=$ $e^{-i\beta}\frac{h_{SQL}}{\sqrt{2\kappa}\cos\theta}$
$\quad\times\left[e^{+2i\beta}\left[\sin\theta\hat{d}_{1}(\Omega)+\cos\theta\hat{a}_{2}(\Omega)\right]\right$.
$\quad\quad\quad\left.-e^{+2i\beta}\cos\theta\kappa\hat{a}_{1}(\Omega)\right].$
Moreover, in Ref. [8], the quantum noise spectral density is estimated by
$\frac{1}{2}2\pi\delta(\Omega-\Omega^{\prime})S_{(K)}(\Omega)$ (196)
$:=$ $\frac{1}{2}\langle\mbox{in}|\hat{h}_{n}(\Omega)\hat{h}_{n}^{\dagger}(\Omega^{\prime})+\hat{h}_{n}^{\dagger}(\Omega^{\prime})\hat{h}_{n}(\Omega)|\mbox{in}\rangle.$
without any reason except for the citation of Refs. [30, 31]. Incidentally, in Refs. [30, 31], there is no description that the noise spectral density $S_{(K)}(\Omega)$ defined by Eq. (196) is the actual quantum noise in measurements of gravitational waves. They only suggested that $S_{(K)}(\Omega)$ is an estimation measure of the amplitude of the quantum noises in the lasers of interferometers.
Besides the problem whether $S_{(K)}(\Omega)$ defined by Eq. (196) is an physically reasonable noise estimation or not, we can calculate $S_{(K)}(\Omega)$ through the definitions (196) of $S_{(K)}(\Omega)$ and the definition (195) of the noise operator $\hat{h}_{n}$. As a result of this calculation, we obtain the result
$S_{(K)}(\Omega)$ $=$ $\frac{h_{SQL}^{2}(\Omega)}{2\cos^{2}\theta}\left(\frac{1}{\kappa(\Omega)}+\cos^{2}\theta\kappa(\Omega)\right).$ (197)
It is a famous fact that $1/\kappa$ in the bracket of the right-hand side of Eq. (197) is the contribution from the shot noise of the coherent state of the laser and $\kappa$ in the bracket of the right-hand side of Eq. (197) is the contribution from the radiation pressure noise. It is also a famous fact that $\kappa$ depends on the frequency $\Omega$, which corresponds to the frequency of the gravitational-wave signal, and the laser power $I_{0}$ as shown in Eq. (187). Furthermore, we can also show that $S_{(K)}$ is bounded below as
$S_{(K)}(\Omega)\geq h_{SQL}^{2}.$ (198)
This is called “standard quantum limit” in the context of the measurement of gravitational-waves [8, 9]. Here, we have to mention that these variables $\beta$, $\kappa$, and $h_{SQL}$ defined by Eqs. (186), (187), and (188), respectively, are defined so that Kimble’s noise spectral density $S_{(K)}$ is realized in the form of Eq. (197).
One of the main problems that we discuss in this paper is whether the Kimble noise spectral density $S_{(K)}(\Omega)$ is regarded as the noise in the measurement process of gravitational waves in some sense, or not.
Before discussing this main problem within this paper, we compare the variables $\beta$, $\kappa$, and $h_{SQL}$ defined by Eqs. (186), (187), and (188), with those in Ref. [8]. First, we evaluate the phase $\beta$ and obtain the result
$\beta$ $=$ $-\frac{1}{2}\arctan\left(2\frac{Tc}{4\Omega L}\right)$ (199)
$+O(T^{2},(2\Omega\tau)^{3},T(2\Omega\tau))$
and we may regard
$\beta=-\frac{Tc}{4\Omega L}+O(T^{2},(2\Omega\tau)^{3},T(2\Omega\tau)).$ (200)
This coincides with the Kimble’s $\beta$ in Ref. [8]. In this paper, we denote this $\beta$ as
$\beta^{(K)}:=-\frac{Tc}{4\Omega L}=:-\frac{\gamma}{\Omega},$ (201)
where we defined the resonant frequency $\gamma$ by
$\gamma:=\frac{Tc}{4L}.$ (202)
On the other hand, $\kappa$ defined by Eq. (187) is evaluated as
$\kappa$ $=$ $4\frac{I_{0}\omega_{0}}{L^{2}\Omega^{2}(\Omega^{2}+\gamma^{2})}\left[\frac{1}{m_{EM}}+\frac{1}{m_{ITM}}\right]$ (203)
$+O(T^{3},T(\Omega\tau)^{2},(\Omega\tau)^{4}).$
When $m_{EM}=m_{ITM}=:m$, we obtain
$\kappa$ $=$ $\frac{2(I_{0}/I_{SQL})\gamma^{4}}{\Omega^{2}(\Omega^{2}+\gamma^{2})}+O(T^{3},T(\Omega\tau)^{2},(\Omega\tau)^{4}),$ (204)
where
$I_{SQL}:=\frac{mL^{2}\gamma^{4}}{4\omega_{0}}.$ (205)
Eq. (204) is a realization of Kimble’s $\kappa$ in Ref. [8]. In this paper, we denote Kimble’s $\kappa$ as
$\kappa^{(K)}$ $:=$ $\frac{2(I_{0}/I_{SQL})\gamma^{4}}{\Omega^{2}(\Omega^{2}+\gamma^{2})}$ (206)
$=$ $\frac{4I_{0}\omega_{0}}{mc^{2}\Omega^{2}}\frac{32}{\left(16(\Omega\tau)^{2}+T^{2}\right)}.$
Furthermore, we also define $\kappa^{(EQ)}$ from $\kappa$ defined by Eq. (187) as
$\kappa^{(EQ)}:=\left.\kappa\right|_{m=m_{EM}=m_{ITM}}.$ (207)
The ratio $\kappa^{(EQ)}/\kappa^{(K)}$ are depicted in Fig. 3. In Fig. 3, we choose the transmissivity $T=10^{-2}$ and $\tau=L/c=10^{-5}$s. This figure indicates that $\kappa^{EQ}$ is 10% greater than $\kappa^{(K)}$ at the high frequency region ($\sim 20$kHz) and 1% smaller than $\kappa^{(K)}$ at low frequency region ($\sim 10$Hz).
Figure 3: The ratio $\kappa^{(EQ)}/\kappa^{(K)}$ given by Eqs. (206) and (207). We choose the transmissivity $T=10^{-2}$ and $\tau=L/c=10^{-5}$s. This figure indicates that $\kappa^{EQ}$ is 10% greater than $\kappa^{(K)}$ at the high frequency region ($\sim 20$kHz) and 1% smaller than $\kappa^{(K)}$ at low frequency region ($\sim 10$Hz).
Next, we consider $h_{SQL}$ defined by Eq. (188), which is estimated as
$h_{SQL}$ $=$ $\sqrt{\frac{4\hbar}{\Omega^{2}L^{2}}}\left(\frac{1}{m_{EM}}+\frac{1}{m_{ITM}}\right)^{1/2}$ (208)
$+O(T,(\Omega\tau)^{2}).$
Here again, we consider the case $m:=m_{EM}=m_{ITM}$ and we define
$h_{SQL}^{(K)}$ $:=$ $\sqrt{\frac{8\hbar}{m\Omega^{2}L^{2}}},$ (209)
$h_{SQL}^{(EQ)}$ $:=$ $\left.h_{SQL}\right|_{m=m_{EM}=m_{ITM}}.$ (210)
$h_{SQL}^{(K)}$ is the $\kappa$ in Ref. [8]. Then, Eq. (208) yields that the leading term of $h_{SQL}$ defined by Eq. (188) coincides with the $h_{SQL}$ in Ref. [8]. We also show the ratio $h_{SQL}^{(EQ)}/h_{SQL}^{(K)}$ in Fig. 4. The difference between $h_{SQL}^{(EQ)}$ and $h_{SQL}^{(K)}$ is $\sim 15\%$ at the high frequency range $\sim 20$kHz. This is due to the $\cos(\Omega\tau)$-dependence in Eq. (188) which is introduced by the phase difference in the radiation pressure forces affecting the EM and ITM. Essentially, this difference was introduced by the motion of ITMs in the input-output relation, which are ignored in Ref. [8].
Figure 4: The ratio $h_{SQL}^{(EQ)}/h_{SQL}^{(K)}$ is depicted. $h_{SQL}^{(K)}$ and $h_{SQL}^{(EQ)}$ are defined in Eqs. (209) and (209), respectively. The difference between $h_{SQL}^{(EQ)}$ and $h_{SQL}^{(K)}$ is $\sim 15\%$ at the high frequency range $\sim 20$kHz. This is due to the $\cos(\Omega\tau)$-dependence in Eq. (188).
Although $\kappa$ and $h_{SQL}$ were defined so that the noise spectral density $S_{(K)}$ is given by Eq. (197) which leads to the inequality (198), the input-output relation (179) is more accurate than that in Ref. [8]. The $10\%\sim 20\%$ difference from those in Ref. [8] can be seen in the high-frequency region.
Finally, we consider the difference $S_{(K)}(\Omega)$ defined by Eq. (197) from the noise spectral density in Ref. [8]. To discuss this difference, we again consider the case $m:=m_{EM}=m_{ITM}$ and we define $S_{(K)}^{(EQ))}(\Omega)$ by
$S_{(K)}^{(EQ)}(\Omega):=\left.S_{(K)}(\Omega)\right|_{m:=m_{EM}=m_{ITM},\theta=0}.$ (211)
We also define the other noise spectral density $S_{(K)}^{(K)}(\Omega)$ by
$S_{(K)}^{(K)}(\Omega):=\frac{(h_{SQL}^{(K)})^{2}}{2}\left(\frac{1}{\kappa^{(K)}}+\kappa^{(K)}\right),$ (212)
where $\kappa^{(K)}$ and $h_{SQL}^{(K)}$ are defined Eqs. (206) and (209), respectively. This $S_{(K)}^{(K)}(\Omega)$ coincides with the noise spectral density for the Fabri-Pérot interferometer in Ref. [8].
The square root of the ratio $\sqrt{S_{(K)}^{(EQ)}(\Omega)/S_{(K)}^{(K)}(\Omega)}$ of these noise spectral densities is depicted in Fig. 5. Fig. 5 shows that the noise spectral density $\sqrt{S_{(K)}^{(EQ)}}$ is 20% smaller than the Kimble’s noise spectral density $\sqrt{S_{(K)}^{(K)}}$ at the high frequency range $\sim 20$kHz. We also observe a small power dependence within the range of 1Hz to 100 Hz. This difference in the range 1Hz to 100Hz is roughly $1\%$.
Figure 5: The square root of the ratio $\sqrt{S_{(K)}^{(EQ)}/S_{(K)}^{(K)}}$ defined by Eqs. (211) and (212) in the range 1Hz to 20kHz. $\sqrt{S_{(K)}^{(EQ)}}$ is 20% smaller than $\sqrt{S_{(K)}^{(K)}}$ at the high frequency range $\sim 20$kHz. We also observe a small power dependence in the range 1Hz to 100 Hz. This difference in the range 1Hz to 100Hz is roughly $1\%$.
[…]
In the next section, we derive the quantum noise spectral density from the quantum fluctuations of photons at the photo-detector through a DC readout scheme, which is discussed in Sec. III.
VII DC readout scheme for a Fabri-Pérot Graviational-Wave Detector
Here, we apply the general arguments on the DC readout scheme for a Fabri-Pérot interferometer with the input-output relation (189) or equivalently Eqs. (193) and (194). Comparing Eqs. (19) and (20) and (189), we obtain
${{\mathfrak{A}}}(\omega_{0}\pm\Omega)$ $=$ $-ie^{\pm i\beta}\sqrt{\kappa}\cos\theta\frac{H(\pm\Omega,L)}{h_{SQL}},$ (213)
${{\mathfrak{B}}}$ $=$ $i\sin\theta\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\left[1+i{{\mathfrak{R}}}\right],$ (214)
$D_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega)D_{d}=D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$ (215)
$=$ $e^{\pm 2i\beta}\left[i\sin\theta\hat{d}_{\pm}(\Omega)+\cos\theta\hat{a}_{\pm}(\Omega)\right]$
$+e^{\pm 2i\beta}\frac{\kappa}{2}\left[\sin\theta\left(\hat{d}_{\mp}^{\dagger}(\Omega)+\hat{d}_{\pm}(\Omega)\right)\right$.
$\quad\quad\quad\quad\quad\left.-i\cos\theta\left(\hat{a}_{\mp}^{\dagger}(\Omega)+\hat{a}_{\pm}(\Omega)\right)\right].$
Here, we note that $\mathfrak{R}$}} in Eq. (214) is an extra-term which is not included in the previous works [8]. Therefore, we estimate the order of magnitude of $\mathfrak{R}$}} as
${{\mathfrak{R}}}$ $=$ $\frac{32I_{0}\omega_{0}}{c^{2}\omega_{p}^{2}T^{2}}\left[\frac{1}{m_{EM}}+\frac{1}{m_{ITM}}\right$.
$\quad\quad\quad\quad\left.-\left(\frac{1}{m_{EM}}+\frac{3}{2m_{ITM}}\right)T+O(T^{2})\right].$
Here, we consider the case where $m:=m_{EM}=m_{ITM}$. In the case where $m:=m_{EM}=m_{ITM}$, Eq. (LABEL:eq:FrakR-defq-2-approx) is given by
${{\mathfrak{R}}}:=\frac{64I_{0}\omega_{0}}{mc^{2}\omega_{p}^{2}T^{2}}\left[1-\frac{5}{4}T+O(T^{2})\right].$ (217)
Then, we have the leading term of $\mathfrak{R}$}} for the situation $T\ll 1$ as
${{\mathfrak{R}}}$ $\sim$ $\frac{64I_{0}\omega_{0}}{mc^{2}\omega_{p}^{2}T^{2}}$ (218)
$=$ $3\times 10^{2}\times\left(\frac{I_{0}}{10^{2}W}\right)\left(\frac{\omega_{0}}{2\pi\times 10^{14}\mbox{Hz}}\right)\left(\frac{40\mbox{kg}}{m}\right)$
$\quad\quad\quad\quad\times\left(\frac{2\pi\times 1\mbox{Hz}}{\omega_{p}}\right)^{2}\left(\frac{10^{-2}}{T}\right)^{2}.$
Now, we consider the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)$ defined by Eq. (25), its expectation value (26) and the stationary noise spectral density (48). First, from Eqs. (213) and (214), the expectation value (26) of the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)$ for the DC readout is given by
$\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle$ $=$ $-\omega_{0}\sin(2\theta)\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\frac{\sqrt{\kappa}}{h_{SQL}}e^{+i\beta}H(\Omega,L)$ (219)
$+O((I_{0})^{0}).$
We note that the additional classical part $\mathfrak{R}$}} defined by Eq. (LABEL:eq:FrakR-def) in $\mathfrak{B}$}} of Eq. (213) does not contribute to the expectation value (219) due to the reality condition of the Fourier transformation $H(\Omega,L)$ of the gravitational-wave signal. Here, we note that the expectation value (219) is maximized when the offset $\theta=\pi/4$, while we do not measure $H(\Omega,L)$ when the complete dark port $\theta=0$. In the case of $\theta=0$, the gravitational-wave signal $H(\Omega,L)$ is included in the term of $O((I_{0})^{0})$. This is the well-known fact that the complete dark-port condition $\theta=0$ is meaningless in the DC-readout scheme. The factor $\sin(2\theta)$ in Eq. (219) comes from the fact that ${{\mathfrak{A}}}(\omega_{0}\pm\Omega)$ in Eq. (213) depends on $\cos\theta$ and $\mathfrak{B}$}} in Eq. (214) is proportional to $\sin\theta$. Thus, we concludes that $\theta=\pi/4$ maximize the expectation value (219).
From Eq. (219), we obtain the gravitational-wave signal $H(\Omega,L)$ through the expectation value (219) of the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)$ of the DC-readout by
$H(\Omega,L)$ $=$ $-e^{-i\beta}\frac{1}{\omega_{0}\sqrt{\kappa}\sin(2\theta)}\sqrt{\frac{\hbar\omega_{0}}{I_{0}}}h_{SQL}\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle$ (220)
$+O((I_{0})^{-1/2}).$
The factor
$-\omega_{0}\sin(2\theta)\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\frac{\sqrt{\kappa}}{h_{SQL}}e^{+i\beta}$ (221)
in Eq. (219) is regarded as the response function for the gravitational-wave signal.
Next, we consider the stationary noise spectral density (48). From the explicit expression (215) of the output operator $D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$, the amplitude- and phase-quadrature $D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$ defined by Eqs. (45) and (46), respectively, are given by
$D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$
$=$ $e^{+2i\beta}\left[-\sin\theta\hat{d}_{2}(\Omega)+\cos\theta\hat{a}_{1}(\Omega)+\kappa\sin\theta\hat{d}_{1}(\Omega)\right],$
$D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$
$=$ $e^{+2i\beta}\left[\sin\theta\hat{d}_{1}(\Omega)+\cos\theta\hat{a}_{2}(\Omega)-\kappa\cos\theta\hat{a}_{1}(\Omega)\right].$
From the expressions (LABEL:eq:B1-noise-final) and (LABEL:eq:B2-noise-final) of the amplitude- and phase-quadrature, the operator $D_{d}^{\dagger}\hat{b}_{n\Theta}D_{d}$ defined by Eq. (47) is given by
$D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$
$=$ $e^{+2i\beta}\left[\sin\theta\left(-\cos\Theta\hat{d}_{2}(\Omega)+\sin\Theta\hat{d}_{1}(\Omega)\right)\right$.
$\quad\quad\quad\left.+\cos\theta\left(\cos\Theta\hat{a}_{1}(\Omega)+\sin\Theta\hat{a}_{2}(\Omega)\right)\right$.
$\quad\quad\quad\left.+\kappa\left(\cos\Theta\sin\theta\hat{d}_{1}(\Omega)-\sin\Theta\cos\theta\hat{a}_{1}(\Omega)\right)\right].$
We also have
$D_{d}^{\dagger}\hat{b}_{n\Theta}^{\dagger}(\Omega)D_{d}$
$=$ $e^{-2i\beta}\left[\sin\theta\left(-\cos\Theta\hat{d}_{2}^{\dagger}(\Omega)+\sin\Theta\hat{d}_{1}^{\dagger}(\Omega)\right)\right$.
$\quad\quad\quad\left.+\cos\theta\left(\cos\Theta\hat{a}_{1}^{\dagger}(\Omega)+\sin\Theta\hat{a}_{2}^{\dagger}(\Omega)\right)\right$.
$\quad\quad\quad\left.+\kappa\left(\cos\Theta\sin\theta\hat{d}_{1}^{\dagger}(\Omega)-\sin\Theta\cos\theta\hat{a}_{1}^{\dagger}(\Omega)\right)\right].$
From the commutation relations (4) and (5), the definitions of the sideband quadrature (44), and the amplitude- and phase-quadratures (45) and (46), or equivalently Eqs. (190)–(192), non-vanishing commutation relations for $\hat{a}_{n1}$ and $\hat{a}_{n2}$ are summarized as
$\left[\hat{a}_{n1}(\Omega),\hat{a}_{n2}^{\dagger}(\Omega^{\prime})\right]=2\pi i\delta(\Omega-\Omega^{\prime}),$ (226)
$\left[\hat{a}_{n2}(\Omega),\hat{a}_{n1}^{\dagger}(\Omega^{\prime})\right]=-2\pi i\delta(\Omega-\Omega^{\prime}),$ (227)
$\left[\hat{a}_{n1}^{\dagger}(\Omega),\hat{a}_{n2}(\Omega^{\prime})\right]=2\pi i\delta(\Omega^{\prime}-\Omega),$ (228)
$\left[\hat{a}_{n2}^{\dagger}(\Omega),\hat{a}_{n1}(\Omega^{\prime})\right]=-2\pi i\delta(\Omega-\Omega^{\prime}).$ (229)
We also obtain the corresponding commutation relations for the amplitude- and phase-quadratures $\hat{d}_{n1}$ and $\hat{d}_{n2}$. From these commutation relations (226)–(229) for the operators $\hat{a}_{n1}$ and $\hat{a}_{n2}$, and corresponding commutation relations for the operators $\hat{d}_{n1}$ and $\hat{d}_{n2}$, we can confirm the commutation relation (49).
Through the operators $D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$ above, the straightforward calculations for the right-hand side of Eq. (48) yields
$S_{Nn}(\Omega)$ $=$ $\omega_{0}^{2}|{{\mathfrak{B}}}|^{2}\left[1-\kappa\sin(2\Theta)\cos(2\theta)\right$.
$\quad\quad\quad\quad\left.+\frac{1}{2}\kappa^{2}\left(1-\cos(2\Theta)\cos(2\theta)\right)\right].$
From this expression (LABEL:eq:Glaubers-photon-number-noise-spectral-density) of the stationary noise-spectral density and comparing with Eq. (197), we can see that $\Theta$ == $\pi/2$ gives $\left.S_{Nn}(\Omega)\right|_{\Theta=\pi/2}$ $\propto$ $S_{(K)}$.
Figure 6: The square root of the noise spectral density $\sqrt{S_{H}}$ for the $m_{EM}=m_{ITM}$ and the optimal offset $\theta=\pi/4$ case. (EQ) indicates $m_{EM}=m_{ITM}$ case. The noise spectral density $S_{H}$ is defined by Eq. (234). $\epsilon$ is the incomplete parameter discussed Sec. VIII.
Figure 7: The square root of the noise spectral density $\sqrt{S_{H}}/{{\mathfrak{R}}}$ with equal mass $(EQ)$ case where $m_{EM}=m_{ITM}$ and the offset is $\theta=\frac{\pi}{4}$. The behavior of this ratio is similar to Kimble’s signal-referred noise spectral density $\sqrt{S_{K}}$.
From Eq. (26) with $\Omega\ll\omega_{0}$, Eqs. (213) and (214), the square of the absolute value of Eq. (219) is also given by
$|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$ $=$ $\omega_{0}^{2}|{{\mathfrak{B}}}|^{2}\frac{1}{1+{{\mathfrak{R}}}^{2}}\kappa\cos^{2}\theta\frac{4}{h_{SQL}^{2}}\left|H(\Omega,L)\right|^{2}$ (231)
$+O(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0}).$
Furthermore, through the definition of $\Theta$ (${{\mathfrak{B}}}=:|{{\mathfrak{B}}}|e^{i\Theta}$), we obtain
$\sin(2\Theta)=\frac{2{{\mathfrak{R}}}}{{{\mathfrak{R}}}^{2}+1},\quad\cos(2\Theta)=\frac{{{\mathfrak{R}}}^{2}-1}{{{\mathfrak{R}}}^{2}+1}.$ (232)
Then, the signal-to-noise ratio at the photodetector is evaluated as $S_{Nn}(\Omega)/|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$. Since we regard the Fourier transformation $H(\Omega,L)$ of the gravitational waves as the signal of our measurement, the signal-to-noise ratio at the photodetector is converted to the signal-referred noise $S_{H}(\Omega)$ by
$\frac{S_{H}(\Omega)}{|H(\Omega,L)|^{2}}:=\frac{S_{Nn}(\Omega)}{|\langle\hat{s}_{{{\mathcal{N}}}}(\Omega)\rangle|^{2}}.$ (233)
Through Eqs. (LABEL:eq:Glaubers-photon-number-noise-spectral-density), (231), and (232), the signal-referred noise spectral density $S_{H}(\Omega)$ defined by Eq. (233) is given by
$S_{H}(\Omega)$ $=$ $\frac{h_{SQL}^{2}}{4\cos^{2}\theta}\left[\left(\frac{1}{\kappa}+\frac{\kappa}{2}\left(1+\cos(2\theta)\right)\right)-2{{\mathfrak{R}}}\cos(2\theta)\right$. (234)
$\quad\quad\quad\quad\left.+\left(\frac{1}{\kappa}+\frac{\kappa}{2}\left(1-\cos(2\theta)\right)\right){{\mathfrak{R}}}^{2}\right].$
In the case where ${{\mathfrak{R}}}\rightarrow 0$, we realize Kimble’s noise spectral density up to the factor $1/2$: $S_{H}(\Omega)$ == $S_{(K)}(\Omega)/2$. In the situation $\theta=\pi/4$, the signal referred stationary noise-spectral density $S_{H}(\Omega)$ is given by
$S_{H}(\Omega,\theta=\pi/4)$ $=$ $\frac{h_{SQL}^{2}}{2}\left(\frac{1}{\kappa}+\frac{\kappa}{2}\right)\left(1+{{\mathfrak{R}}}^{2}\right).$ (235)
The square root signal-referred stationary noise spectral density $S_{H}(\Omega)$ with the offset $\theta=\pi/4$ in the case $m_{EM}=m_{ITM}$ is depicted in Fig. 6. Since the noise spectral density $S_{H}$ coincides with $S_{K}/2$ and ${{\mathfrak{R}}}\neq 0$ case, the dominant order of $\mathfrak{R}$}} in $S_{H}$ is $O({{\mathfrak{R}}}^{2})$. Therefore, we show the ratio
$\sqrt{\frac{2S_{H}}{{{\mathfrak{R}}}^{2}S_{(K)}^{(K)}}}$ (236)
in Fig. 8 to clarify the difference between the derived noise-spectral density $S_{H}$ and the original Kimble’s noise spectral density. This ratio is approximately unity except for the high frequency range $\sim 20$kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$ defined by Eq. (212). We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$. Since the expectation value $|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$ given by Eq. (219) is the largest when $\theta=\pi/4$, we show Figures 6, 7 and 8 with $\theta=\pi/4$.
Figure 8: The square root ratio of the noise spectral density $S_{H}^{(EQ)}/{{\mathfrak{R}}}^{2}$ and $S_{(K)}^{(K)}$ where $S_{(K)}^{(K)}$ is defined by Eq. (212). This ratio is approximately unity except for the high frequency range $\sim 20$kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$. We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$.
The factor $\mathfrak{R}$}} in the signal referred noise spectral density $S_{H}(\Omega)$ comes from (214). Equation (214) indicates that the classical carrier field is modified by $\mathfrak{R}$}}. In DC readout schemes, we use the classical carrier field as the reference to measure gravitational-wave signals. Due to this modification of the classical carrier field appears to the signal referred noise spectral density $S_{H}(\Omega)$, although it does not appear in the expectation value (219) of the output signal due to the reality condition of the gravitational-wave signals $H(\Omega)=H^{*}(-\Omega)$. Furthermore, we also note that the existence of the factor $\mathfrak{R}$}} leads to the signal-referred noise spectral density in which the shot noise is not reduced even if the injected power $I_{0}$ is increased. This contradicts the consensus in the community of gravitational-wave experiments.
Originally, the additional classical carrier field $\mathfrak{R}$}} comes from the solutions (172) and (174) to the Heisenberg equations as the terms proportional to $2\pi\delta(\Omega)$. Furthermore, these terms come from the radiation pressure forces (162)–(165) as the classical constant force in the time domain. This classical constant force in the time domain appears in the Heisenberg equations of motion (141), (142), (145), and (146). In both classical and quantum mechanics of a forced harmonic oscillator, the constant force in the harmonic oscillator leads to changes in the equilibrium points of the oscillator. In actual experiments, the interferometers should be operated at these equilibrium points of mirrors. Therefore, to take into account the effects of the constant forces, we have to change the tuning condition (161) so that the laser is tuned at the equilibrium points of mirrors. Therefore, we have to consider the change of the tuning condition (161) in the next section VIII. To do this, we have to reconsider the Heisenberg equations for mirrors in Sec. V.
VIII Changing Tuning Point
VIII.1 Complete equilibrium tuning
Now, we consider the modification of the tuning points from the tuning conditions (161) and (177). First, we consider the modification of the tuning condition (161). To do this, we note that the cavity propagation conditions (78) and (79) yields that the retarded effects in the electric field operators are determined by
$\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)$ $=$ $\frac{L+\hat{X}_{XEM}(t-\tau)-\hat{X}_{XITM}(t-\tau)}{c},$
$\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)$ $=$ $\frac{L+\hat{X}_{YEM}(t-\tau)-\hat{X}_{YITM}(t-\tau)}{c}.$
Even when we evaluate the radiation pressure forces which affect XEM and YEM, we use Eqs. (157) and (158) in which the retarded effects in the electric field operator are determined by Eqs. (LABEL:eq:retarded-effect-cavity-X) and (LABEL:eq:retarded-effect-cavity-Y). From Eqs. (LABEL:eq:retarded-effect-cavity-X) and (LABEL:eq:retarded-effect-cavity-Y), we can see that if the operators $\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{YEM}-\hat{X}_{YITM}$ have the constant terms in time, we may always include these constant terms in $L$, i.e., in $\tau=L/c$.
As we see in Appendix B, the Fourier transformations (336) and (337) of the radiation pressure forces to XEM and YEM, respectively, include the term proportional to $2\pi\delta(\Omega)$. In the time domain, these terms represent the constant forces in time. These terms are the same in the radiation pressure forces for XEM and YEM. We denote these terms, which are proportional to $2\pi\delta(\Omega)$, as
$\omega_{p}^{2}{{\mathscr{D}}}_{EM}2\pi\delta(\Omega)$ (239)
$:=$ $N^{2}T^{2}\frac{\hbar\omega_{0}}{m_{EM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\times\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}2\pi\delta(\Omega).$
On the other hand, for ITMs, the Fourier transformations (338) and (339) of the radiation pressure forces affect XITM and YITM, respectively, and include the term proportional to $2\pi\delta(\Omega)$. In the time domain, these terms represent the constant forces in time. These terms are the same in the radiation pressure forces for XITM and YITM. We denote these terms, which are proportional to $2\pi\delta(\Omega)$, as
$\omega_{p}^{2}{{\mathscr{D}}}_{ITM}2\pi\delta(\Omega)$ (240)
$:=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\times\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\times\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega).$
Keep in our mind the existence of these constant terms ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$, the Heisenberg equations (141)–(144) of mirrors’ motion are given by
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{XITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XITM}+\omega_{p}^{2}{{\mathscr{D}}}_{ITM}$ (241)
$+\frac{1}{m_{ITM}}\hat{F}^{(fluc)}_{rpXITM}$,
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{YITM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XITM}+\omega_{p}^{2}{{\mathscr{D}}}_{ITM}$ (242)
$+\frac{1}{m_{ITM}}\hat{F}^{(fluc)}_{rpYITM}$,
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{XEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{XEM}+\omega_{p}^{2}{{\mathscr{D}}}_{EM}$ (243)
$+\frac{1}{m_{EM}}\hat{F}^{(fluc)}_{rpXEM}+\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L),$
$\!\!\!\!\frac{d^{2}}{dt^{2}}\hat{X}_{YEM}$ $=$ $-\omega_{p}^{2}\hat{X}_{YEM}+\omega_{p}^{2}{{\mathscr{D}}}_{EM}$ (244)
$+\frac{1}{m_{EM}}\hat{F}^{(fluc)}_{rpYEM}-\frac{1}{2}L\frac{d^{2}}{dt^{2}}h(t,L),$
where $\hat{F}^{(fluc)}_{rpXITM}$, $\hat{F}^{(fluc)}_{rpYITM}$, $\hat{F}^{(fluc)}_{rpXEM}$, and $\hat{F}^{(fluc)}_{rpYEM}$ the fluctuation parts of the radiation pressure forces. These do not include the classical constant part of the forces. The solutions to Eqs. (241)–(244) are given by
$\hat{X}_{XITM}$ $=:$ $\hat{X}^{(fluc)}_{XITM}+{{\mathscr{D}}}_{ITM},$ (245)
$\hat{X}_{YITM}$ $=:$ $\hat{X}^{(fluc)}_{XITM}+{{\mathscr{D}}}_{ITM},$ (246)
$\hat{X}_{XEM}$ $=:$ $\hat{X}^{(fluc)}_{XEM}+{{\mathscr{D}}}_{EM},$ (247)
$\hat{X}_{YEM}$ $=:$ $\hat{X}^{(fluc)}_{YEM}+{{\mathscr{D}}}_{EM},$ (248)
where $\hat{X}^{(fluc)}_{XITM}$, $\hat{X}^{(fluc)}_{YITM}$, $\hat{X}^{(fluc)}_{XEM}$, and $\hat{X}^{(fluc)}_{YEM}$ are solutions to Eqs. (241)–(244) with ${{\mathscr{D}}}_{ITM}={{\mathscr{D}}}_{EM}=0$.
Through the expressions (245)–(248), the retarded effects (LABEL:eq:retarded-effect-cavity-X) and (LABEL:eq:retarded-effect-cavity-Y) are given by
$\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)$ $=$ $\frac{L+{{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM}}{c}$
$+\frac{1}{c}\hat{X}_{x}^{(fluc)}(t-\tau)$
$=:$ $\tau^{(o)}+\frac{1}{c}\hat{X}_{x}^{(fluc)}(t-\tau)$
$=$ $\tau^{(o)}+\frac{1}{c}\hat{X}_{x}^{(fluc)}(t-\tau^{(o)})+O(X_{x}^{2}),$
$\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)$ $=$ $\frac{L+{{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM}}{c}$
$+\frac{1}{c}\hat{X}_{y}^{(fluc)}(t-\tau)$
$=:$ $\tau^{(o)}+\frac{1}{c}\hat{X}_{y}^{(fluc)}(t-\tau)$
$=$ $\tau^{(o)}+\frac{1}{c}\hat{X}_{y}^{(fluc)}(t-\tau^{(o)})+O(X_{y}^{2}).$
where $\hat{X}_{x}^{(fluc)}:=\hat{X}^{(fluc)}_{XEM}-\hat{X}_{XITM}^{(fluc)}$ and $\hat{X}_{y}^{(fluc)}:=\hat{X}^{(fluc)}_{YEM}-\hat{X}_{YITM}^{(fluc)}$. Here, $\tau^{(o)}$ is regarded as the operation point of the cavity length of the Fabri-Pérot interferometer. Equations (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom) indicate that we may replace
$\tau\rightarrow\tau^{(o)},$ (251)
$\hat{X}_{x,y}(t-\tau)\rightarrow\hat{X}_{x,y}^{(fluc)}(t-\tau^{(o)})$ (252)
in Eqs. (78), (79), (80)–(LABEL:eq:yarm-prop-g-and-f-g), (157), (158), and any equations derived from them within the accuracy up to $O(X_{x,y}^{2})$.
On the other hand, the retarded effects between BS and ITMs, which are evaluated in Eqs. (62), (63), (69), and (70) are determined by
$\tau_{x}^{\prime}+\frac{1}{c}\hat{X}_{XITM}(t)$ $=$ $\frac{l_{x}+\hat{X}_{XITM}(t)}{c},$ (253)
$\tau_{y}^{\prime}+\frac{1}{c}\hat{X}_{YITM}(t)$ $=$ $\frac{l_{y}+\hat{X}_{YITM}(t)}{c}.$ (254)
Furthermore, we applied the tuning conditions (177) to $(l_{x}+l_{y})/(2c)$. However, similarly to the cases of Eqs. (LABEL:eq:retarded-effect-cavity-X) and (LABEL:eq:retarded-effect-cavity-Y), $l_{x}$ and $l_{y}$ are also modified by the constant term in $\hat{X}_{XITM}(t)$ and $\hat{X}_{YITM}(t)$. As shown in Eqs. (241) and (242), we can separate the solutions $\hat{X}_{XITM}$ and $\hat{X}_{YITM}$ to Eqs. (241) and (242) into the constant terms ${{\mathscr{D}}}_{ITM}$ and the fluctuation terms $\hat{X}_{XITM}^{(fluc)}$ and $\hat{X}_{YITM}^{(fluc)}$. Then, as in the cases of Eqs. (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), we may replace Eqs. (253) and (254) as
$\tau_{x}^{\prime}+\frac{1}{c}\hat{X}_{XITM}(t)$ $=$ $\frac{l_{x}+{{\mathscr{D}}}_{ITM}}{c}+\frac{1}{c}\hat{X}_{XITM}^{(fluc)}(t)$ (255)
$=:$ $\tau_{x}^{(o)^{\prime}}+\frac{1}{c}\hat{X}_{XITM}^{(fluc)}(t),$
$\tau_{y}^{\prime}+\frac{1}{c}\hat{X}_{YITM}(t)$ $=$ $\frac{l_{y}+{{\mathscr{D}}}_{ITM}}{c}+\frac{1}{c}\hat{X}_{YITM}^{(fluc)}(t)$ (256)
$=:$ $\tau_{y}^{(o)^{\prime}}+\frac{1}{c}\hat{X}_{YITM}^{(fluc)}(t).$
As in the cases of Eqs. (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), $\tau_{x}^{(o)^{\prime}}$ and $\tau_{y}^{(o)^{\prime}}$ is regarded as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer. Equation (255) and (256) indicate that we may replace
$\tau_{x}^{\prime}\rightarrow\tau_{x}^{(o)^{\prime}},$ (257)
$\tau_{y}^{\prime}\rightarrow\tau_{y}^{(o)^{\prime}},$ (258)
$\hat{X}_{XITM}(t)\rightarrow\hat{X}_{XITM}^{(fluc)}(t),$ (259)
$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{(fluc)}(t)$ (260)
in Eqs. (62), (63), (69), (70), and any equations derived from them.
We note that Eqs.(251), (257), and (258) are regarded as the “renormalizations” of $L$, $l_{x}$, and $l_{y}$, respectively. After these renormalizations of $L$, $l_{x}$, and $l_{y}$, we apply the tuning conditions
$\omega_{0}\tau^{(o)}=2n\pi,\quad n\in{{\mathbb{N}}},$ (261)
$\omega_{0}\frac{\tau_{x}^{(o)^{\prime}}+\tau_{y}^{(o)^{\prime}}}{2}=2m\pi,\quad m\in{{\mathbb{N}}},$ (262)
instead of the tuning conditions (161) and (177). After these renormalizations, the phase offset $\theta$ is defined by (176) is replaced by
$\theta:=\omega_{0}(\tau_{y}^{(o)^{\prime}}-\tau_{x}^{(o)^{\prime}}).$ (263)
Furthermore, through the renormalization (251) and the tuning condition (261), ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ defined by Eqs. (239) and (240) are given by
${{\mathscr{D}}}_{EM}$ $=$ $T^{2}\frac{I_{0}}{m_{EM}c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}$ (264)
$\sim$ $+\frac{4I_{0}}{m_{EM}c\omega_{p}^{2}}\frac{1}{T}$ (266)
$\sim$ $+3.0\times 10^{-7}\;\mbox{m}\left(\frac{I_{0}}{10^{2}\;\mbox{W}}\right)\left(\frac{40\;\mbox{kg}}{m_{EM}}\right)$
$\quad\quad\quad\times\left(\frac{2\pi\times 1\;\mbox{Hz}}{\omega_{p}}\right)\left(\frac{10^{-2}}{T}\right).$
and
${{\mathscr{D}}}_{ITM}$ $=$ $-2\sqrt{1-T}\frac{I_{0}}{m_{ITM}c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-1}$ (267)
$\sim$ $-\frac{4I_{0}}{m_{ITM}c\omega_{p}^{2}}\frac{1}{T}$ (269)
$\sim$ $-3.0\times 10^{-7}\;\mbox{m}\left(\frac{I_{0}}{10^{2}\;\mbox{W}}\right)\left(\frac{40\;\mbox{kg}}{m_{ITM}}\right)$
$\quad\quad\quad\times\left(\frac{2\pi\times 1\;\mbox{Hz}}{\omega_{p}}\right)\left(\frac{10^{-2}}{T}\right).$
Since all $\hat{X}_{x,y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$ are replaced by $\hat{X}_{x,y}^{(fluc)}$, $\hat{X}_{XITM}^{(fluc)}$, and $\hat{X}_{YITM}^{(fluc)}$, respectively, their Fourier transformations $\hat{Z}_{x,y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ are also replaced to $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$. These $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$ does not includes the classical term which proportional to $2\pi\delta(\Omega)$. Therefore, the resulting input-output relation corresponding to Eq. (189) is given by the expression of Eq. (189) with ${{\mathfrak{R}}}=0$. In this expression of the input-output relation, there is no additional term in the classical carrier fields $\mathfrak{R}$}}, we obtain the noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ instead of Eq. (234). This recover the conventional noise-spectral density except for the fact that we have to use the definitions (187) and (188) of the variables $h_{SQL}$ and $\kappa$ with the replacement $\tau\rightarrow\tau^{(o)}$, respectively. The difference between the resulting noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ and Kimble’s noise spectral density $S_{(K)}^{(K)}$ [8] are depicted in Fig. 5 except for the factor $1/2$ in $S_{H}(\Omega)$ as discussed in Sec. VII.
Since $\hat{X}_{XITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{YITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{XEM}={{\mathscr{D}}}_{EM}$, and $\hat{X}_{YEM}={{\mathscr{D}}}_{EM}$ are equilibrium points of the pendulum for mirrors’, we can completely exclude the classical signals due to the classical radiation pressure forces $\mathfrak{R}$}} in the feedback electric current. For this reason, we call the set of the tuning conditions (261) and (262) as “complete equilibrium tuning.” This complete equilibrium tuning is one of the justifications of the ignorance of the effects of the classical part $\mathfrak{R}$}}, which comes from the radiation pressure forces.
VIII.2 Incomplete equilibrium tuning
If we achieve the above complete equilibrium tuning in experiments, we can realize Kimble’s noise spectral density [8] with the modifications of $\kappa$ and $h_{SQL}$ as an idealized case. However, in some experiments, we might not be able to accomplish the above complete equilibrium tuning for some reasons. In this subsection, we consider the case where we cannot realize the above complete equilibrium tuning and how this incompleteness appears in the noise spectral densities through our consideration within this paper.
To consider the incomplete equilibrium tuning, we introduce “operation points” so that
$\frac{l_{x}+\hat{X}_{XITM}(t)}{c}$ $=:$ $\frac{l_{x}+(1-\epsilon){{\mathscr{D}}}_{ITM}}{c}+\frac{\hat{X}_{XITM}^{\epsilon(o)}(t)}{c}$ (270)
$=:$ $\frac{l_{x}^{\epsilon(o)}}{c}+\frac{1}{c}\hat{X}_{XITM}^{\epsilon(o)}(t)$
$=:$ $\tau_{x}^{\epsilon(o)^{\prime}}+\frac{1}{c}\hat{X}_{XITM}^{\epsilon(o)}(t),$
$\hat{X}_{XITM}^{\epsilon(o)}$ $:=$ $\epsilon{{\mathscr{D}}}_{ITM}+\hat{X}_{XITM}^{(fluc)}(t),$ (271)
$\frac{l_{y}+\hat{X}_{XITM}(t)}{c}$ $=:$ $\frac{l_{y}+(1-\epsilon){{\mathscr{D}}}_{ITM}}{c}+\frac{1}{c}\hat{X}_{YITM}^{\epsilon(o)}(t)$ (272)
$=:$ $\frac{l_{y}^{\epsilon(o)}}{c}+\frac{1}{c}\hat{X}_{YITM}^{\epsilon(o)}(t)$
$=:$ $\tau_{y}^{\epsilon(o)^{\prime}}+\frac{1}{c}\hat{X}_{YITM}^{\epsilon(o)}(t),$
$\hat{X}_{YITM}^{\epsilon(o)}$ $:=$ $\epsilon{{\mathscr{D}}}_{ITM}+\hat{X}_{YITM}^{(fluc)}(t),$ (273)
through ${{\mathscr{D}}}_{ITM}$, $\hat{X}_{XITM}^{(fluc)}(t)$, and $\hat{X}_{YITM}^{(fluc)}(t)$ defined by Eqs. (240), (245), and (246), respectively. Here, $\epsilon$ is a phenomenological parameter within $[0,1]$.
As in the case of the complete equilibrium tuning, we regard $\tau_{x}^{\epsilon(o)^{\prime}}$ and $\tau_{y}^{\epsilon(o)^{\prime}}$ as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer. Equations (270) and (272) indicate that we may replace
$\tau_{x}^{\prime}\rightarrow\tau_{x}^{\epsilon(o)^{\prime}},$ (274)
$\tau_{y}^{\prime}\rightarrow\tau_{y}^{\epsilon(o)^{\prime}},$ (275)
$\hat{X}_{XITM}(t)\rightarrow\hat{X}_{XITM}^{\epsilon(o)}(t),$ (276)
$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{\epsilon(o)}(t)$ (277)
in Eqs. (62), (63), (69), (70), and any equations derived from them.
Similarly, we define the incomplete equilibrium operation point for the retarded effect $\tau+\hat{X}_{x}(t-\tau)$ and $\tau+\hat{X}_{y}(t-\tau)$ so that
$\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)$ $=$ $\frac{1}{c}\left(L+\hat{X}_{XEM}-\hat{X}_{XITM}\right)$
$=:$ $\frac{1}{c}\left(L+(1-\epsilon)({{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM})\right$.
$\quad\left.+\hat{X}_{XEM}^{\epsilon(o)}-\hat{X}^{\epsilon(o)}_{XITM}\right)$
$=:$ $\tau^{\epsilon(o)}+\frac{1}{c}\hat{X}_{x}^{\epsilon(o)}(t-\tau^{\epsilon(o)})+O(X_{x}^{2}),$
$\hat{X}_{XEM}^{\epsilon(o)}$ $:=$ $\epsilon{{\mathscr{D}}}_{EM}+\hat{X}_{XEM}^{(fluc)},$ (279)
$\hat{X}_{x}^{\epsilon(o)}$ $:=$ $\hat{X}^{\epsilon(o)}_{XEM}-\hat{X}_{XITM}^{\epsilon(o)},$ (280)
$\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)$ $=$ $\frac{1}{c}\left(L+\hat{X}_{YEM}-\hat{X}_{YITM}\right)$
$=:$ $\frac{1}{c}\left(L+(1-\epsilon)({{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM})\right$.
$\quad\left.+\hat{X}_{YEM}^{\epsilon(o)}-\hat{X}^{\epsilon(o)}_{YITM}\right)$
$=:$ $\tau^{\epsilon(o)}+\frac{1}{c}\hat{X}_{y}^{\epsilon(o)}(t-\tau^{\epsilon(o)})+O(X_{y}^{2}),$
$\hat{X}_{YEM}^{\epsilon(o)}$ $:=$ $\epsilon{{\mathscr{D}}}_{EM}+\hat{X}_{YEM}^{(fluc)},$ (282)
$\hat{X}_{y}^{\epsilon(o)}$ $:=$ $\hat{X}^{\epsilon(o)}_{YEM}-\hat{X}_{YITM}^{\epsilon(o)}.$ (283)
Here, we emphasize that $\tau^{\epsilon(o)}$ includes the relative constant displacements between XEM and XITM (or equivalently, the relative constant displacements between YEM and YITM) due to the constant displacements ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ from the radiation pressure forces as
$\tau^{\epsilon(o)}=\frac{1}{c}\left(L+(1-\epsilon)({{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM})\right),$ (284)
while the operator $\hat{X}_{x}^{\epsilon(o)}$ and $\hat{X}_{y}^{\epsilon(o)}$ includes constant displacements ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ as
$\hat{X}_{x}^{\epsilon(o)}:=\epsilon\left({{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM}\right)+\hat{X}_{x}^{(fluc)}.$ (285)
As in the case of the complete equilibrium tuning, we regard $\tau^{\epsilon(o)}$ as the operation point of the cavity length between ITMs and EMs in the Fabri-Pérot interferometer. Equations (278) and (LABEL:eq:L-incomplete-operation-point-y) indicate that we may replace
$\tau\rightarrow\tau^{\epsilon(o)},$ (286)
$\hat{X}_{x,y}(t-\tau)\rightarrow\hat{X}_{x,y}^{\epsilon(o)}(t-\tau^{(o)})$ (287)
in Eqs. (78), (79), (80)–(LABEL:eq:yarm-prop-g-and-f-g), (157), (158), and any equations derived from them within the accuracy up to $O(X_{x,y}^{2})$.
We note that Eqs. (274), (275), and (286) are regarded as the “renormalizations” of $l_{x}$, $l_{y}$, and $L$ which are different from that in Sec. VIII.1. After these renormalizations of $l_{x}$, $l_{y}$, and $L$, we apply the tuning condition
$\omega_{0}\tau^{\epsilon(o)}=2n\pi,\quad n\in{{\mathbb{N}}},$ (288)
$\omega_{0}\frac{\tau_{x}^{\epsilon(o)^{\prime}}+\tau_{y}^{\epsilon(o)^{\prime}}}{2}=2m\pi,\quad m\in{{\mathbb{N}}},$ (289)
instead of the tuning conditions (161) and (177). After these renormalizations, the phase offset $\theta$ is defined by (176) is replaced by
$\theta:=\omega_{0}(\tau_{y}^{\epsilon(o)^{\prime}}-\tau_{x}^{\epsilon(o)^{\prime}}).$ (290)
Figure 9: The schematic picture of the incomplete equilibrium tuning. The phenomenological parameter $\epsilon$ represents the deviations of the tuning point from the equilibrium point of the pendulum for the mirrors. $\epsilon=0$ corresponds to the complete equilibrium tuning. Even if we keep the mirrors at the operation point $\epsilon\neq 0$, the mirrors still feel acceleration and try to move. These accelerations for each mirror are included in the feedback electric current.
As in the case of the above complete equilibrium tuning, all $\hat{X}_{x,y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$ are replaced by $\hat{X}^{\epsilon(o)}_{x,y}$, $\hat{X}_{XITM}^{\epsilon(o)}$, and $\hat{X}_{YITM}^{\epsilon(o)}$, respectively. Furthermore, their Fourier transformations $\hat{Z}_{x,y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ are also replaced by $\hat{Z}_{x,y}^{\epsilon(o)}$, $\hat{Z}_{XITM}^{\epsilon(o)}$, and $\hat{Z}_{YITM}^{\epsilon(o)}$. These $\hat{Z}_{x,y}^{\epsilon(o)}$, $\hat{Z}_{XITM}^{\epsilon(o)}$, and $\hat{Z}_{YITM}^{\epsilon(o)}$ includes the classical term proportional to $2\pi\delta(\Omega)$ in the order of $O(\epsilon)$. Therefore, the resulting input-output relation corresponding to Eq. (189) is given by the expression of Eq. (189) with the replacement $\mathfrak{R}$}} $\rightarrow$ $\epsilon{{\mathfrak{R}}}$. In this new input-output relation, $\tau$ in the expressions of $e^{\pm 2i\beta}$ given by Eq. (186), $\kappa$ given by Eq. (187), and $h_{SQL}$ given by Eq. (188) must be replaced as Eq. (286). We note that $\epsilon=0$ corresponds to the complete equilibrium tuning discussed in Sec. VIII.1.
Note that $\hat{X}_{XITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{YITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{XEM}={{\mathscr{D}}}_{EM}$, and $\hat{X}_{YEM}={{\mathscr{D}}}_{EM}$ are equilibrium points of the pendulum for mirrors, as mentioned in Sec. VIII.1. Even if we make the mirrors stay at the operation point so that $\epsilon\neq 0$, the mirrors feel their accelerations due to the classical radiation pressure and try to move with the frequency $\omega_{p}$ of the pendulum. In this case, these accelerations are measured through the feedback control system, and its signal is included in the feedback electric current. In this case, the noise in the feedback electric current may deviate from the noise at the photodetection. For this reason, we call the tuning conditions $\epsilon\neq 0$ as “incomplete equilibrium tuning.” The schematic picture of the incomplete equilibrium tuning is depicted in Fig. 9.
Figure 10: The square-root of the explicit signal referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=0)}$ for each $\epsilon$ is depicted in the range1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (293) with $m=m_{EM}=m_{ITM}$.
We also note that the classical part $\mathfrak{R}$}} from the radiation pressure forces are given by the displacements ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ as
${{\mathfrak{R}}}$ $=$ $\frac{2\omega_{0}T}{c}\left[1-\sqrt{1-T}\right]^{-2}$
$\quad\times\left[{{\mathscr{D}}}_{EM}-\left(1-\frac{[1-\sqrt{1-T}]^{2}}{T}\right){{\mathscr{D}}}_{ITM}\right].$
These displacements ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ are given by Eqs. (264) and (267), respectively. Equation (LABEL:eq:FrakR-scrDEM-scrDITM-relation) shows that the origin of the additional classical carrier $\mathfrak{R}$}} from the radiation pressure forces is determined by the deviation from the complete equilibrium tuning.
Even in this incomplete equilibrium tuning, the expectation value of the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)$ defined by Eq. (25) is also given by Eq. (219). This is because the expectation value $\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle$ does not depend on $\mathfrak{R}$}} due to the reality condition of the gravitational-wave signal $H(\Omega)=H^{*}(-\Omega)$.
On the other hand, the strain-referred noise spectral density (233) defined by the signal-to-noise ratio at the photodetector is given by Eq. (234) with the replacement ${{\mathfrak{R}}}\rightarrow\epsilon{{\mathfrak{R}}}$, i.e.,
$\!\!\!\!\!\!\!S_{H}(\Omega)$ (292)
$=$ $\frac{h_{SQL}^{2}}{4\cos^{2}\theta}\left[\left(\frac{1}{\kappa}+\frac{\kappa}{2}\left(1+\cos(2\theta)\right)\right)-2\epsilon{{\mathfrak{R}}}\cos(2\theta)\right$.
$\quad\quad\quad\quad\left.+\left(\frac{1}{\kappa}+\frac{\kappa}{2}\left(1-\cos(2\theta)\right)\right)\epsilon^{2}{{\mathfrak{R}}}^{2}\right].$
Here, we consider the case of the complete dark port limit of the expectation value (219) and the signal-referred stationary noise spectrum density (292). In this limit, the expectation value (219) vanishes as expected. This is the property of the DC-readout. In the DC-readout scheme, we cannot measure the signal, i.e., gravitational waves in the complete dark port $\theta=0$. However, if we consider the situation where $\theta\ll 1$ but $\theta\neq 0$, we can measure the small expectation value (219) and we can estimate the gravitational-wave signal from this small expectation value. In this case, the noise spectral density (292) has its meaning and it is given by
$S_{H}(\Omega)=\frac{h_{SQL}^{2}}{4}\left[\frac{1}{\kappa}+\kappa-2\left(\epsilon{{\mathfrak{R}}}\right)+\frac{1}{\kappa}\left(\epsilon{{\mathfrak{R}}}\right)^{2}\right].$ (293)
In the case of the complete equilibrium tuning $\epsilon=0$, the noise spectral density (293) yields Kimble’s noise spectral density $S_{(K)}/2$ [8] apart from the overall factor $1/2$. Here, we consider the case where $\epsilon\neq 0$. In this case, the noise spectral density (293) is also expressed as
$S_{H}(\Omega)=\frac{h_{SQL}^{2}}{4\kappa}\left[1+\left(\kappa-\epsilon{{\mathfrak{R}}}\right)^{2}\right]\geq\frac{h_{SQL}^{2}}{4\kappa}.$ (294)
This indicates that the noise spectral density (294) have the minimum value $h_{SQL}^{2}/(4\kappa)$. The behavior of the noise spectral density (294) depicted in Fig. 10.
Equality in the last inequality in Eq. (294) is achieved when
$\kappa=\epsilon{{\mathfrak{R}}}.$ (295)
We evaluate Eq. (295) through the definition (187) of $\kappa$ and the definition (LABEL:eq:FrakR-def) of $\mathfrak{R}$}}. Here, we consider the situation where $\Omega\tau^{\epsilon(o)}\ll 1$ and use the approximation forms (204) with Eq. (205) and (LABEL:eq:FrakR-defq-2-approx) of $\kappa$ and $\mathfrak{R}$}} with the equal mass condition $m:=m_{EM}=m_{ITM}$. Then, we obtain the estimation
$1=\frac{\kappa}{\epsilon{{\mathfrak{R}}}}\sim\frac{2\omega_{p}^{2}\gamma^{2}}{\Omega^{2}(\Omega^{2}+\gamma^{2})}\sim\frac{2\omega_{p}^{2}}{\epsilon\Omega^{2}}.$ (296)
From this evaluation the minimum of $S_{H}(\Omega,\theta=0)$ is achieved at
$\Omega\sim\frac{\sqrt{2}}{\epsilon}\omega_{p}.$ (297)
This estimation is supported by the profiles in Fig. 10.
In Fig. 10, the left-top panel shows the conventional Kimble’s noise spectral density with the modification of $\kappa$ and $h_{SQL}$. However, if the incompleteness $\epsilon$ increases, the deviation from Kimble’s noise spectral density can be seen. After all, in the high-frequency region, the noise spectral density shows that the shot noise does not decrease due to the increase of the incident laser power $I_{0}$. This is merely due to the contribution of the additional classical carrier field $\epsilon{{\mathfrak{R}}}$, which may be dominant in the incomplete equilibrium tuning.
Furthermore, the minimum at Eq. (297) of the noise spectral density can be seen as a dip around the frequency (297) in Fig. 10. This dip shows a violation of the so-called “standard quantum limit” which is the envelope of the left-top figure in Fig. 10. However, as discussed in Sec. V.1, this effect is not the violation of the conventional arguments of the Heisenberg uncertainty principle which arise from the non-commutation of the position operator $\hat{X}$ and the momentum operator $\hat{P}$, i.e., $\hat{X},\hat{P}]=i\hbar$. This is because the quantum mirrors’ initial conditions, which have the information $\hat{X},\hat{P}]=i\hbar$, do not affect our consideration as discussed in Sec. V.1. Therefore, we have to say that this dip has nothing to do with the arguments of the violation of the conventional arguments of the Heisenberg uncertainty principle in Ref. [12]. In the gravitational-wave community, it is well-known that the standard quantum limit can be violated through appropriate changes of the interferometer configurations. Therefore, we may regard the dip in Fig. 10 as one of them.
Figure 11: The square-root of the explicit signal-referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=\pi/4)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (298).
Next, we examine the situation where $\langle\hat{s}_{{{\mathcal{N}}}}(\Omega)\rangle$ is the maximum for the fixed gravitational-wave signal $H(\Omega,L)$, i.e., $\theta=\pi/4$. In this case, the noise spectral density $S_{H}(\Omega)$ in Eq. (292) is given by
$S_{H}(\Omega)$ $=$ $\frac{h_{SQL}^{2}}{2}\left[\left(\frac{1}{\kappa}+\frac{\kappa}{2}\right)+\left(\frac{1}{\kappa}+\frac{\kappa}{2}\right)\epsilon^{2}{{\mathfrak{R}}}^{2}\right]$ (298)
$=$ $\frac{h_{SQL}^{2}}{2}\left(\frac{1}{\kappa}+\frac{\kappa}{2}\right)\left[1+\epsilon^{2}{{\mathfrak{R}}}^{2}\right].$
The behavior of the noise-spectral density $S_{H}(\Omega,\theta=\pi/4)$ in Eq. (298) is depicted in Fig. 11. Since the effects of the incomplete equilibrium tuning are factorized $1+\epsilon^{2}{{\mathfrak{R}}}^{2}]$, the shape of the noise spectral densities does not show any drastic changes. However, we can see that if the effect of the additional classical carrier $\epsilon{{\mathfrak{R}}}$ is considered, the power dependence of the noise spectral density is changed. As a result, the bottom figures in Fig. 11 show the property that the shot noise is not reduced even if the input power $I_{0}$ is increased. Incidentally, the bottom right panel in Fig. 11 is the same figure as Fig. 6.
Finally, we discuss the realization of the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}=0$. As shown in both Fig. 10 and Fig. 11, if we have sufficiently small $\epsilon$, we can realize the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}\sim 0$. From the input-output relation (189), if the term $\mathfrak{R}$}} in the classical carrier field, which is proportional to $2\pi\delta(\Omega)$ is negligible compared with $1$ in the same term, we can realize the ideal input-output relation. In the case of the incomplete equilibrium tuning, the phenomenological parameter $\epsilon$, which we introduced, appears as $\epsilon{{\mathfrak{R}}}$ in this classical carrier field of the input-output relation as mentioned above. Therefore, if we achieve the $\epsilon$ so that
$\epsilon{{\mathfrak{R}}}\lesssim O(10^{-1}),$ (299)
we can realize the idealized input-output relation which as shown in Ref. [8]. From the order of magnitude (218) of $\mathfrak{R}$}}, $\epsilon$ can be estimated as
$\epsilon$ $\lesssim$ $O(10^{-1}){{\mathfrak{R}}}^{-1}$ (301)
$\sim$ $O(10^{-1})\frac{mc^{2}\omega_{p}^{2}T^{2}}{64I_{0}\omega_{0}}$
$\sim$ $3\times 10^{-4}\times\left(\frac{10^{2}W}{I_{0}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right)\left(\frac{m}{40\mbox{kg}}\right)$
$\quad\quad\quad\quad\times\left(\frac{\omega_{p}}{2\pi\times 1\mbox{Hz}}\right)^{2}\left(\frac{T}{10^{-2}}\right)^{2}$
for the equal mass case $m=m_{EM}=m_{ITM}$. In terms of the mirror displacements $\epsilon{{\mathscr{D}}}_{EM}$ and $\epsilon{{\mathscr{D}}}_{ITM}$, which are estimated in Eqs. (264)–(269) as
$\epsilon{{\mathscr{D}}}_{EM}$ $\lesssim$ $+O(10^{-1})\frac{cT}{16\omega_{0}}$
$\sim$ $+3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right)$
and
$\epsilon{{\mathscr{D}}}_{ITM}$ $\lesssim$ $-O(10^{-1})\frac{cT}{16\omega_{0}}$
$\sim$ $-3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right).$
Thus, if we can control the mirrors’ positions whose deviations from the equilibrium points of the pendulum and the radiation pressure forces are less than $\epsilon{{\mathscr{D}}}_{EM}$ and $\epsilon{{\mathscr{D}}}_{ITM}$, we can realize the ideal Kimble noise spectral density with the modification of $\kappa$ and $h_{SQL}$. These $\epsilon{{\mathscr{D}}}_{EM}$ and $\epsilon{{\mathscr{D}}}_{ITM}$ are given by Eqs. (VIII.2) and (VIII.2), respectively.
IX Summary and Discussion
The primary purpose of this paper was to develop the theoretical description of the quantum noise in gravitational-wave detectors in detail, which was motivated by the theoretical developments of the measurement theory in quantum field theories. In the mathematically rigorous measurement theory in quantum mechanics [12], the specification of the finally observed quantum operator is an essential issue in the theory. We must extend this mathematical theory to quantum field theories to apply it to gravitational-wave detectors. Besides this extension of the quantum measurement theory to quantum field theories, the above specification of the finally observed quantum operator is crucially important. Actually, in quantum measurement theories, there is a famous problem of the von Neumann chain [17], i.e., where we should regard the measurement outcomes as the classical information in the sequence of the quantum measurements as emphasized in Sec. I. In this paper, we adopt the standing point that the von Neumann chain should be cut at the moment that the photodetectors convert the information in photon signals to the electric currents. Note that actual gravitational-wave detectors employ a feedback system to control the stability of the detector. However, from our standpoint, we regard the feedback current as a classical electric current. Therefore, we do not have to care about the quantum properties of the feedback control system. If this premise is wrong, we have to discuss quantum feedback control systems [18] as emphasized in Sec. I, which is beyond the current scope of this paper.
After the introduction of the notions which was used within this paper in Sec. II, we first developed the description of the DC readout scheme in terms of quantum electrodynamics. Although the analysis in these arguments of the DC readout scheme is just an extension of the balanced homodyne detection, which was discussed in Ref. [19], the essence of the DC readout scheme is the leakage of the large classical carrier field from the output port. Due to this large classical carrier field, we can separate the output electric field into orders of the large classical carrier field. After excluding the first dominant term, we can measure the output field, which includes gravitational signals as the second dominant term. Since the leaked large classical carrier field from the output port serves as the reference for the output signal in the DC readout scheme, the leaked large classical carrier field is essential in the DC readout scheme.
After the general arguments of the DC readout scheme in Sec. III, we carefully discussed the input-output relation of the Fabri-Pérot gravitational-wave detectors, because the details of the input-output relation are necessary to examine the properties of the DC readout scheme. Therefore, in Secs. IV, V, and VI, we described the detailed theoretical description of the derivation of the input-output relation of the Fabri-Pérot interferometer.
In Sec. IV, we derived the input-output relation through the propagation effects and the reflection effects due to the mirrors in the Fabri-Pérot interferometer. In the previous works of gravitational-wave detectors, it has been considered only the end mirrors’ motions. However, in this paper, we also take into account the motions of the intermediate mirrors. We obtained the input-output relation without any specification of the mirror displacements. The equations of motion of mirrors determine these mirror displacements. Since we discussed quantum motions of mirrors, we used the Heisenberg equations of motion in quantum mechanics as the equation of motion for mirrors. Furthermore, we introduce the fundamental frequency $\omega_{p}$ of the mirrors’ pendulum, where $\omega_{p}$ is estimated as $\sim 1$Hz.
In Sec. V, we briefly reviewed a quantum mechanical forced harmonic oscillator, which is essential to our arguments on the motions of mirrors in gravitational-wave detectors. Based on the understanding of this quantum forced harmonic oscillator, we explicitly derived the equations of motion for the mirrors in the Fabri-Përot interferometer in Sec. V.2. Then, we pointed out that the evaluation of the radiation pressure effects is essential. Although the details are described in Appendix B, we review the outline of the derivation of the radiation pressure forces to the intermediate and end mirrors in Fabri-Pérot gravitational-wave detectors in Sec. V.3. From the derived radiation pressure forces, we obtained the solutions to the Heisenberg equations in Sec. V.4.
[…]
Together with the input-output relation derived in Sec. IV and the solutions to the Heisenberg equation derived in Sec. V.4, we obtain the final input-output relation of the Fabri-Pérot interferometer in Sec. VI. Then, we reach the input-output relation for the Fabri-Pérot interferometer (189). Although the parameters $\beta$, $\kappa$, and $h_{SQL}$ in the input-output relation are different from these parameters in Ref. [8], the resulting input-output relation is almost the same as that derived in Ref. [8]. The difference between $\kappa$ in this paper and the $\kappa$ in Ref. [8], which denote $\kappa^{(K)}$ in this paper are depicted in Fig. 3. The difference between $h_{SQL}^{(EQ)}$ ($m=m_{EM}=m_{ITM}$) and $h_{SQL}$ in Ref. [8] which denote $h_{SQL}^{(K)}$ within this paper are depicted in Fig. 4. The frequency dependence of these differences comes from the phase difference between the motion of EMs and the motion of ITMs. Due to these differences, we reach the difference of the $S_{(K)}^{(EQ)}$ and $S_{(K)}^{(K)}$ as depicted in Fig. 5, where $S_{(K)}^{(EQ)}$ is the equal mass version of Eq. (197) and $S_{(K)}^{(K)}$ is defined by Eq. (212).
In addition to the above difference between the input-output relation (189), the difference also exists in the classical carrier part as the parameter $\mathfrak{R}$}}. This classical carrier field with the modification $\mathfrak{R}$}} is used as the reference field to measure the gravitational-wave signal in the DC readout scheme. Therefore, the classical carrier field is essential in the DC readout scheme. As mentioned in Sec. VII, the expectation value of the signal is not affected by the additional parameter $\mathfrak{R}$}} in the classical carrier field due to the reality condition of the gravitational-wave signal $H(\Omega)=H^{*}(-\Omega)$. The contribution of $\mathfrak{R}$}} in the classical carrier field affects the signal-referred noise spectral density (234). For this reason, we discussed the difference from the Kimble noise spectral density in Ref. [8]. As a result of the additional classical carrier field $\mathfrak{R}$}}, we have to conclude that the shot noise in the high-frequency range does not decrease even if the incident laser power is increased. This contradicts to the description in Ref. [8] which is the common knowledge in the community of gravitational-wave experiments.
In this situation, we reconsidered the tuning points (161) and (177) in Sec. VIII. The origin of the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation (189) is the leakage of the classical carrier from the radiation pressure forces, which are proportional to $2\pi\delta(\Omega)$ in the Fourier transformation of the radiation pressure forces. In the time domain, these terms, proportional to $2\pi\delta(\Omega)$, represent the constant forces in the radiation pressure forces. Here, we note that the equations of motion for mirrors are given in the form of a forced harmonic oscillator. In simple classical mechanics, or even in quantum mechanics, a constant force in a harmonic oscillator yields the deviation of the equilibrium point. If there exists a constant force in the Heisenberg equation of motion for a harmonic oscillator, we have to take into account the change of the equilibrium points of the mirror displacements. For this reason, we reconsidered the Heisenberg equation of motion for the mirrors and the tuning points (161) and (177) in Sec. VIII.
First, we considered the complete elimination of the deviations of the mirror displacements from their equilibrium points in Sec. VIII. Then, we showed that the deviation of the equilibrium point due to the laser radiation pressure forces is eliminated through the “renormalization” of the cavity arm length $L$, the distances $l_{x}$ and $l_{y}$ between BS and ITMs. Then, we change the tuning point as Eqs. (261) and (262). We call this elimination of the classical forces from the radiation pressure forces the “complete equilibrium tuning.” We estimate the deviations ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$ of the equilibrium points as Eqs. (266) and (269).
We also considered the incomplete equilibrium tuning in which the elimination of the deviations of the mirror displacements from their equilibrium points is not complete through the introduction of the phenomenological parameter $\epsilon$. We call this elimination the “incomplete equilibrium tuning.” If $\epsilon=0$, we can realize the complete equilibrium tuning discussed above. Instead of the complete equilibrium tuning (261) and (262), we apply the tuning condition (288) and (289). We also showed that the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation is explicitly expressed by the displacement of the equilibrium points ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$. In this incomplete equilibrium tuning, the stationary noise spectral density is given by (292). To clarify the properties of the noise spectral density (292), we consider the typical two cases. The first case is $\theta\ll 1$ but $\theta\neq 0$ and the second case is $\theta=\pi/4$. These are depicted in Figs. 10 and 11.
In the first case, the noise spectral density is given by Eq. (294). If $\epsilon$ is sufficiently large, Fig. 10 indicates the dip in the low-frequency region. Although this dip violates the “standard quantum limit” which is the envelope of the curves with different powers $I_{0}$ in $\epsilon=0$ case, this is not surprising fact that there is no contradiction with the commutation relation $\hat{X}(t),\hat{P}(t)]=i\hbar$ as discussed in Sec. V. In the second case, the effect of $\mathfrak{R}$}} appears as the factor of the noise spectral density as Eq. (298). In both cases, if the incompleteness $\epsilon$ is sufficiently large, there are cases where the shot noise in the high-frequency region does not decrease even if the incident power is increased. We also evaluate the incompleteness $\epsilon$ when we can realize the behavior near the complete equilibrium tuning in Eqs. (LABEL:eq:classical-terms-of-RPF-EM-renorm-retune-require-leading-order) and (LABEL:eq:classical-terms-of-RPF-ITM-renorm-retune-require-leading-order). This is the main result of this paper.
Here, we note that we cannot discuss the recent experimental results of the violation of the “standard quantum limit” [10, 11] through the ingredients of this paper, because we did not discuss the power recycling, the signal recycling, or the squeezed input techniques. We cannot say any relation between the dips depicted in Fig. 10 and works in Refs. [10, 11] due to the same reason. In other words, there are many rooms to be developed through an accurate understanding within this paper. We leave these rooms for future work.
Going back to the mathematically rigorous quantum measurement theory in quantum mechanics, the developed mathematical measurement theory [12] was correct. This is supported by the experimental realizations [14, 15, 16] of the derived error-disturbance relations, which are different from the Heisenberg uncertainty principle [34]. However, due to the arguments within this paper, the uncertainty relation of the position and the momentum for a point mass has nothing to do with the so-called “standard quantum limit” in the gravitational-wave detection community. We may emphasize that the developed mathematical quantum measurement theory [12] is physically correct. However, their motivation in the gravitational-wave detectors, as the precise measurement of the mirrors’ positions, was pointless.
Of course, it is well-known that this “standard quantum limit” can be estimated based on the position and the momentum uncertainty in quantum mechanics [9]. The arguments presented in this paper do not provide any answer or insight into this point. At this moment, we have to say that this point requires the delicate discussions on the quantum measurement theory in the quantum fields, because it is true that the “standard quantum limit” in gravitational-wave detectors arises from the noncommutativity (4) and (5) in the quantum electrodynamics for the optical field, namely the uncertainty relation of the electromagnetic field. In this sense, we have to say that the application of the mathematical rigorous quantum measurement theory in quantum mechanics [12] requires their extension to the quantum measurement theory for quantum field theories. Although the answer to this question is beyond the current scope of this paper, we expect that the mathematically rigorous quantum measurement theory in quantum field theory will exist. We hope that our consideration within this paper will motivate us to develop the measurement theory in quantum field theories.
Finally, we have to emphasize that the purpose of this paper is not to point out new techniques nor to develop new ideas for gravitational-wave detectors themselves. However, in this paper, we aim to explicitly show the derivation of the common knowledge in the community of the gravitational-wave detections to clarify their concepts from a more theoretically accurate point of view. Due to the more precise theoretical arguments on the noise spectral density, we could discuss the incomplete equilibrium tuning of the Frabri-Pérot interferometer as an imperfection of the interferometric gravitational-wave detectors. We hope the ingredients of this paper and this kind of research will be helpful for the further development of detector science in gravitational-wave detectors.
Appendix A Power counting photodetection
In this appendix, we consider the photodetector model in which the photocurrent is proportional to the power operator $\hat{P}_{b}(t)$ of the output optical field $\hat{E}_{b}(t)$,
$\hat{P}_{b}(t):=\frac{\kappa_{p}c}{4\pi\hbar}{\cal A}\left(\hat{E}_{b}(t)\right)^{2},$ (306)
while we discussed the model in which the photocurrent is proportional to the Glauber photon number (LABEL:eq:Glauber-photon-number). In Eq. (306), $\kappa_{p}$ is a phenomenological constant whose dimension is [time]. This coefficient $\kappa_{p}$ includes so-called “quantum efficiency.” However, $\kappa_{p}$ is not important within our discussion, though quantum efficiency is crucial for the actual experiment. We use the notations in Sec. II.1 for the output electric field $\hat{E}_{b}(t)$ for the laser.
In terms of the quadrature $\hat{B}(\omega)$ defined by Eq. (51), the power operator (306) is given by
$\hat{P}_{b}(t)$ $=$ $\frac{\kappa_{p}}{2}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}$ (307)
$\quad\times\hat{B}(\omega_{1})\hat{B}(\omega_{2})e^{-i(\omega_{1}+\omega_{2})t},$
and its Fourier transformation is given by
$\hat{{{\mathcal{P}}}}_{b}(\Omega)$
$:=$ $\int_{-\infty}^{+\infty}dt\hat{P}_{b}(t)e^{+i\Omega t}$
$=$ $\frac{\kappa_{p}}{2}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\Omega-\omega_{1})|}\hat{B}(\omega_{1})\hat{B}(\Omega-\omega_{1}).$
Here, we assume the output quadrature $\hat{b}(\Omega)$ is given in the form (19) with the expectation value (20). Substituting Eqs. (19) and (20) through Eq. (51), and considering the situation $\Omega\ll\omega_{0}$, we obtain the expectation value of the power operator $\hat{{{\mathcal{P}}}}_{b}(\Omega)$ as
$\langle\hat{{{\mathcal{P}}}}_{b}(\Omega)\rangle$ $=$ $\frac{\kappa_{p}}{2}\omega_{0}\left({{\mathfrak{B}}}^{2}2\pi\delta(\Omega-2\omega_{0})+|{{\mathfrak{B}}}|^{2}2\pi\delta(\Omega)\right$. (309)
$\quad\quad\quad\left.+|{{\mathfrak{B}}}|^{2}2\pi\delta(\Omega)+({{\mathfrak{B}}}^{*})^{2}2\pi\delta(\Omega+2\omega_{0})\right)$
$\quad+\kappa_{p}\omega_{0}\left[+{{\mathfrak{B}}}{{\mathfrak{A}}}^{*}(\omega_{0}-\Omega)+{{\mathfrak{B}}}^{*}{{\mathfrak{A}}}(\Omega+\omega_{0})\right]$
$\quad+O\left(|{{\mathfrak{B}}}|^{0}\right).$
Since gravitational-wave signals are included in ${{\mathfrak{A}}}(\omega)$, we may define the signal operator $\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)$ in the situation $\Omega\ll\omega_{0}$ by
$\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)$
$:=$ $\frac{1}{\kappa_{p}}\left[\hat{{{\mathcal{P}}}}_{b}(\Omega)\right$.
[…]
$\quad\quad\quad\quad\quad\left.\left.+|{{\mathfrak{B}}}|^{2}2\pi\delta(\Omega)+({{\mathfrak{B}}}^{*})^{2}2\pi\delta(\Omega+2\omega_{0})\right)\right].$
For the frequency range $\Omega\ll\omega_{0}$, we obtain
$\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)$ $\sim$ $\omega_{0}\left[{{\mathfrak{B}}}{{\mathfrak{A}}}^{*}(\omega_{0}-\Omega)+{{\mathfrak{B}}}^{*}{{\mathfrak{A}}}(\Omega+\omega_{0})\right$. (311)
$\quad\quad\left.+{{\mathfrak{B}}}\hat{b}_{n}^{\dagger}(\omega_{0}-\Omega)+{{\mathfrak{B}}}^{*}\hat{b}_{n}(\omega+\Omega_{0})\right]$
$+O\left(|{{\mathfrak{B}}}|^{0}\right).$
In this situation, the expectation value of $\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)$ is given by
$\langle\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)\rangle$ $\sim$ $\omega_{0}\left[{{\mathfrak{B}}}{{\mathfrak{A}}}^{*}(\omega_{0}-\Omega)+{{\mathfrak{B}}}^{*}{{\mathfrak{A}}}(\omega+\Omega_{0})\right]$ (312)
$\quad+O\left(|{{\mathfrak{B}}}|^{0}\right).$
The dominant term in Eq. (312) is same as Eqs. (26) in the case of the Glauber photon number ${{\mathcal{N}}}_{b}(\Omega)$ with the situation $\Omega\ll\omega_{0}$. The signal operator in the time domain $\hat{s}_{P_{b}}(t)$ is given by the inverse Fourier transformation of $\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega)$ as
$\hat{s}_{P_{b}}(t)$ $=$ $\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}e^{-i\Omega t}\hat{s}_{{{\mathcal{P}}}_{b}}(\Omega).$ (313)
From the definition of the signal operator (LABEL:eq:output-power-signal-operator-def-freq-domain), we can also define the noise operator $\hat{s}_{Pn}(t)$ for this signal operator $\hat{s}_{P_{b}}(t)$ by
$\hat{s}_{Pn}(t)$ (314)
$:=$ $\hat{s}_{P_{b}}(t)-\langle\hat{s}_{P_{b}}(t)\rangle$
$=$ $\sqrt{\omega_{0}}\left({{\mathfrak{B}}}e^{-i\omega_{0}t}+{{\mathfrak{B}}}^{*}e^{+i\omega_{0}t}\right)$
$\times\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}|}\left[\hat{b}_{n}(\omega_{1})\Theta(\omega_{1})\right$.
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\hat{b}_{n}^{\dagger}(-\omega_{1})\Theta(-\omega_{1})\right]e^{-i\omega_{1}t}$
$+O\left(|{{\mathfrak{B}}}|^{0}\right).$
Through this definition of the noise operator $\hat{s}_{Pn}(t)$, we can evaluate the time-averaged noise correlation function as in Sec. III by
$C_{({\rm av})s_{Pn}}(\tau)$
$:=$ $\lim_{T\rightarrow\infty}\frac{1}{T}\int_{-T/2}^{T/2}dt\frac{1}{2}\langle\mbox{in}|\hat{s}_{Pn}(t+\tau)\hat{s}_{Pn}(t)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\hat{s}_{Pn}(t)\hat{s}_{Pn}(t+\tau)|\mbox{in}\rangle.$
Similarly to the arguments in Sec. III, the noise spectral density $S_{s_{P}}(\omega)$ is also defined as
$S_{s_{P_{b}}}(\Omega)$
$:=$ $\int_{-\infty}^{+\infty}d\tau C_{({\rm av})s_{N}}(\tau)e^{+i\Omega\tau}$
$=$ $\int_{-\infty}^{+\infty}d\tau e^{+i\omega\tau}\lim_{T\rightarrow\infty}\frac{1}{T}\int_{-T/2}^{T/2}dt$
$\times\frac{1}{2}\left[\langle\hat{s}_{Pn}(t+\tau)\hat{s}_{Pn}(t)\rangle+\langle\hat{s}_{Pn}(t)\hat{s}_{Pn}(t+\tau)\rangle\right].$
Through the original definition (LABEL:eq:power-operator-quadrature-Fourier) of the power operator is different from the definition of Glauber’s photon number (III), in the situation $\omega_{0}\gg\Omega$, the tedious but straightforward calculations lead to the expression of the stationary noise spectral density including the one-point support function defined by Eq. (33). As in the case of Glauber’s photon number, using the formulae
$\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega_{0}+\omega-\omega_{1}).$ (318)
$\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}^{\dagger}(-\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega_{0}+\omega+\omega_{1}).$ (319)
$\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}(\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega-\omega_{0}+\omega_{1}).$ (320)
$\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}(\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}-\omega_{0}),$ (321)
we reach the expression
$2\pi\delta(\Omega-\Omega^{\prime})S_{s_{P}}(\Omega)$ (322)
[…]
$\quad\quad\quad\quad\left.\left.+\hat{b}_{n+}^{\dagger}(\Omega^{\prime})\hat{b}_{n+}(\Omega)+e^{-2i\Theta}\hat{b}_{n-}(\Omega^{\prime})\hat{b}_{n+}(\Omega)\right]\right\rangle$
$+O\left(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0}\right),$
where we used the sideband picture (44) and the situation $\Omega\ll\omega_{0}$. Furthermore, we introduce the amplitude- and the phase-quadratures $\hat{b}_{n1}(\Omega)$ and $\hat{b}_{n2}(\Omega)$ by Eqs. (45) and (46). Moreover, we define the operator $\hat{b}_{n\Theta}(\Omega)$ defined by Eq. (47). Then, we obtain
$2\pi\delta(\Omega-\Omega^{\prime})S_{s_{P}}(\Omega)$ (323)
$=$ $\omega_{0}^{2}|{{\mathfrak{B}}}|^{2}\left\langle\left[\hat{b}_{n\Theta}(\Omega)\hat{b}_{n\Theta}^{\dagger}(\Omega)+\hat{b}_{n\Theta}^{\dagger}(\Omega)\hat{b}_{n\Theta}(\Omega)\right]\right\rangle$
$+O\left(|{{\mathfrak{B}}}|^{1},|{{\mathfrak{B}}}|^{0}\right),$
with the commutation relation (49). This coincides with the expression of the noise spectral density (48) in the case of Glauber’s photon number model of the photodetection.
Appendix B Explicit evaluation of the radiation pressure forces to the mirrors
In this appendix, we show the evaluation of the radiation pressure forces on the mirrors in Sec. V.2. The crucial premise of this evaluation is that the radiation pressure forces on the mirrors are determined by the power of the laser, which corresponds to the pointing flux of the laser that affects the mirrors. As described in Ref. [8], the power operator of the laser is given by ${\cal A}\hat{E}^{2}(t)/(4\pi)$, where $\hat{E}(t)$ is the electric field operator which touch to mirrors and $\cal A$} is the cross-sectional area of the optical beam which introduced in Eq. (3).
As depicted in Fig. 2, the electric field operator of the laser at EMs is determined by $\hat{E}_{j_{x,y}}(t)$. As depicted in Fig. 2, the electric field $\hat{E}_{j_{x,y}}(t)$ at EMs are the propagated fields of $\hat{E}_{g_{x,y}}(t)$ at ITMs as
$\hat{E}_{j_{x,y}}(t)=\hat{E}_{g_{x,y}}\left[t-\left(\tau+\frac{1}{c}\hat{X}_{x,y}(t-\tau)\right)\right].$ (324)
In this paper, we assumed the perfect reflection at EMs as shown in Eq. (77). If we take into account the imperfection of the mirrors, we have to change the condition (77) as in Ref. [8].
For example, the radiation pressure force to XEM is given by Eq. (157). Through the notation introduced in Sec. II.1 and Eq. (84), the Fourier transformation of the radiation pressure force $\hat{F}_{rpXEM}(t)$ to XEM per the XEM mass $m_{EM}$ is given by
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}\hat{F}_{rpXEM}(t)$ $=$ $\frac{\hbar}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\hat{G}_{x}(\omega_{1})\hat{G}_{x}(\omega-\omega_{1})$ (325)
$+i\frac{2\hbar}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\omega_{1}\sqrt{|\omega_{1}\omega_{2}|}\hat{G}_{x}(\omega_{1})\hat{G}_{x}(\omega_{2})\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})$
$+O\left(\hat{X}_{x}^{2}\right).$
Substituting Eq. (57), (71), and (88) into Eq. (325), we obtain
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}\hat{F}_{rpXEM}(t)$ (326)
$=$ $T\frac{\hbar}{2m_{EM}c}e^{+i\omega\tau}e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega-\omega_{1})-\hat{A}(\omega-\omega_{1})\right)$
$+iT\frac{\hbar}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{2}\omega_{1}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})-\hat{A}(\omega_{2})\right)\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})$
$+iT\frac{\hbar}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1+2\sqrt{1-T}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})-\hat{A}(\omega_{2})\right)\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})$
$+O\left(\hat{X}^{2}_{x},\hat{X}_{XITM}^{2}\right).$
Similarly, the radiation pressure force on YEM is given by (158). Substituting Eqs. (58), (72), and (88) with the replacement $x\rightarrow y$, we obtain
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}\hat{F}_{rpYEM}(t)$ (327)
$=$ $T\frac{\hbar}{2m_{EM}c}e^{+i\omega\tau}e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega-\omega_{1})+\hat{A}(\omega-\omega_{1})\right)$
$+iT\frac{\hbar}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})+\hat{A}(\omega_{2})\right)\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})$
$+iT\frac{\hbar}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1+2\sqrt{1-T}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})+\hat{A}(\omega_{2})\right)\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})$
$+O\left(\hat{X}_{y}^{2},\hat{X}_{YITM}^{2}\right).$
Next, we consider the radiation pressure force on XITM. As depicted in Fig. 2, the radiation pressure force to XITM is given by Eq. (LABEL:eq:total-radiation-pressure-force-XITM). Through the junction condition (75) for the electric field operator at the XITM and the propagation effects (78), the radiation pressure force (LABEL:eq:total-radiation-pressure-force-XITM) on XITM yields
$\hat{F}_{rpXITM}(t)$ $=$ $\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{x}}(t)\right)^{2}-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{x}}(t)\right)^{2}+\frac{{\cal A}}{4\pi}\left(-\sqrt{1-T}\hat{E}_{f_{x}}(t)+\sqrt{T}\hat{E}_{g_{x}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)\right)\right]\right)^{2}$ (328)
$-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{x}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)\right)\right]\right)^{2}.$
Through the notation same as Eq. (7), the Fourier transformation (84), Eqs. (57), (65), (71), and (88), the tedious calculations lead to the Fourier transformation of the radiation pressure force $\hat{F}_{rpXITM}(t)$ as
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}\hat{F}_{rpXITM}(t)$ (329)
$=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[2e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega-\omega_{1})-\hat{A}(\omega-\omega_{1})\right)$
$-i\sqrt{1-T}\frac{\hbar}{m_{ITM}c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})-\hat{A}(\omega_{2})\right)\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})$
$-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})-\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})-\hat{A}(\omega_{2})\right)\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})$
$+O\left(\hat{X}_{x}^{2},\hat{X}_{XITM}^{2}\right).$
Similarly, the radiation pressure force on the YITM is given by Eq. (LABEL:eq:total-radiation-pressure-force-YITM), which yields
$\hat{F}_{rpYITM}(t)$ $=$ $\frac{{\cal A}}{4\pi}\left(\hat{E}_{f_{y}}(t)\right)^{2}-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{y}}(t)\right)^{2}+\frac{{\cal A}}{4\pi}\left(-\sqrt{1-T}\hat{E}_{f_{y}}(t)+\sqrt{T}\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right]\right)^{2}$ (330)
$-\frac{{\cal A}}{4\pi}\left(\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right]\right)^{2}.$
As in the case of $\hat{F}_{rpXITM}(t)$, we used Eqs. (79) and (82). Furthermore, through the notation same as Eq. (7), the Fourier transformation (84), Eqs. (58), (66), (72), and (88) with the replacement $x\rightarrow y$, the tedious calculations lead to the Fourier transformation of the radiation pressure force $\hat{F}_{rpYITM}(t)$ as
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}\hat{F}_{rpYITM}(t)$ (331)
$=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[2e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega-\omega_{1})+\hat{A}(\omega-\omega_{1})\right)$
$-i\sqrt{1-T}\frac{\hbar}{m_{ITM}c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})+\hat{A}(\omega_{2})\right)\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})$
$-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left(\hat{D}(\omega_{1})+\hat{A}(\omega_{1})\right)\left(\hat{D}(\omega_{2})+\hat{A}(\omega_{2})\right)\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})$
$+O\left(\hat{X}_{y}^{2},\hat{X}_{YITM}^{2}\right).$
Here, we note that the radiation pressure forces (326), (327), (329), and (331) are given by the input quadrature $\hat{D}(\omega)$ and $\hat{A}(\omega)$ with the Fourier transformations of the displacements $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$. Within this paper, we do not consider the power recycling technique, nor the signal recycling techniques [8, 9, 22, 23, 24]. If we take into account the power recycling technique, we have to use the reflected optical field quadrature by the power recycling mirror as the input quadrature $\hat{D}(\omega)$. On the other hand, if we take into account the signal recycling technique, we have to use the reflected optical field quadrature by the signal recycling mirror as the input quadrature $\hat{A}(\omega)$. In any case, we have to note that even when we consider the power recycling technique and the signal recycling technique, the expressions of the radiation pressure forces (326), (327), (329), and (331) should be used.
We also note that we have to evaluate the Fourier transformations of the mirror displacement operators $D_{d}^{\dagger}\hat{Z}_{x}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{y}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{XITM}D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{YITM}D_{d}$ instead of operators $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ when we evaluate the input-output relation (117). Here, the operator $D_{d}$ is the displacement operator associated with the coherent state for the optical quadrature $\hat{D}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ which is defined by Eq. (101). For the quadrature $\hat{D}(\omega)$, $D_{d}^{\dagger}\hat{D}(\omega)D_{d}$ are given by Eq. (104). Trivially, the quadrature $\hat{A}(\omega)$ commutes with the displacement operator $D_{d}$. Keep in our mind the properties of the displacement operator $D_{d}$, we evaluate the radiation pressure forces (326), (327), (329), and (331) by the operations $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right. Furthermore, we ignore the terms include $\hat{D}_{v}(\omega)\hat{D}_{v}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{A}(\omega^{\prime})$, and $\hat{A}(\omega)\hat{A}(\omega^{\prime})$. We symbolically denote these terms $\hat{D}_{v}^{2}$, $\hat{D}_{v}\hat{A}$, and $\hat{A}^{2}$. Moreover, we also ignored the terms that include $\hat{D}_{v}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$. We symbolically denote these terms as $\hat{D}_{v}\hat{X}$ and $\hat{A}\hat{X}$. Through these evaluations, the radiation pressure forces (326), (327), (329), and (331) are given by
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (332)
$=$ $\frac{\hbar T}{2m_{EM}c}e^{+i\omega(\tau+\tau_{x}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega-\omega_{1})$
$+\frac{\hbar T}{m_{EM}c}e^{+i\omega(\tau+\tau_{x}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\left(\hat{D}_{v}(\omega-\omega_{1})-\hat{A}(\omega-\omega_{1})\right)$
$+i\frac{\hbar T}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{2}\omega_{1}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
$+i\frac{\hbar T}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1+2\sqrt{1-T}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
[…]
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (333)
$=$ $\frac{\hbar T}{2m_{EM}c}e^{+i\omega(\tau+\tau_{y}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega-\omega_{1})$
$+\frac{\hbar T}{m_{EM}c}e^{+i\omega(\tau+\tau_{y}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega-\omega_{1})\left(\hat{D}_{v}(\omega_{1})+\hat{A}(\omega_{1})\right)$
$+i\frac{\hbar T}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
$+i\frac{\hbar T}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1+2\sqrt{1-T}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
[…]
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$ (334)
$=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[2e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega-\omega_{1})$
$-2\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[e^{+2i\omega_{1}\tau}+e^{+2i(\omega-\omega_{1})\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\left(\hat{D}_{v}(\omega-\omega_{1})-\hat{A}(\omega-\omega_{1})\right)$
$-i\sqrt{1-T}\frac{\hbar}{m_{ITM}c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
$-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
[…]
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$ (335)
$=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[2e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega-\omega_{1})$
$-2\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[e^{+2i\omega_{1}\tau}+e^{+2i(\omega-\omega_{1})\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\left(\hat{D}_{v}(\omega-\omega_{1})+\hat{A}(\omega-\omega_{1})\right)$
$-i\sqrt{1-T}\frac{\hbar}{m_{ITM}c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
$-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
[…]
Next, we consider the case of the monochromatic incident laser. In this case, the classical part $\hat{D}_{c}(\omega)$ of the incident quadrature $\hat{D}(\omega)$ is given by Eq. (IV.5.1) with Eq. (114). Substituting Eq. (IV.5.1) into Eqs. (332)–(335), we can evaluate the linearized radiation pressure forces in the case of the monochromatic incident laser with the central frequency $\omega_{0}$. The resulting expressions are naturally expressed in the sideband picture for the frequency of $\omega_{0}+\Omega$ with the sideband frequency $\Omega$. As discussed in Sec. IV.5.2, we ignore the rapidly oscillating terms in the frequencies $2\omega_{0}\pm\Omega$. Then, we obtain
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (336)
$=$ $+\frac{N^{2}T\hbar\omega_{0}}{m_{EM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}2\pi\delta(\Omega)$
$+\frac{NT\hbar}{m_{EM}c}e^{+i\Omega(\tau+\tau_{x}^{\prime})}\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)$
$+\frac{NT\hbar}{m_{EM}c}e^{+i\Omega(\tau+\tau_{x}^{\prime})}\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)$
$+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+i\Omega\tau}\left[1-e^{+2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
$-4N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+2i\Omega\tau}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (337)
$=$ $+\frac{N^{2}T\hbar\omega_{0}}{m_{EM}c}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}2\pi\delta(\Omega)$
$+\frac{NT\hbar}{m_{EM}c}e^{+i\Omega(\tau+\tau_{y}^{\prime})}\sqrt{|(\Omega-\omega_{0})\omega_{0}|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)$
$+\frac{NT\hbar}{m_{EM}c}e^{+i\Omega(\tau+\tau_{y}^{\prime})}\sqrt{|(\Omega+\omega_{0})\omega_{0}|}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)$
$+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+i\Omega\tau}\left[1-e^{+2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
$-4N^{2}T\sqrt{1-T}e^{+2i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (338)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$
$=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega)$
$-N\sqrt{1-T}\frac{\hbar}{m_{ITM}c}e^{+i\Omega\tau_{x}^{\prime}}\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[e^{+2i\omega_{0}\tau}+e^{+2i(\Omega-\omega_{0})\tau}-\sqrt{1-T}(1+e^{+2i\Omega\tau})\right]\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)$
$-N\sqrt{1-T}\frac{\hbar}{m_{ITM}c}e^{+i\Omega\tau_{x}^{\prime}}\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[e^{-2i\omega_{0}\tau}+e^{+2i(\Omega+\omega_{0})\tau}-\sqrt{1-T}(1+e^{+2i\Omega\tau})\right]\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)$
$-N^{2}T\sqrt{1-T}2\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
$+4N^{2}T\sqrt{1-T}e^{+i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (339)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$
$=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega)$
$-N\sqrt{1-T}\frac{\hbar}{m_{ITM}c}e^{+i\Omega\tau_{y}^{\prime}}\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[e^{+2i\omega_{0}\tau}+e^{+2i(\Omega-\omega_{0})\tau}-\sqrt{1-T}(1+e^{+2i\Omega\tau})\right]\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)$
$-2N\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\Omega\tau_{y}^{\prime}}\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[e^{-2i\omega_{0}\tau}+e^{+2i(\Omega+\omega_{0})\tau}-\sqrt{1-T}(1+e^{+2i\Omega\tau})\right]\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)$
$-2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
$+4N^{2}T\sqrt{1-T}e^{+i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"}.$
Here, ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$ are defined in Eq. (147) and (148), respectively. Furthermore, from the definition (LABEL:eq:radiation-pressure-Fourier-X-def) and (LABEL:eq:radiation-pressure-Fourier-Y-def) of ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$, we obtain
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (340)
$=$ $\frac{N^{2}\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
$+\frac{N\hbar}{c}e^{+i\Omega\tau_{x}^{\prime}}e^{+i\Omega\tau}\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos((2\omega_{0}-\Omega)\tau)-\sqrt{1-T}\cos(\Omega\tau)\right]\right]\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)$
$+\frac{N\hbar}{c}e^{+i\Omega\tau_{x}^{\prime}}e^{+i\Omega\tau}\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos((2\omega_{0}+\Omega)\tau)-\sqrt{1-T}\cos(\Omega\tau)\right]\right]\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)$
$+\sin(2\omega_{0}\tau)\frac{2N^{2}T\sqrt{1-T}\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
$-\sin(2\omega_{0}\tau)e^{+i\Omega\tau}\frac{4N^{2}T\sqrt{1-T}\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ (341)
$=$ $N^{2}\frac{\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
$+N\frac{\hbar}{c}e^{+i\Omega\tau}e^{+i\Omega\tau_{y}^{\prime}}\sqrt{|(\Omega-\omega_{0})\omega_{0}|}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos((2\omega_{0}-\Omega)\tau)-\sqrt{1-T}\cos(\Omega\tau)\right]\right]\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)$
$+N\frac{\hbar}{c}e^{+i\Omega\tau}e^{+i\Omega\tau_{y}^{\prime}}\sqrt{|(\Omega+\omega_{0})\omega_{0}|}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos((2\omega_{0}+\omega)\tau)-\sqrt{1-T}\cos(\omega\tau)\right]\right]\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)$
$+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
$-4N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)e^{+i\Omega\tau}\frac{\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
[…]
$+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"}.$
In these expressions (338), (339), (340), and (341), we have the terms proportional to $D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z_{y}}(\Omega)D_{d}$ which includes the factor $\sin(2\omega_{0}\tau)$. These terms correspond to the optical spring effects of the Fabri-Pérot interferometer [32, 33]. However, these optical spring effects can be ignored through the tuning condition (161) due to the factor $\sin(2\omega_{0}\tau)$. In this paper, we only consider the situation of this tuning, in which there is no optical spring effect.
Appendix C Consistency relation (V.1) of the operator $\hat{{{\mathcal{F}}}}(\omega)$
In this appendix, we consider the consistency relation (V.1) through the radiation pressure forces (162)–(165). However, the confirmation of the consistency relation (V.1) for these four radiation pressure forces are essentially identical. Therefore, in this appendix, we evaluate the consistency relation (V.1) only for the radiation pressure force (162).
The radiation pressure force (162) includes a classical part. We consider the commutation relation of the radiation pressure force, and this classical part does not contribute to the commutation relation. Therefore, we ignore the classical part in Eq. (162). Furthermore, we only consider the situation $\Omega\ll\omega_{0}$. In this case, the radiation pressure force (162) is symbolically represented by
$\!\!\!\!\!\!\!\!\!\!\!\!\!-\frac{1}{\mu}\hat{\cal F}(\Omega)$ (342)
$=$ $\frac{\omega_{0}}{\mu_{XITM}(\Omega)}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right$.
$\quad\quad\quad\quad\quad\quad\left.+\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)$
$=$ $\frac{\omega_{0}}{\mu_{XITM}(\Omega)}\left(\hat{d}^{\dagger}(\omega_{0}-\Omega)-\hat{a}^{\dagger}(\omega_{0}-\Omega)\right$.
$\quad\quad\quad\quad\quad\quad\left.+\hat{d}(\omega_{0}+\Omega)-\hat{a}(\omega_{0}+\Omega)\right),$
where
$\frac{1}{\mu_{XITM}(\Omega)}$ $:=$ $\frac{2N\hbar\sqrt{1-T}}{m_{ITM}c}e^{+i\Omega(\tau+\tau_{x}^{\prime})}$
$\quad\times\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau).$
Now, we evaluate the consistency relation (V.1) for the radiation pressure force (342) as
$\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\frac{\omega-\omega_{1}}{\left(\omega_{1}^{2}-\omega_{p}^{2}\right)\left((\omega-\omega_{1})^{2}-\omega_{p}^{2}\right)}$ (344)
$\quad\quad\times\left[\frac{1}{\mu}\hat{{{\mathcal{F}}}}(\omega_{1}),\frac{1}{\mu}\hat{{{\mathcal{F}}}}(\omega-\omega_{1})\right]$
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\frac{\omega-\omega_{1}}{\left(\omega_{1}^{2}-\omega_{p}^{2}\right)\left((\omega-\omega_{1})^{2}-\omega_{p}^{2}\right)}$
$\quad\quad\times\frac{\omega_{0}^{2}}{\mu_{XITM}(\omega_{1})\mu_{XITM}(\omega-\omega_{1})}$
$\quad\quad\times\left[\hat{d}^{\dagger}(\omega_{0}-\omega_{1})-\hat{a}^{\dagger}(\omega_{0}-\omega_{1})\right$.
$\quad\quad\quad\quad\left.+\hat{d}(\omega_{0}+\omega_{1})-\hat{a}(\omega_{0}+\omega_{1}),\right$.
$\quad\quad\quad\quad\left.\hat{d}^{\dagger}(\omega_{0}-(\omega-\omega_{1}))-\hat{a}^{\dagger}(\omega_{0}-(\omega-\omega_{1}))\right$.
$\quad\quad\quad\quad\left.+\hat{d}(\omega_{0}+(\omega-\omega_{1}))-\hat{a}(\omega_{0}+(\omega-\omega_{1}))\right]$
$=$ $\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\frac{\omega-\omega_{1}}{\left(\omega_{1}^{2}-\omega_{p}^{2}\right)\left((\omega-\omega_{1})^{2}-\omega_{p}^{2}\right)}$
$\quad\quad\times\frac{\omega_{0}^{2}}{\mu_{XITM}(\omega_{1})\mu_{XITM}(\omega-\omega_{1})}$
$\quad\quad\times\left\{-2\pi\delta(\omega)-2\pi\delta(\omega)+2\pi\delta(\omega)+2\pi\delta(\omega)\right\}$
$=$ $0.$
Then, we have confirmed the consistency relation (V.1) for the radiation pressure force (342), i.e., the radiation pressure force (162).
For the other radiation pressure forces (163)–(165), the evaluations of the consistency relation (V.1) are similar to that for the radiation pressure force (162) shown above.
Acknowledgements.
The author acknowledges Masa-Katsu Fujimoto for valuable discussions and his continuous encouragement. He also thanks Tomotada Akutsu, Takayuki Tomaru, and colleagues in the Gravitational Science Project at the National Astronomical Observatory of Japan for their continuous encouragement and conversations.

### v2

@@
-Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector
+Theoretical Detailed Analyses for DC readout and a Fabry-Pérot gravitational-wave detector
 Abstract
-The quantum expectation value and the stationary noise spectral density for a Fabri-Pérot gravitational-wave detector with a DC readout scheme are discussed in detail only through the quantum electrodynamics of lasers and the Heisenberg equations of mirrors’ motion.
+The quantum expectation value and the stationary noise spectral density for a Fabry-Pérot gravitational-wave detector with a DC readout scheme are discussed in detail only through the quantum electrodynamics of lasers and the Heisenberg equations of mirrors’ motion.
 We demonstrate that the initial conditions of the mirrors’ motion concentrate around the fundamental frequency of the pendulum and are not related to the frequency range of our interest.
-Although in the ideal case, there is a consensus that the shot noise for the laser in the high-frequency range in the signal-referred noise spectral density decreases if the injected laser power is increased, our obtained noise spectral density shows that the shot noise does not decrease even if the injected laser power is increased.
-This is due to the leakage of the classical radiation pressure forces from the classical carrier field to the output port, and the classical carrier field is used as the reference in the DC readout scheme.
-Since the classical radiation pressure forces affect the mirror motion as a classical constant force, this classical constant force changes the equilibrium point of the pendulum of mirrors’ motion.
-To recover the ideal case, we must consider adjusting the interferometer’s tuning point to place the mirrors in equilibrium points.
+Although, in the ideal case, there is consensus that the shot-noise contribution from the laser to the high-frequency range of the signal-referred noise spectral density decreases as the injected laser power increases, our derived noise spectral density shows that the shot-noise contribution does not decrease.
+This is due to leakage of classical radiation pressure forces from the carrier field to the output port, and the carrier field is used as the reference in the DC readout scheme.
+Since classical radiation pressure acts as a constant force, it shifts the pendulum’s equilibrium point of the mirrors’ motion.
+To recover the ideal case, we must consider adjusting the interferometer’s tuning point to place the mirrors at their equilibrium positions.
 We investigate the case where the equilibrium tuning is incomplete and show that the behavior of the above shot noise is due to this incompleteness.
-We also discuss the maximal deviation of the mirror displacements from the equilibrium point in the incomplete tuning to recover the near ideal case.
+We also discuss the maximum deviation of the mirror displacements from the equilibrium point during incomplete tuning to recover a near-ideal case.
 gravitational-wave detector; quantum noise; DC readout scheme
@@
 A decade ago, the first detection of the gravitational-wave signal GW150914 from a black hole-black hole binary was achieved.
-This event marked the beginning of gravitational-wave astronomy and multi-messenger astronomy including gravitational-wave detection [1, 2].
-We are now at the stage where we can directly measure gravitational waves, and we can conduct scientific research through these gravitational-wave events.
+This event was the beginning of gravitational-wave astronomy and multi-messenger astronomy including gravitational-wave detection [1, 2].
+We are now at a stage where we can directly measure gravitational waves and conduct scientific research using these events.
 Many events, mainly from black hole-black hole binaries, have already been detected.
-We can also expect that one of the future directions of gravitational-wave astronomy is the development as a “precise science.” This will involve detailed studies of source science from the astrophysical point of view, the tests of general relativity, and the developments of the global network of gravitational-wave detectors [2, 3, 4, 5].
-In addition to the current network of ground-based detectors, as future ground-based gravitational-wave detectors, the projects of Einstein Telescope [6] and Cosmic Explorer [7] are in progress, aiming to achieve more sensitive detections.
+We can also expect that one of the future directions of gravitational-wave astronomy is to develop as a “precise science”, as we can easily extrapolate from the current achievements [3].
+These involve detailed studies of source science from the astrophysical point of view, the tests of general relativity, and the developments of the global network of gravitational-wave detectors [2, 4, 5, 6].
+In addition to the current network of ground-based detectors, as future ground-based gravitational-wave detectors, the projects of Einstein Telescope [7] and Cosmic Explorer [8] are in progress, aiming to achieve better sensitivity.
 To advance the field of gravitational-wave detections as a more precise science, it is essential to improve detector sensitivity.
 We note that theoretical developments in detector science are also crucial for enhancing our basic understanding of these gravitational-wave detectors.
-The current gravitational-wave detectors are limited by the shot noise, which is one of the quantum noises in the interferometer.
-In interferometric gravitational detectors, it is considered that there are two kinds of quantum noises.
-One is the shot noise of the laser, which constrains the sensitivity in the high-frequency range of the gravitational-wave detectors.
-On the other hand, the quantum radiation pressure forces are considered as a constraint on the sensitivity in the low-frequency region.
-In the common understanding in the gravitational-wave detection community, if the injected power is increased, the shot noise is decreased, while the quantum radiation pressure forces are increased [8].
-The envelop of this trade-off relation in the signal-referred noise spectral density is regarded as the “standard quantum limit.” It is also a common understanding in the gravitational-wave detection community that this standard quantum limit is also estimated from the Heisenberg uncertainty relation of the noncommutativity of the quantum positions and the momenta of test masses [9].
-Furthermore, in the current research on the gravitational-wave detection, there are some reports which state that the current LIGO gravitational-wave detectors already violate the above “standard quantum limit” [10, 11].
+In the high-frequency part of the noise spectral density, current gravitational-wave detectors are limited by shot noise, one of the quantum noises in the interferometer.
+In interferometric gravitational detectors, it is considered that there are two kinds of quantum noises, although many plots in the gravitational-wave community do not distinguish these two quantum noises, e.g.
+see Ref.
+[9].
+One is the shot noise of the laser, which limits the sensitivity of gravitational-wave detectors at high frequencies.
+On the other hand, quantum radiation pressure forces are considered a constraint on sensitivity in the low-frequency region.
+Note that, in the current experimental situation, this quantum radiation-pressure noise remains subdominant in this low-frequency region.
+In the common understanding in the gravitational-wave detection community, if the injected power is increased, the shot-noise contribution to the signal-referred noise spectral density is decreased, while the contribution of the quantum radiation pressure forces is increased [10].
+The envelop of this trade-off relation in the signal-referred noise spectral density is regarded as the “standard quantum limit.” It is also a common understanding in the gravitational-wave detection community that this standard quantum limit is estimated from the Heisenberg uncertainty relation arising from the noncommutativity of the quantum positions and momenta of free test masses [11].
+Furthermore, in the current research on gravitational-wave detection, there are some reports which state that the current LIGO gravitational-wave detectors already surpassed the above “standard quantum limit” [12, 13].
 Even from a theoretical perspective, this presents an opportune moment to revisit and refine the theoretical arguments, aiming for greater accuracy than those presented in previous works in Refs.
@@
 On the other hand, a mathematically rigorous quantum measurement theory in quantum mechanics has also been developed (Ref.
-[12] and the references therein), which introduces new error-disturbance relations, distinct from the Heisenberg uncertainty relation.
+[14] and the references therein), introducing new error-disturbance relations distinct from the Heisenberg uncertainty relation.
 These new error-disturbance relations have already been confirmed through experiments [16, 17, 18].
-One of the motivations of this development was the detection of gravitational-waves [13].
+One of the motivations of this development was the detection of gravitational waves [15].
 However, the actual application of this mathematical theory to the gravitational-wave detectors requires its extension to the quantum field theories, because the quantum noise in gravitational-wave detectors is analyzed through the quantum field theories of lasers [10].
 Moreover, in the quantum measurement theory, it is essential to specify the final measured operator in the quantum measurement process due to the von Neumann chain problem [19].
-In interferometric gravitational-wave detectors, the directly measured quantum operator is identified within the “readout scheme” of the detectors.
-While current gravitational-wave detectors utilize a feedback control system, and the final measured data of gravitational-wave detectors consist of the electric currents from this feedback control, the readout scheme in gravitational-wave detectors is the optical system that determines the finally measured quantum operator in the optical fields detected at the photodetectors, which receive signals from the main interferometer.
+In interferometric gravitational-wave detectors, the directly measured quantum operator is identified within the detectors’ “readout scheme.” While current gravitational-wave detectors utilize a feedback control system, and the final measured data of gravitational-wave detectors consist of the electric currents from this feedback control, the readout scheme in gravitational-wave detectors is the optical system that determines the final measured quantum operator in the optical fields detected at the photodetectors, which receive signals from the main interferometer.
 Therefore, research into this readout scheme is crucial for advancing mathematical quantum measurement theory and its application to gravitational-wave detections.
-Due to the issue of the von Neumann chain, this paper adopts the perspective that the quantum properties in the laser interferometers are preserved until the photons are detected.
-At the photo-detectors, the quantum information of photons is transformed into an electric current.
-The processes that occur within the photodetectors are complex and vary depending on the type of device used.
-As a result, we speculate that quantum decoherence takes place during this photodetection process, and the resulting electric current, generated from the statistical results of the detection, can be regarded as classical currents.
+Due to the von Neumann chain issue, this paper adopts the perspective that the quantum properties of the photons in the laser interferometers are preserved until they are detected.
+At the photodetectors, the quantum information of photons is transformed into an electric current.
+The processes that occur within photodetectors are complex and vary depending on the device type.
+As a result, we speculate that quantum decoherence occurs during this photodetection process, and the resulting electric current, derived from the statistical results of the detection, can be regarded as a classical current.
 If this perspective is incorrect and the quantum nature is retained even in the feedback electric current, we would need to explore the concept of a quantum feedback control system [20].
@@
 Our first step is to investigate the DC-readout scheme from a quantum theoretical perspective.
-We previously discussed the mathematically rigorous quantum description of the balanced homodyne detection as a readout scheme in gravitational-wave detectors [19], since the balanced homodyne detection is planned for installation in future gravitational-wave detectors.
-In this paper, we will consider the DC readout scheme as a readout scheme for gravitational-wave detectors.
+We previously discussed the mathematically rigorous quantum description of the balanced homodyne detection as a readout scheme in gravitational-wave detectors [21], since the balanced homodyne detection is planned for installation in future gravitational-wave detectors [22].
+In this paper, we will consider the DC readout scheme for gravitational-wave detectors.
 Our investigation in this paper is a natural extension of the quantum theoretical arguments of a balanced homodyne detection [21].
 The key aspect of our argument lies in the specification of the quantum operator that is finally observed.
-We assume that the finally observed quantum operator in the photodetectors is Glauber’s photon number [20], as discussed in Sec.
+We assume that the final observed quantum operator in the photodetectors is Glauber’s photon number [23], as discussed in Sec.
 III.
-It is often stated that the finally observed quantum operator is the power operator of the laser.
+For the monochromatic case, we consider the number operator $\hat{a}^{\dagger}(\omega)\hat{a}(\omega)$ for the quantum field associated with the annihilation operator $\hat{a}(\omega)$.
+Glauber’s photon number is a natural extension of this single-mode number operator to the multi-mode electric field.
+It is often stated that the final observed quantum operator is the power operator of the laser.
 However, even when we utilize the laser’s power operator, the conclusions we arrive at for the DC readout scheme are consistent with those related to Glauber’s photon number.
@@
 This variation depends on the photodetection device, as mentioned above.
-Therefore, the specification of our finally measured quantum operator is essential to our arguments.
-After discussing the general arguments related to the DC readout scheme, we will examine the input-output relation for a Fabri-Pérot gravitational-wave detector.
-A specific input-output relation is essential for a detailed analysis of the DC readout scheme.
-We aim to derive this input-output relation of a Fabri-Pérot gravitational-wave detector, starting without specifying the small motions of the mirrors.
-These small motions of mirrors encapsulate information about external forces, which include gravitational-wave signals and the radiation pressure forces from the laser.
-In this paper, we describe the small motions of mirrors using the Heisenberg equation in quantum mechanics for a forced harmonic oscillator.
-While previous literature has often considered the mirrors’ small motions as free except for the gravitational-wave signal and the radiation pressure forces, we have introduced the fundamental frequency $\omega_{p}$ of the vibration isolation system of the pendulum supporting the mirrors.
+Therefore, the specification of our final measured quantum operator is essential to our arguments.
+After discussing the general arguments related to the DC readout scheme, we will examine the input-output relation for a Fabry-Pérot gravitational-wave detector.
+A specific input-output relation is essential for a detailed analysis of the DC readout scheme [25].
+We aim to derive this input-output relation of a Fabry-Pérot gravitational-wave detector, starting without specifying the small motions of the mirrors.
+These small mirror motions encapsulate information about external forces, including gravitational-wave signals and radiation pressure from the laser.
+In this paper, we describe the small motions of mirrors using the Heisenberg equation in quantum mechanics for a forced harmonic oscillator [26].
+While the previous literature [10, 11] has often considered the mirrors’ small motions as free except for the gravitational-wave signal and the radiation pressure forces, we have introduced the fundamental frequency $\omega_{p}$ of the vibration isolation system of the pendulum supporting the mirrors.
 After solving the Heisenberg equations for the motion of the mirrors, we find that the initial conditions for the forced harmonic oscillator concentrate at the frequency $\omega_{p}$, which is outside the frequency range of our interest.
 This indicates that the mirrors’ initial conditions, which have information about the noncommutativity of the position and the momentum in quantum mechanics, do not relate to the frequency range relevant to gravitational-wave detectors.
-We find that the classical radiation pressure force also affects the mirror, which is finite due to the introduced fundamental frequency $\omega_{p}$ of pendulums.
+We find that the classical radiation pressure force also acts on the mirror, which is finite due to the introduced fundamental frequency $\omega_{p}$ of the pendulums.
 This classical radiation pressure manifests as a classical carrier field that leaks to the output port of the interferometer.
@@
 We discuss the effects of this leakage of classical radiation pressure forces in the stationary noise spectral densities.
-Notably, when we consider the effects of this leakage, there is a situation where the shot noise contribution to the signal-referred noise spectral density in the high-frequency range does not decrease, even when the incident laser power is increased.
+Notably, when we consider the effects of this leakage, there is a situation where the shot-noise contribution to the signal-referred noise spectral density in the high-frequency range does not decrease, even when the incident laser power is increased.
 To avoid the effects of the leakage of the classical radiation pressure forces, we have to adjust the tuning point of the Fabry-Pérot interferometer very carefully.
-We also estimate the maximum deviation from the ideal mirror position which realizes the near ideal case.
+We also estimate the maximum deviation from the ideal mirror position, which realizes the near ideal case.
 Figure 1: The setup for the Fabry-Pérot interferometer includes notations for photoelectric quadratures.
-We define the classical distances $L$, $l_{x}$, and $l_{y}$, along with the small quantum displacements represented by $\hat{X}_{XEM}$, $\hat{X}_{XITM}$, $\hat{X}_{YEM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ for the end mirrors (EMs) and the intermediate mirrors (ITMs).
+We define the classical distances $L$, $l_{x}$, and $l_{y}$, along with the small quantum displacements represented by $\hat{X}_{XEM}$, $\hat{X}_{XITM}$, $\hat{X}_{YEM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ for the end test masses (EMs) and the input test masses (ITMs).
 Additionally, the quadratures for the laser are denoted by $\hat{d}$ and $\hat{a}$ as the input quadratures to the interferometer, while $\hat{b}$ represents the output quadrature from the interferometer.
@@
 We need to emphasize that we do not discuss the power recycling, the signal recycling, or the squeezed state input techniques [10, 11, 27, 28, 29] in this paper.
-Our focus is solely on re-evaluating a simple Fabri-Pérot gravitational-wave detector.
+Our focus is solely on re-evaluating a simple Fabry-Pérot gravitational-wave detector.
 Therefore, within this paper, we cannot discuss the recent findings regarding the “violation of the standard quantum limit” in LIGO gravitational-wave detectors [12, 13].
-However, we expect that the arguments presented in this paper can be extended to include techniques such as power recycling, signal recycling, or squeezed state input techniques.
+However, the main purpose of this paper is to point out two facts.
+One is the fact that the quantum uncertainty due to the motion of test masses does not affect the noise spectral density.
+The other is the fact the additional classical-carrier-field leakage due to the imperfect tuning of the interferometer at the equilibrium point of the pendulum affect the noise spectral density as an imperfection of the gravitational-wave detectors.
+The realization of the currently operating gravitational-wave detectors is beyond the current scope of this paper.
+Furthermore, we expect that the arguments presented in this paper can be extended to include techniques such as power recycling, signal recycling, and squeezed-state input.
 In this regard, we are confident that the arguments in this paper are meaningful.
@@
 III and VI.
-In this section, the resulting noise spectral density shows that the shot noise in the high-frequency region does not decrease even if the injected laser power is increased.
+In this section, the resulting noise spectral density shows that the shot-noise contribution to the high-frequency region of the signal-referred noise spectral density does not decrease even as the injected laser power is increases.
 In Sec.
@@
 $\hat{E}_{a}^{(+)}(t-z)=\int_{0}^{\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\hat{a}(\omega)e^{-i\omega(t-z)},$ (3)
-where $\hat{a}(\omega)$ is the photon annihilation operator associated with the electric field $\hat{E}_{a}(t-z)$, which satisfies the commutation relation
+where $\hat{a}(\omega)$ is the photon annihilation operator associated with the electric field $\hat{E}_{a}(t-z)$, which satisfies the commutation relations
 $\left[\hat{a}(\omega),\hat{a}^{\dagger}(\omega^{\prime})\right]=2\pi\delta(\omega-\omega^{\prime}),$ (4)
@@
 $\cal A$} is the cross-sectional area of the optical beam.
-To discuss the input-output relation of the interferometer, based on one-photon formulation, it is convenient to introduce the operator $\hat{A}(\omega)$ defined by
+To discuss the input-output relation of the interferometer, based on one-photon formulation.
+Although the two-photon formulation [35, 36] is widely used in the gravitational-wave community, this formulation includes some approximations.
+Since we also want to examine the validity of the two-photon formulation itself in this paper, we derive the input-output relation through the one-photon formulation starting from the point without any approximations.
+In the one-photon formulation, it is convenient to introduce the operator $\hat{A}(\omega)$ defined by
 $\hat{A}(\omega):=\hat{a}(\omega)\Theta(\omega)+\hat{a}^{\dagger}(-\omega)\Theta(-\omega)$ (6)
@@
 The subscription “aa” of the function $\Delta_{a}(t-t^{\prime})$ indicates that this is the vacuum fluctuation originating from the electric field $\hat{E}_{a}$ with the quadrature $\hat{a}(\omega)$.
-We note that the function $\Delta_{a}(t-t^{\prime})$ has an ultraviolet divergence.
+We note that the function $\Delta_{a}(t-t^{\prime})$ has an ultraviolet divergence in its integration when $\omega\rightarrow\infty$.
 However, in the actual measurements of the time sequence of the variables, the time in a measurement is discrete with a finite time bin.
 This time bin gives the maximum frequency $\omega_{\max}$, which becomes the natural ultraviolet cut-off of the frequency in the obtained data.
+If the upper frequency of the signal of our interest is 20kHz, $\omega_{\max}$ must be larger than $2\pi\times 20\times 10^{3}$Hz.
 Incidentally, in the actual measurements of the time sequence of the variables, the whole measurement time is also finite.
 It gives the minimum frequency $\omega_{\min}$ which corresponds to a natural infrared cut-off in frequency.
+If the lower frequency of the signal of our interest is 20Hz, $\omega_{\min}$ must be smaller than $2\pi\times 20$Hz.
 Therefore, we may regard that the integration range over $\omega$ in the definition of the function $\Delta_{a}(t-t^{\prime})$ in Eq.
@@
 (1)–(3).
-One of the quantum states of the electric fields which considered within this paper is the vacuum state $|0\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
-$\hat{a}(\omega)|0\rangle:=0\quad{}^{\forall}\omega>0.$ (14)
+One of the quantum states of the electric fields which is considered within this paper is the vacuum state $|0\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
+$\hat{a}(\omega)|0\rangle_{a}:=0\quad{}^{\forall}\omega>0.$ (14)
 On the other hand, we also consider the coherent state $|\alpha\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
@@
 As well-known, the coherent state $|\alpha\rangle_{a}$ and the vacuum state $|0\rangle_{a}$ are related through the displacement operator $D_{a}[\alpha]$ as [10]
-$|\alpha\rangle_{a}$ $=$ $D_{a}[\alpha]|0\rangle_{a},$ (16)
-$D_{a}[\alpha]$ $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)\hat{a}^{\dagger}(\omega)-\alpha^{*}(\omega)\hat{a}(\omega)\right)\right].$
-Here, we note that the subscription “aa” in the state $|0\rangle_{a}$ and $|\alpha\rangle_{a}$ indicate that these states are associated with the electric field operator $\hat{E}_{a}(t)$, with the quadrature $\hat{a}(\omega)$.
+$|\alpha\rangle_{a}$ $=$ ${{\mathfrak{D}}}_{a}[\alpha]|0\rangle_{a},$ (16)
+${{\mathfrak{D}}}_{a}[\alpha]$ $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)\hat{a}^{\dagger}(\omega)-\alpha^{*}(\omega)\hat{a}(\omega)\right)\right].$
+Here, we note that the subscription “aa” in the state $|0\rangle_{a}$ and $|\alpha\rangle_{a}$ indicate that these states are associated with the electric field operator $\hat{E}_{a}(t)$ with the quadrature $\hat{a}(\omega)$.
 II.2 Multi-mode number and power operators
@@
 $+O\left(\left|{{\mathfrak{B}}}\right|^{0}\right).$
-Here, we ignore the remaining term $O\left(\left|{{\mathfrak{B}}}\right|^{0}\right)$. The time-domain version of this signal operator is given by
+Since we consider the situation where $|{{\mathfrak{B}}}|$ is so large that we can distinguish the effects due to the different order of $|{{\mathfrak{B}}}|$, we ignore the remaining term $O\left(\left|{{\mathfrak{B}}}\right|^{0}\right)$ because we want to measure the $O(|{{\mathfrak{B}}}|^{1})$ effect in the expectation value (26).
+The time-domain version of this signal operator is given by
 $\hat{s}_{N_{b}}(t):=\frac{1}{\kappa_{n}}\hat{N}_{b}(t)-\omega_{0}|{{\mathfrak{B}}}|^{2}.$ (27)
@@
 $f(a)$ $:=$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{-T/2}^{T/2}dte^{-iat}$ (33)
-$=$ $\left\{\begin{array}[]{ccccc}1&&\mbox{for}&&a=0,\\ 0&&\mbox{for}&&a\neq 0,\end{array}\right$. (36)
+$=$ $\left\{\begin{array}[]{ccccc}1&\quad&\mbox{for}&\quad&a=0,\\ 0&\quad&\mbox{for}&\quad&a\neq 0,\end{array}\right$. (36)
 as discussed in Ref. [21]. We also note that for a finite function $g(a)$,
@@
 $\int_{-\infty}^{+\infty}\delta(b)g(b)f(a)db=f(a)g(0).$ (38)
-From the input-output relation which will be derived in Sec. VI, we can confirm the following expectation values:
+Although the final input-output relation will be derived in Sec.
+VI, we compare the final input-output relation (189) and Eqs.
+(19) and (20) above.
+Then, we obtain the expression of the operator $\hat{b}_{n}(\omega)$, more precisely ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$.
+Through this obtained expression of $\hat{b}_{n}(\omega)$, we can confirm the following expectation values:
 $\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{0}-\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}),$ (39)
@@
 [21].
-IV Input-Output relation of the Fabri-Pérot interferometer
-In this section, we derive the input-output relation of the Fabri-Pérot interferometer, which is used in the setup of gravitational-wave detectors.
-The Fabri-Pérot interferometer consists of the beam splitter (BS), intermediate masses (ITMs), and the end mirrors (EMs).
-The injected laser to the Fabri-Pérot interferometer is separated into the directions of the $x$- and the $y$-arms.
+IV Input-Output relation of the Fabry-Pérot interferometer
+In this section, we derive the input-output relation of the Fabry-Pérot interferometer, which is used in the setup of gravitational-wave detectors.
+The Fabry-Pérot interferometer consists of the Beam Splitter (BS), Input Test Masses (ITMs), and the End test Masses (EMs).
+The injected laser to the Fabry-Pérot interferometer is separated into the directions of the $x$- and the $y$-arms.
 The laser is amplified between the ITM and EM.
 In gravitational-wave detectors, BS, ITMs, and EMs are suspended through the vibration-isolation system to measure the relative positions of these mirrors precisely.
-In these relative positions, gravitational-wave signals are included.
+In these relative positions, gravitational-wave signals are included through the equation of mirrors’ motion as extensively discussed in Sec.
+V.
 IV.1 Mirror Displacements
@@
 1, we denote the coordinate values in the proper reference frame of ITM (XITM) and EM (XEM) along $x$-arm by $l_{x}+\hat{X}_{XITM}$ and $l_{x}+L+\hat{X}_{XEM}$, respectively.
-Here, $l_{x}$ is the distance between BS and XITM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals.
-$L$ is the distance between XITM and XEM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals.
+Here, $l_{x}$ is the distance between BS and XITM, and $L$ is the distance between XITM and XEM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals.
 We regard $l_{x}$ and $L$ as classical distances.
-In addition to the $l_{x}$ and $L$, we introduce the quantum displacement $\hat{X}_{XITM}$ and $\hat{X}_{EM}$, which are induced by the injected laser and other forces, including the gravitational-wave signal.
-In this paper, we regard $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ as quantum operators that describe the quantum mirror positions.
+In addition to the $l_{x}$ and $L$, we introduce the quantum displacement $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$, which are induced by the injected laser and other forces, including the gravitational-wave signal.
+In this paper, we regard $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$ as quantum operators that describe the quantum mirror positions.
 In Sec.
-V, the operators $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ are determined as the solution to the Heisenberg equation of motion.
+V, the operators $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$ are determined as the solution to the Heisenberg equation of motion.
 Similarly, we also introduce the coordinate values in the proper reference frame of ITM (YITM) and EM (YEM) along $y$-arm $l_{y}+\hat{X}_{YITM}$ and $l_{y}+L+\hat{X}_{YEM}$, respectively.
@@
 V, the operators $\hat{X}_{YITM}$ and $\hat{X}_{YEM}$ are determined as the solution to the Heisenberg equation of motion.
-Based on the above setup, we consider the electric field of the laser through quantum electrodynamics.
+Based on the above setup, we calculate the laser’s electric field using quantum electrodynamics.
 IV.2 Beam Splitter Junction
@@
 Here, we defined $l_{x}=:c\tau_{x}^{\prime}$ and $l_{y}=:c\tau_{y}^{\prime}$.
-In the Fourier space, we Eq.
+In the Fourier space, Eq.
 (62) is given by
@@
 IV.3.2 Junction at ITMs and cavity propagation
-Next, we consider the arm propagation of the laser along each arm. We denote the notation of the electric field of the electric field as depicted in Fig. 2.
+Next, we consider the laser’s propagation along each arm. We denote the notation of the electric field of the electric field as depicted in Fig. 2.
 Figure 2: Arm propagation in the Fabry-Pérot interferometer and photoelectric quadratures notations.
 The quadratures $\hat{f}$ and $\hat{f}^{\prime}$ are those of the laser incident from the BS to ITM and the laser from ITM to BS, respectively.
-The $\hat{g}$ and $\hat{g}^{\prime}$ are the quadratures for the lasers from ITM to the Fabri-Pérot cavity and from the Fabri-Pérot cavity to ITM, respectively.
+The $\hat{g}$ and $\hat{g}^{\prime}$ are the quadratures for the lasers from ITM to the Fabry-Pérot cavity and from the Fabry-Pérot cavity to ITM, respectively.
 The quadratures $\hat{j}$ and $\hat{j}^{\prime}$ are quadratures for lasers that reach the EM and are reflected by the EM, respectively.
@@
 $\hat{E}_{j^{\prime}_{x,y}}(t)=\hat{E}_{j_{x,y}}(t).$ (77)
-If we have to consider the loss model of the imperfection of EM as in Ref. [8], we have to change Eq. (77).
+If we have to consider the loss model of the imperfection of EM as in Ref. [10], we have to change Eq. (77) as described in Ref. [10].
 Furthermore, the derivations of the electric field relations between $\hat{E}_{g_{x,y}}$ , $\hat{E}_{j_{x,y}}$, $\hat{E}_{j^{\prime}_{x,y}}$, and $\hat{E}_{g^{\prime}_{x,y}}$ are delicate.
-However, the physical effects are just free propagation of the laser.
+However, the physical effects are just the free propagation of the laser.
 This is because the time-dependence of the relative positions $\hat{X}_{x}(t):=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}(t):=\hat{X}_{YEM}-\hat{X}_{YITM}$.
-We have to be careful to treat this time dependence.
-However, we apply the perfect reflection of EMs and the relation of $\hat{E}_{g_{x,y}}$ and $\hat{E}_{g^{\prime}_{x,y}}$ should be identical to the Michelson interferometer from the setup depicted in Fig.
+We have to be careful to treat this time dependence correctly to estimate the retarded effects on the electric fields in the interferometer.
+Since we apply the perfect reflection of EMs, the relation of $\hat{E}_{g_{x,y}}$ and $\hat{E}_{g^{\prime}_{x,y}}$ is identical to the case of the Michelson interferometer from the setup depicted in Fig.
 2.
@@
 $\hat{E}_{g^{\prime}_{y}}(t)$ $=$ $\hat{E}_{g_{y}}\left[t-2\left(\tau+\frac{1}{c}\hat{X}_{y}(t-\tau)\right)\right],$ (79)
-where $\tau=L/c$. Furthermore, through Eqs. (74), (75), (78), and (79), we obtain the relations
+where $\tau=L/c$. Moreover, through Eqs. (74), (75), (78), and (79), we obtain the relations
 $\hat{E}_{f^{\prime}_{x}}(t)$ $=$ $-\sqrt{1-T}\hat{E}_{f_{x}}(t)$ (80)
@@
 (82)).
-To obtain the expressions of $\hat{E}_{g_{x,y}}$ in terms of $\hat{E}_{f_{x,y}}$, the treatment of the Fourier transformation is convenient.
+To obtain the expressions of $\hat{E}_{g_{x,y}}$ in terms of $\hat{E}_{f_{x,y}}$, the treatment of the Fourier transformation is convenient, since we only consider the linear perturbation level with respect to $\hat{X}_{x,y}$ within this paper.
 First, we consider the Fourier transformation of Eq.
@@
 $\hat{X}_{x,y}(t)$ $=:$ $\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\hat{Z}_{x,y}(\omega_{2})e^{-i\omega_{2}t},$ (84)
-we only considered the effects of the linear level of $\hat{X}_{x}$ in Eq. (80). Furthermore, through the inverse Fourier transformation, we obtain
+we only consider the effects of the linear level of $\hat{X}_{x}$ in Eq.
+(80), since we regard that the higher-order of $\hat{X}_{x}$ are negligible.
+Furthermore, through the inverse Fourier transformation, we obtain
 $\hat{F}^{\prime}_{x}(\omega)$ $=$ $-\sqrt{1-T}\hat{F}_{x}(\omega)+\sqrt{T}\hat{G}_{x}(\omega)e^{+2i\omega\tau}$ (85)
@@
 $+O\left(\left(\hat{X}\right)^{2}\right).$
-The last integral in Eq.
-(94) which includes $\hat{Z}_{YITM}$ and $\hat{Z}_{XITM}$ are the retarded effects during the propagation of the laser from the intermediate masses to the beam splitter due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively.
-The third integral in Eq.
-(94) which includes $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{XEM}-\hat{X}_{XITM}$, is also the retarded effect during the propagation of the laser in the Fabri-Pérot cavities.
-The second integral in Eq.
-(94) are the retarded effect during the propagation from the beam splitter to the intermediate masses due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively.
 The first line is the direct shot noise of this interferometer.
+The integral in the second and third lines in Eq.
+(94) is the retarded effect during the propagation from the beam splitter to the input test masses due to the modification of the proper distance between the beam splitter and the input test masses $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively.
+We note that these second and third lines do not include the propagation of the Fabry-Perot cavities between the ITMs and EMs, respectively.
+The integral in the fourth and fifth line in Eq.
+(94), which includes $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{XEM}-\hat{X}_{XITM}$, is also the retarded effect during the propagation of the laser in the Fabry-Pérot cavities.
+The integral in the sixth and seventh line of Eq.
+(94), which includes $\hat{Z}_{YITM}$ and $\hat{Z}_{XITM}$, are the retarded effects during the propagation of the laser from the input test masses to the beam splitter due to the modification of the proper distance between the beam splitter and the input test masses $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively.
+The difference of these terms from the second and third lines in Eq.
+(94) is the fact that the sixth and seventh lines of Eq.
+(94) include the information of the propagation in the Fabry-Pérot cavities.
 The last line in Eq.
@@
 $\hat{Z}_{y}(\omega)$ $=:$ $\hat{Z}_{com}(\omega)-\hat{Z}_{diff}(\omega),$ (96)
-where $\hat{Z}_{com}(\omega)$ represents the $x$-arm and $y$-arm common motion of the relative motion between the end-mirror and the intermediate mass and $\hat{Z}_{dif}(\omega)$ represents the $x$-arm and $y$-arm differential motion of the relative motion the end-mirror and intermediate mass.
+where $\hat{Z}_{com}(\omega)$ represents the $x$-arm and $y$-arm common motion of the relative motion between the end test mass and the input test mass and $\hat{Z}_{diff}(\omega)$ represents the $x$-arm and $y$-arm differential motion of the relative motion the end test mass and input test mass.
 Furthermore, we also define $\hat{Z}_{comITM,diffITM}(\omega)$ by
@@
 (99) implies that the output operator $\hat{B}$ is given by the operators $\hat{A}$, $\hat{D}$, $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$.
-Later, we see that the displacements $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$ are given by $\hat{A}$ and $\hat{D}$ together with the gravitational-wave signal through the equations of motions for end-mirrors and the intermediate mirrors.
+Later, we see that the displacements $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$ are given by $\hat{A}$ and $\hat{D}$ together with the gravitational-wave signal through the equations of motions for end test masses and the input test masses.
 Therefore, to discuss the information from the output operator $\hat{B}$, we have to specify the quantum states associated with the operators $\hat{A}$ and $\hat{D}$.
 The state associated with the operator $\hat{A}$ is the state of the electric field that is injected from the anti-symmetric port.
-At this anti-symmetric port, the photo-detector is located as in Fig.
+At this anti-symmetric port, the photodetector is located as in Fig.
 1.
@@
 Then, the total state $|\mbox{in}\rangle$ of the photon is given by the direct product of the photon states of each frequency as
-$|\mbox{in}\rangle$ $=$ $\prod_{\omega}|\alpha(\omega)\rangle_{d}\otimes|0\rangle_{a}=\prod_{\omega}D_{d}[\alpha(\omega)]|0\rangle_{d}\otimes|0\rangle_{a}$ (100)
-$=:$ $D_{d}|0\rangle_{d}\otimes|0\rangle_{a},$
-$D_{d}$ $:=$ $\prod_{\omega}D_{d}(\alpha(\omega))$ (101)
+$|\mbox{in}\rangle$ $=$ $\prod_{\omega}|\alpha(\omega)\rangle_{d}\otimes|0\rangle_{a}=\prod_{\omega}{{\mathfrak{D}}}_{d}[\alpha(\omega)]|0\rangle_{d}\otimes|0\rangle_{a}$ (100)
+$=:$ ${{\mathfrak{D}}}_{d}|0\rangle_{d}\otimes|0\rangle_{a},$
+${{\mathfrak{D}}}_{d}$ $:=$ $\prod_{\omega}{{\mathfrak{D}}}_{d}(\alpha(\omega))$ (101)
 $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)d^{\dagger}(\omega)-\alpha^{*}(\omega)d(\omega)\right)\right].$
+These states are the reasonable realization of electric field operators that represent the injection of the coherent state laser from the light source, and no input sources except for vacuum fluctuations from the output port.
+We also note that the unitarity of the operation of the beam splitter on the electric field is guaranteed by the injection of the vacuum state associated with the operator $\hat{a}$ from the output port.
+Therefore, these setups are essential to the interferometer’s arguments.
 In the Heisenberg picture, the operator $\hat{d}$ is replaced as
-$D_{d}^{\dagger}\hat{d}(\omega)D_{d}=\hat{d}(\omega)+\alpha(\omega),$ (102)
-$D_{d}^{\dagger}\hat{d}^{\dagger}(\omega)D_{d}=\hat{d}^{\dagger}(\omega)+\alpha^{*}(\omega),$ (103)
-through the displacement operator $D_{d}$ defined by Eq. (101). Then, the operator $\hat{D}(\omega)$ defined by Eq. (61) is transformed as
-$D_{d}^{\dagger}\hat{D}(\omega)D_{d}=\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega),$ (104)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{d}(\omega){{\mathfrak{D}}}_{d}=\hat{d}(\omega)+\alpha(\omega),$ (102)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{d}^{\dagger}(\omega){{\mathfrak{D}}}_{d}=\hat{d}^{\dagger}(\omega)+\alpha^{*}(\omega),$ (103)
+through the displacement operator ${{\mathfrak{D}}}_{d}$ defined by Eq. (101). Then, the operator $\hat{D}(\omega)$ defined by Eq. (61) is transformed as
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{D}(\omega){{\mathfrak{D}}}_{d}=\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega),$ (104)
 where
@@
 $\hat{D}_{v}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega).$ (106)
-To evaluate the output signal expectation value and its fluctuations from the input-output relation (99) under the coherent state of the quadrature $\hat{D}(\omega)$, the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ is useful instead of $\hat{B}(\omega)$, because we treat operators in the Heisenberg picture.
-Furthermore, we regard that $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$ are small correction due to the radiation-pressure noise and gravitational-wave signals.
-Since $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$ also describe small fluctuations, we neglect the quadratic terms $\hat{D}_{v}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ and $\hat{A}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ in the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$.
-Operating $D_{d}^{\dagger}$ and $D_{d}$ to Eq.
+To evaluate the output signal expectation value and its fluctuations from the input-output relation (99) under the coherent state of the quadrature $\hat{D}(\omega)$, the expression of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$ is useful instead of $\hat{B}(\omega)$, because we treat operators in the Heisenberg picture.
+Furthermore, we regard that ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ are small correction due to the radiation-pressure noise and gravitational-wave signals.
+Since $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$ also describe small fluctuations, we neglect the quadratic terms $\hat{D}_{v}(\omega){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{*}(\Omega){{\mathfrak{D}}}_{d}$ and $\hat{A}(\omega){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{*}(\Omega){{\mathfrak{D}}}_{d}$ in the expression of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$.
+Operating ${{\mathfrak{D}}}_{d}^{\dagger}$ and ${{\mathfrak{D}}}_{d}$ to Eq.
 (99), substituting Eqs.
-(104) into this $D_{d}^{\dagger}$-$D_{d}$-operated version of Eq.
+(104) into this ${{\mathfrak{D}}}_{d}^{\dagger}$-${{\mathfrak{D}}}_{d}$-operated version of Eq.
 (99), and applying the above approximation, then, we obtain the input-output relation as
-$\!\!\!\!\!\!D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ (107)
+$\!\!\!\!\!\!{{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$ (107)
 $=$ $i\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))\hat{D}_{c}(\omega)$
@@
 $+\frac{i}{c}e^{+2i\omega\tau}e^{+i\omega(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i\omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\omega_{1}\hat{D}_{c}(\omega_{1})$
-$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
+$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
 $+i\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i\omega\tau}\right]^{-1}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+i(\omega+\omega_{1})\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}\hat{D}_{c}(\omega_{1})$
-$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\omega-\omega_{1})D_{d}\right]$
+$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
 $+\frac{i}{c}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\omega_{1}\sqrt{\frac{|\omega_{1}|}{|\omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{1}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}e^{+2i\omega_{1}\tau}e^{+i(\omega+\omega_{1})\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}\hat{D}_{c}(\omega_{1})$
-$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
+$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
 […]
@@
 (107) is the leakage of the classical carrier field due to the phase offset $\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})$.
-The terms in the second line are the vacuum fluctuations which correspond to the shot noise in the conventional input-output relations in Refs.
+The terms in the second line are the vacuum fluctuations that correspond to the shot noise in the conventional input-output relations in Refs.
 [10].
-The terms in the third- and fourth-lines are the response of the intermediate mirror motion which includes the radiation pressure noise through the motions of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$.
-The terms in the fifth-, sixth-lines are the response of the relative motion of the end mirrors and the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$.
-Finally, the term in the seventh- and eighth-lines are response of the intermediate mirror motion which includes the radiation pressure noise through the motion of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$.
-These come from the retarded effects which was explained after Eq.
+The terms in the third and fourth lines are the response of the input-test-mass motion, which includes the radiation pressure noise through the motions of the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$.
+The terms in the fifth and sixth lines are the response of the relative motion of the end test masses and the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$.
+Finally, the terms in the seventh and eighth lines are responses of the input-test-mass motion, which includes the radiation pressure noise through the motion of the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$.
+These come from the retarded effects which were explained after Eq.
 (94).
@@
 The input-output relation (107) is one of the main results of this paper.
-To conduct the actual evaluation of the input-output relation (107), we have to evaluate $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in some way.
-In the case of gravitational-wave detectors, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ are evaluated through the equations of motions for the end-mirrors as discussed in the next section.
+To conduct the actual evaluation of the input-output relation (107), we have to evaluate ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ in some way.
+In the case of gravitational-wave detectors, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ are evaluated through the equations of motions for the end test masses as discussed in the next section.
 IV.5.1 Monochromatic coherent amplitude
-Before the evaluations of the mirror displacements, we consider the coherent state of the incident laser.
+Before evaluating the mirror displacements, we consider the coherent state of the incident laser.
 In the conventional gravitational-wave detectors, the state of the optical beam from the light source is in the single-mode coherent state with the complex amplitude
 $\alpha(\omega)=2\pi N\delta(\omega-\omega_{0}).$ (108)
+Here, $N$ is given by
+$N=\sqrt{\frac{I_{0}}{\hbar\omega_{0}}},$ (109)
+where $I_{0}$ is the averaged classical power of the laser defined by Eq.
+(114) below.
 We note that $\alpha(\omega)$ is real.
@@
 Then, we have obtain
-$D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{E}_{d}(t){{\mathfrak{D}}}_{d}$
 $=$ $\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}\sqrt{\frac{2\pi\hbar|\omega|}{{\cal A}c}}\left(\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega)\right)e^{-i\omega t}.$
@@
 $=$ $2\pi N\left\{\delta(\omega-\omega_{0})\Theta(\omega)+\delta(\omega+\omega_{0})\Theta(-\omega)\right\}.$
-The quantum expectation value $I$ of the power of the electric field $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$ is given by
+The quantum expectation value $I$ of the power of the electric field ${{\mathfrak{D}}}_{d}^{\dagger}\hat{E}_{d}(t){{\mathfrak{D}}}_{d}$ is given by
 $I=N^{2}\hbar\omega_{0}(1+\cos(2\omega_{0}t))+\frac{\hbar}{2}\int_{0}^{+\infty}\frac{d\omega}{2\pi}\omega.$ (113)
 The first term in Eq.
-(112) is the contribution from the classical carrier field $\alpha(\omega)$, and the second term comes from the vacuum fluctuations of the electric field operator $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$.
+(113) is the contribution from the classical carrier field $\alpha(\omega)$, and the second term comes from the vacuum fluctuations of the electric field operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{E}_{d}(t){{\mathfrak{D}}}_{d}$.
 The second term diverges as well-known.
@@
 $=$ $N^{2}\hbar\omega_{0}.$
-Then, we denote the factor $N$ in the function $\alpha(\omega)$ is given by
-$N=\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}.$ (114)
+Thus, we denote the factor $N$ in the function $\alpha(\omega)$ is given by Eq. (109).
 IV.5.2 Final Input-output relation with mirrors’ motion
@@
 Furthermore, we consider the sideband picture $\omega=\omega_{0}\pm\Omega$ of the input-output relation under this monochromatic condition (IV.5.1).
-In this sideband picture, the terms include $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, or $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ appears in the expression of the input-output relation.
-Here, we neglect these terms because their behavior as
+In this sideband picture, the terms include $\delta(2\omega_{0}\pm\Omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$, or ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ appears in the expression of the input-output relation.
+Here, we neglect these terms because they behave as
 $\int_{-\infty}^{+\infty}\frac{d\Omega}{2\pi}f(2\omega_{0}+\Omega)e^{-i(\Omega t}$
@@
 $=$ $e^{+2i\omega_{0}t}\int_{-\infty}^{+\infty}\frac{d\omega^{\prime}}{2\pi}f(\omega^{\prime})e^{+i\omega^{\prime}t}.$ (116)
-These show that the terms of $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ have the behavior of the rapid oscillation with the frequency $2\omega_{0}$ in the time domain.
+These show that the terms of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ have the behavior of the rapid oscillation with the frequency $2\omega_{0}$ in the time domain.
 Since we concentrate only on the behavior of the sideband frequency $\Omega$, these rapid oscillation terms are outside of the frequency range of interest.
-For these reasons, we ignore the terms $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ as an approximation [26].
+For these reasons, we ignore the terms $\delta(2\omega_{0}\pm\Omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ as an approximation [31].
 Applying this approximation, the input-output relation (107) is given by
-$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{D}$ (117)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (117)
 $=$ $N\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+2i\omega_{0}\tau}e^{+i\omega_{0}(\tau_{y}^{\prime}+\tau_{x}^{\prime})}i\sin(\omega_{0}(\tau_{y}^{\prime}-\tau_{x}^{\prime}))2\pi\delta(\Omega)$
@@
 $+Nie^{+2i(\omega_{0}\pm\Omega)\tau}e^{+i(\omega_{0}\pm\Omega)(\tau_{y}^{\prime}+\tau_{x}^{\prime})}\left[1-\sqrt{1-T}e^{-2i(\omega_{0}\pm\Omega)\tau}\right]\left[1-\sqrt{1-T}e^{+2i(\omega_{0}\pm\Omega)\tau}\right]^{-1}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\frac{\omega_{0}}{c}$
-$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
+$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
 $+Ni\frac{2}{c}T\left[1-\sqrt{1-T}e^{+2i(\omega_{0}\pm\Omega)\tau}\right]^{-1}\omega_{0}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+i(2\omega_{0}\pm\Omega)\tau}e^{+i(2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
-$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
+$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
 $+Ni\frac{\omega_{0}}{c}\sqrt{\frac{|\omega_{0}|}{|\omega_{0}\pm\Omega|}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}e^{+2i\omega_{0}\tau}e^{+i(2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}+\tau_{x}^{\prime}}{2}}$
-$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
+$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"}.$
-To complete the evaluation of this input-output relation (117), we have to specify the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators.
+To complete the evaluation of this input-output relation (117), we have to specify the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators.
 These are determined through the Heisenberg equation of motion in the next section.
 V Equations for mirrors’ motions and their solutions
-The purpose of this section is the evaluation of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators.
+The purpose of this section is the evaluation of the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators.
 To discuss this evaluation within quantum mechanics, we first discuss a quantum mechanical model of a forced harmonic oscillator in Sec.
@@
 In Sec.
-V.4, we derived the explicit form of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators based on the arguments in Sec.
+V.4, we derived the explicit form of the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators based on the arguments in Sec.
 V.1.
-The results of this section are used to evaluate the input-output relation of the Fabri-Pérot interferometer in Sec.
+The results of this section are used to evaluate the input-output relation of the Fabry-Pérot interferometer in Sec.
 VI.
@@
 In this subsection, we consider the quantum mechanical Hamiltonian $\hat{H}$ for a forced harmonic oscillator with the position operator $\hat{X}(t)$ and the momentum operator $\hat{P}(t)$ with external quantum force $\hat{F}(t)$:
-$\hat{H}=\frac{\hat{P}^{2}}{2\mu}+\frac{1}{2}\mu m_{p}^{2}\hat{X}^{2}-\hat{F}(t)\hat{X}$. (118)
+$\hat{H}=\frac{\hat{P}^{2}}{2\mu}+\frac{1}{2}\mu\omega_{p}^{2}\hat{X}^{2}-\hat{F}(t)\hat{X}$. (118)
 The position operator $\hat{X}$ corresponds to the mirror displacement operators $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$.
-$\omega_{p}$ corresponds to the pendulum fundamental frequency of mirrors.
+$\mu$ and $\omega_{p}$ correspond to the mass and the pendulum fundamental frequency of mirrors, respectively.
 The operators $\hat{X}(t)$ and $\hat{P}(t)$ satisfy the canonical commutation relation
@@
 $\left[\hat{X},\hat{F}(t)\right]=\left[\hat{P},\hat{F}(t)\right]=0.$ (120)
+This assumption (120) is valid for the mirrors’ motion if we ignore the optical spring effects in Appendix B due to the appropriate tuning conditions as discussed in Sec.
+VIII.
 From the Hamiltonian (118), Heisenberg’s equations of motion are given by
@@
 V.2 Equations of motion for each mirror
-Now, we consider the equation of motion for EMs and ITMs to evaluate the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$, $D_{d}\hat{Z}_{com}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in the input-output relation (117).
-We assume the masses of ITMs are equal to $m_{ITM}$ and those of EMs are equal to $m_{EM}$.
+Now, we consider the equation of motion for EMs and ITMs to evaluate the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ in the input-output relation (117).
+We assume the masses of ITMs are equal to $m_{ITM}$, and those of EMs are equal to $m_{EM}$.
 Since we apply the proper reference frame whose center is BS, the displacement of BS vanishes, which means that we have to control the external forces to BS vanish.
 Furthermore, we only consider the motions of mirrors are restricted so that XITM and XEM move only along the $x$-direction and YITM and YEM move only along the $y$-direction and $x$-direction and $y$-direction is orthogonal.
-In the proper reference frame [25], the effects of gravitational waves are represented by the components $R_{txtx}$, $R_{tyty}$ of the Riemann curvature.
+In the proper reference frame, the effects of gravitational waves are represented by the components $R_{txtx}$, $R_{tyty}$ of the Riemann curvature tensor [30].
 Moreover, we denote the radiation pressure force to mirrors, XITM, YITM, XEM, and YEM from the laser in the interferometer by $\hat{F}_{rpXITM}$, $\hat{F}_{rpYITM}$, $\hat{F}_{rpXEM}$, $\hat{F}_{rpYEM}$, respectively.
@@
 We also applied the approximation $L,l_{x},l_{y}$ $\gg$ $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{XEM}$, $\hat{X}_{YEM}$.
-As the components of the Riemann tensor, we only consider the plus modes of gravitational waves associated with the $x$- and $y$-directions of the interferometer.
-In this case, the components of $R_{txtx}$ and $R_{tyty}$ the Riemann tensor are given by
+As the components of the Riemann curvature tensor, we only consider the plus modes of gravitational waves associated with the $x$- and $y$-directions of the interferometer.
+In this case, the components of $R_{txtx}$ and $R_{tyty}$ the Riemann curvature tensor are given by
 $R_{txtx}=-\frac{1}{2}\frac{\partial^{2}}{\partial t^{2}}h_{+}=-R_{tyty},\quad h:=h_{+}.$ (140)
@@
 $h(t,z)=:\int_{-\infty}^{+\infty}\frac{d\omega}{2\pi}e^{-i\omega_{2}t}H(\omega,z),$ (151)
-where $z$ is the propagation direction of the laser in the interferometer.
+where $z$ is proper distance to the propagation direction of the laser from BS in the interferometer.
 […]
 $\hat{Z}_{comITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpYITM}(\omega)+{{\mathscr{F}}}_{rpXITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
-$\hat{Z}_{difITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpXITM}(\omega)-{{\mathscr{F}}}_{rpYITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
+$\hat{Z}_{diffITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpXITM}(\omega)-{{\mathscr{F}}}_{rpYITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
 $\hat{Z}_{com}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpy}(\omega)+{{\mathscr{F}}}_{rpx}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$ (154)
-$\hat{Z}_{dif}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpx}(\omega)-{{\mathscr{F}}}_{rpy}(\omega)}{2(\omega^{2}-\omega_{p}^{2})}$ (155)
+$\hat{Z}_{diff}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpx}(\omega)-{{\mathscr{F}}}_{rpy}(\omega)}{2(\omega^{2}-\omega_{p}^{2})}$ (155)
 $+\frac{\omega^{2}}{2(\omega^{2}-\omega_{p}^{2})}H(\omega,L)L.$
-In the input-output relation (117), $\hat{Z}_{dif}(\omega)$, $\hat{Z}_{com}(\omega)$, $\hat{Z}_{comITM}(\omega)$, and $\hat{Z}_{difITM}(\omega)$ do not appear in this expression, but $D_{d}^{\dagger}\hat{Z}_{dif}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{difITM}(\omega)D_{d}$ appear in Eq.
+In the input-output relation (117), $\hat{Z}_{diff}(\omega)$, $\hat{Z}_{com}(\omega)$, $\hat{Z}_{comITM}(\omega)$, and $\hat{Z}_{diffITM}(\omega)$ do not appear in this expression, but ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega){{\mathfrak{D}}}_{d}$ appear in Eq.
 (117).
 Through Eqs.
-(154) and (155), $\hat{Z}_{dif}(\omega)$ and $\hat{Z}_{com}(\omega)$ are linearly related to ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$.
+(154) and (155), $\hat{Z}_{diff}(\omega)$ and $\hat{Z}_{com}(\omega)$ are linearly related to ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$.
 Furthermore, Eqs.
-(LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx) and (LABEL:eq:Eq-of-motion-XYITM-dif-Fourier-approx) yields that $\hat{Z}_{comITM}(\omega)$ and $\hat{Z}_{difITM}(\omega)$ linearly depends on ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$.
-Therefore, we have to evaluate $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$ to the direct evaluation of the input-output relation (117).
+(LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx) and (153) yields that $\hat{Z}_{comITM}(\omega)$ and $\hat{Z}_{diffITM}(\omega)$ linearly depends on ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$.
+Therefore, we have to evaluate ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega){{\mathfrak{D}}}_{d}$ to the direct evaluation of the input-output relation (117).
 These evaluations are discussed in the next subsection.
 V.3 Evaluation of Radiation pressure forces
-Now, we evaluate the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$.
+Now, we evaluate the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega){{\mathfrak{D}}}_{d}$.
 The details of the evaluation of these radiation pressure forces are given in Appendix B.
@@
 (329) and (331), respectively.
-To consider the situation where the state for the incident electric field from the light source in the coherent state with the complex amplitude $\alpha(\omega)$, we apply the displacement operator $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right, which are defined by Eq.
+To consider the situation where the state for the incident electric field from the light source in the coherent state with the complex amplitude $\alpha(\omega)$, we apply the displacement operator ${{\mathfrak{D}}}_{d}^{\dagger}$ from the left and ${{\mathfrak{D}}}_{d}$ from the right, which are defined by Eq.
 (101), to the obtained radiation pressure forces (326), (327), (329), and (331).
@@
 (108).
-Thus, we reach the expressions of the Fourier transformation of the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ as Eqs.
+Thus, we reach the expressions of the Fourier transformation of the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ as Eqs.
 (338), (339), (340), and (341), respectively.
-Here, we consider the tuning condition of the Fabri-Pérot cavity by
+Here, we consider the tuning condition of the Fabry-Pérot cavity by
 $\omega_{0}\tau=2n\pi,\quad n\in{{\mathbb{N}}}.$ (161)
-Due to this tuning condition, the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ given by Eqs.
+Due to this tuning condition, the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ given by Eqs.
 (338), (339), (340), and (341), respectively, are expressed as
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (162)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$ (162)
 $=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
@@
 $\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)\right],$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (163)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$ (163)
 $=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
@@
 $\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})+\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})+\hat{A}(\Omega+\omega_{0})\right)\right],$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (164)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$ (164)
 $=$ $+\frac{N^{2}\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[1-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
@@
 $\quad\quad\times\left[\sqrt{|\omega_{0}(\Omega-\omega_{0})|}\left(\hat{D}_{v}(\Omega-\omega_{0})-\hat{A}(\Omega-\omega_{0})\right)+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\hat{D}_{v}(\Omega+\omega_{0})-\hat{A}(\Omega+\omega_{0})\right)\right],$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$ (165)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$ (165)
 $=$ $+N^{2}\frac{\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[1-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
@@
 (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx)–(155), we can obtain the explicit solution to the Heisenberg equations as follows:
-$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ (166)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ (166)
 $=$ $+\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$
@@
 $\quad\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left(\cos\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})+i\sin\left(\Omega\left(\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right)\right],$
-$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ (167)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ (167)
 $=$ $-\frac{N\hbar}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}e^{+i\Omega\left(\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}\right)}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]$
@@
 $+\frac{\Omega^{2}}{2(\Omega^{2}-\omega_{p}^{2})}H(\Omega,L+l)L,$
-$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ (168)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ (168)
 $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
@@
 $\quad\quad\quad\left.+\sqrt{|\omega_{0}(\Omega+\omega_{0})|}\left[\cos\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{D}_{v}(\Omega+\omega_{0})-i\sin\left(\Omega\left(\frac{\tau_{x}^{\prime}-\tau_{y}^{\prime}}{2}\right)\right)\hat{A}(\Omega+\omega_{0})\right]\right],$
-$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ (169)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ (169)
 $=$ $-\frac{2N\sqrt{1-T}\hbar}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)e^{+i\Omega\left(\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}\right)}$
@@
 Through these approximations, (166)–(169) to the Heisenberg equations for mirrors’ motion [33]
-$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ $=$ $\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$ (172)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$ (172)
 $-\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$
 $\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]\left[\hat{D}_{v}(\Omega-\omega_{0})+\hat{D}_{v}(\Omega+\omega_{0})\right],$
-$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ $=$ $\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$ (173)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$ (173)
 $\quad\quad\times\left[\frac{T}{m_{EM}}+\frac{2\cos(\Omega\tau)\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right]$
 $+\frac{\Omega^{2}}{2(\Omega^{2}-\omega_{p}^{2})}H(\Omega,L+l)L,$
-$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
 $+\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\cos(\Omega\tau)\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\left[\hat{D}_{v}(\Omega-\omega_{0})+\hat{D}_{v}(\Omega+\omega_{0})\right],$
-$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ $=$ $-\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right].$ (175)
-We have to note that the classical constant terms in Eq.
-(172) and (174) diverge if we choose $\omega_{p}=0$, i.e., if we assume the mirrors are in completely free motion.
-This is the reason why we introduced the nonvanishing pendulum fundamental frequency $\omega_{p}$ in each equation of motion for mirrors.
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right].$ (175)
+We have to note that the classical constant terms in Eqs.
+(172) and (LABEL:eq:Eq-of-motion-XYITM-com-tuned-wobifreq-sol-sum-approx) diverge if we choose $\omega_{p}=0$, i.e., if we assume the mirrors are in completely free motion.
+This is the reason why we introduced the nonvanishing pendulum fundamental frequency, $\omega_{p}$, in each equation of motion for the mirrors.
 This introduction of the nonvanishing pendulum fundamental frequency $\omega_{p}$ is a natural consequence from the viewpoint of the actual ground-based gravitational-wave detectors.
@@
 However, we can regard the introduction of this regularization as reasonable for the actual ground-based gravitational-wave detectors.
-In this sense, we have to emphasize that this introduction $\omega_{p}$ has a physical meaning more than a mere mathematical regularization.
+In this sense, we have to emphasize that this introduction of $\omega_{p}$ has a physical meaning more than a mere mathematical regularization.
 […]
-VI Final input-output relation for a Fabri-Pérot Graviational-Wave Detector
+VI Final input-output relation for a Fabry-Pérot Graviational-Wave Detector
 Now, we return to the evaluation of the input-output relation (117). First, we introduce the offset phase $\theta$ by
 $\theta:=\omega_{0}(\tau_{y}^{\prime}-\tau_{x}^{\prime}).$ (176)
+In the DC readout scheme, we dare to set up the interferometer to leak the classical carrier field in the dark port to use this classical carrier field as a reference of the measurement of the small fluctuations in the dark port.
+Within this paper, we control this leakage of the classical carrier field through the control the parameter $\theta$ defined by Eq.
+(176), though there are many different methods to make a leakage of the classical carrier field at the dark port.
 We also apply the approximation $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2\sim 0$ and $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2\sim 0$ as estimated in Eqs.
 (LABEL:eq:Omegatauy-taux-half) and (LABEL:eq:Omegatauy+taux-half).
-Furthermore, we consider the tuning condition of the Fabri-Pérot cavity by Eq.
+Furthermore, we consider the tuning condition of the Fabry-Pérot cavity by Eq.
 (161).
@@
 $\omega_{0}\frac{\tau_{x}^{\prime}+\tau_{y}^{\prime}}{2}=2m\pi,\quad m\in{{\mathbb{N}}}.$ (177)
+Even if we specified the offset angle $\theta$ defined by Eq.
+(176), the average $(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2$ is still free.
+Therefore, we have to tune this average distance to guarantee the output port is nearly dark.
 Moreover, we apply the approximation $\Omega\ll\omega_{0}$.
-$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (178)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (178)
 $=$ $+iN\sin\theta 2\pi\delta(\Omega)$
 $+\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}e^{\pm 2i\Omega\tau}\left[i\sin\theta\hat{D}_{v}(\omega_{0}\pm\Omega)+\cos\theta\hat{A}(\omega_{0}\pm\Omega)\right]$
-$+i\frac{N\omega_{0}}{c}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
-$-i\frac{2NT\omega_{0}}{c}e^{\pm i\Omega\tau}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}+\cos\theta D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
-$+\frac{iN\omega_{0}}{c}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right].$
+$+i\frac{N\omega_{0}}{c}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
+$-i\frac{2NT\omega_{0}}{c}e^{\pm i\Omega\tau}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}+\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
+$+\frac{iN\omega_{0}}{c}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right].$
 Substituting the radiation pressure forces (172)–(175) into Eq. (178), and applying the approximation $L\gg l_{x,y}$, we reached the input-output relation
-$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (179)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (179)
 $=$ $+iN\sin\theta\left[1+i\frac{2N^{2}\hbar\omega_{0}^{2}}{c^{2}\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-4}\left[\frac{T^{2}}{m_{EM}}+\frac{4(1-T)\left[1-\sqrt{1-T}\right]^{2}}{m_{ITM}}\right]\right]2\pi\delta(\Omega)$
@@
 $h_{SQL}$ $:=$ $\sqrt{\frac{4\hbar}{\Omega^{2}L^{2}}}\left[\frac{1}{m_{EM}}+\frac{2\sqrt{1-T}\left[1-\sqrt{1-T}\right]\cos(\Omega\tau)}{T}\left[1-\frac{\left[1-\sqrt{1-T}\right]^{2}}{T}\cos(\Omega\tau)\right]\frac{1}{m_{ITM}}\right]^{1/2}.$ (188)
+Here, “SQL” in $h_{SQL}$ means “Standard Quantum Limit” which is given by Eq.
+(198) below.
+We may regard that the definition of $h_{SQL}$ comes from the inequality Eq.
+(198).
 Through the quadratures (180)–(184), variables $\mathfrak{R}$}}, $\beta$, $\kappa$, and $h_{SQL}$, the input-output relation (179) is given by
-$D_{d}^{\dagger}\hat{b}_{\pm}(\Omega)D_{d}$ $=$ $+i\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\sin\theta\left[1+i{{\mathfrak{R}}}\right]2\pi\delta(\Omega)$ (189)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{\pm}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $+i\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\sin\theta\left[1+i{{\mathfrak{R}}}\right]2\pi\delta(\Omega)$ (189)
 $+e^{\pm 2i\beta}\left[i\sin\theta\hat{d}_{\pm}(\Omega)+\cos\theta\hat{a}_{\pm}(\Omega)\right]$
@@
 From these definitions, we obtain
-$D_{d}^{\dagger}\hat{b}_{1}(\Omega)D_{d}$ $=$ $-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}{{\mathfrak{R}}}\sin\theta 2\pi\delta(\Omega)$ (193)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{1}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}{{\mathfrak{R}}}\sin\theta 2\pi\delta(\Omega)$ (193)
 $+e^{+2i\beta}\left[-\sin\theta\hat{d}_{2}(\Omega)+\cos\theta\hat{a}_{1}(\Omega)\right]$
 $+e^{+2i\beta}\kappa\sin\theta\hat{d}_{1}(\Omega).$
-$D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}$ $=$ $\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)$ (194)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)$ (194)
 $+e^{+2i\beta}\left[\sin\theta\hat{d}_{1}(\Omega)+\cos\theta\hat{a}_{2}(\Omega)\right]$
@@
 In Ref.
-[8], $D_{d}^{\dagger}\hat{b}_{2}D_{d}$ is regarded as the signal operator in gravitational-wave detectors.
+[10], ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}{{\mathfrak{D}}}_{d}$ is regarded as the signal operator in gravitational-wave detectors.
 The first line of Eq.
@@
 The second line of Eq.
-(194) corresponds to the shot noise due to the errors of the photon number estimation in the coherent state of the laser.
+(194) corresponds to the shot noise due to the uncertainties of the photon number estimation in the coherent state of the laser.
 The third line of Eq.
@@
 $\hat{h}_{n}$ $:=$ $e^{-i\beta}\frac{h_{SQL}}{\sqrt{2\kappa}\cos\theta}$ (195)
-$\quad\times\left[\left(D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)\right)\right$.
+$\quad\times\left[\left({{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}(\Omega){{\mathfrak{D}}}_{d}-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)\right)\right$.
 $\quad\quad\quad\left.+e^{+i\beta}\sqrt{2\kappa}\cos\theta\frac{H(\Omega,L)}{h_{SQL}}\right]$
@@
 They only suggested that $S_{(K)}(\Omega)$ is an estimation measure of the amplitude of the quantum noises in the lasers of interferometers.
+However, the main purpose of Ref.
+[10] was to show some techniques to surpass the “standard quantum limit” using the squeezed states techniques and comparison with them, and it was not a serious problem of the definition of the noise spectral density at that time.
+After the publication of Ref.
+[10], general arguments of quantum measurement theory were proposed [14].
+Then, the definitions and the meanings of “noises” began to be discussed seriously.
 Besides the problem whether $S_{(K)}(\Omega)$ defined by Eq.
-(196) is an physically reasonable noise estimation or not, we can calculate $S_{(K)}(\Omega)$ through the definitions (196) of $S_{(K)}(\Omega)$ and the definition (195) of the noise operator $\hat{h}_{n}$.
+(196) is a physically reasonable noise estimation or not, we can calculate $S_{(K)}(\Omega)$ through the definitions (196) of $S_{(K)}(\Omega)$ and the definition (195) of the noise operator $\hat{h}_{n}$.
 As a result of this calculation, we obtain the result
@@
 (206) and (209), respectively.
-This $S_{(K)}^{(K)}(\Omega)$ coincides with the noise spectral density for the Fabri-Pérot interferometer in Ref.
+This $S_{(K)}^{(K)}(\Omega)$ coincides with the noise spectral density for the Fabry-Pérot interferometer in Ref.
 [10].
@@
 […]
-In the next section, we derive the quantum noise spectral density from the quantum fluctuations of photons at the photo-detector through a DC readout scheme, which is discussed in Sec. III.
-VII DC readout scheme for a Fabri-Pérot Graviational-Wave Detector
-Here, we apply the general arguments on the DC readout scheme for a Fabri-Pérot interferometer with the input-output relation (189) or equivalently Eqs.
+In the next section, we derive the quantum noise spectral density from the quantum fluctuations of photons at the photodetector through a DC readout scheme, which is discussed in Sec. III.
+VII DC readout scheme for a Fabry-Pérot Graviational-Wave Detector
+Here, we apply the general arguments on the DC readout scheme for a Fabry-Pérot interferometer with the input-output relation (189) or equivalently Eqs.
 (193) and (194).
@@
 ${{\mathfrak{B}}}$ $=$ $i\sin\theta\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\left[1+i{{\mathfrak{R}}}\right],$ (214)
-$D_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega)D_{d}=D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$ (215)
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}={{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\pm}(\Omega){{\mathfrak{D}}}_{d}$ (215)
 $=$ $e^{\pm 2i\beta}\left[i\sin\theta\hat{d}_{\pm}(\Omega)+\cos\theta\hat{a}_{\pm}(\Omega)\right]$
@@
 In the case of $\theta=0$, the gravitational-wave signal $H(\Omega,L)$ is included in the term of $O((I_{0})^{0})$.
-This is the well-known fact that the complete dark-port condition $\theta=0$ is meaningless in the DC-readout scheme.
+It is well-known that the complete dark-port condition $\theta=0$ is meaningless in the DC-readout scheme.
 The factor $\sin(2\theta)$ in Eq.
@@
 (214) is proportional to $\sin\theta$.
-Thus, we concludes that $\theta=\pi/4$ maximize the expectation value (219).
+Thus, we conclude that $\theta=\pi/4$ maximize the expectation value (219).
 From Eq. (219), we obtain the gravitational-wave signal $H(\Omega,L)$ through the expectation value (219) of the signal operator $\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)$ of the DC-readout by
@@
 Next, we consider the stationary noise spectral density (48).
-From the explicit expression (215) of the output operator $D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$, the amplitude- and phase-quadrature $D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$ defined by Eqs.
+From the explicit expression (215) of the output operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\pm}(\Omega){{\mathfrak{D}}}_{d}$, the amplitude- and phase-quadrature ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n1}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n2}(\Omega){{\mathfrak{D}}}_{d}$ defined by Eqs.
 (45) and (46), respectively, are given by
-$D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n1}(\Omega){{\mathfrak{D}}}_{d}$
 $=$ $e^{+2i\beta}\left[-\sin\theta\hat{d}_{2}(\Omega)+\cos\theta\hat{a}_{1}(\Omega)+\kappa\sin\theta\hat{d}_{1}(\Omega)\right],$
-$D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n2}(\Omega){{\mathfrak{D}}}_{d}$
 $=$ $e^{+2i\beta}\left[\sin\theta\hat{d}_{1}(\Omega)+\cos\theta\hat{a}_{2}(\Omega)-\kappa\cos\theta\hat{a}_{1}(\Omega)\right].$
-From the expressions (LABEL:eq:B1-noise-final) and (LABEL:eq:B2-noise-final) of the amplitude- and phase-quadrature, the operator $D_{d}^{\dagger}\hat{b}_{n\Theta}D_{d}$ defined by Eq.
+From the expressions (LABEL:eq:B1-noise-final) and (LABEL:eq:B2-noise-final) of the amplitude- and phase-quadrature, the operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}{{\mathfrak{D}}}_{d}$ defined by Eq.
 (47) is given by
-$D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega){{\mathfrak{D}}}_{d}$
 $=$ $e^{+2i\beta}\left[\sin\theta\left(-\cos\Theta\hat{d}_{2}(\Omega)+\sin\Theta\hat{d}_{1}(\Omega)\right)\right$.
@@
 We also have
-$D_{d}^{\dagger}\hat{b}_{n\Theta}^{\dagger}(\Omega)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}^{\dagger}(\Omega){{\mathfrak{D}}}_{d}$
 $=$ $e^{-2i\beta}\left[\sin\theta\left(-\cos\Theta\hat{d}_{2}^{\dagger}(\Omega)+\sin\Theta\hat{d}_{1}^{\dagger}(\Omega)\right)\right$.
@@
 From these commutation relations (226)–(229) for the operators $\hat{a}_{n1}$ and $\hat{a}_{n2}$, and corresponding commutation relations for the operators $\hat{d}_{n1}$ and $\hat{d}_{n2}$, we can confirm the commutation relation (49).
-Through the operators $D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$ above, the straightforward calculations for the right-hand side of Eq. (48) yields
+Through the operators ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega){{\mathfrak{D}}}_{d}$ above, the straightforward calculations for the right-hand side of Eq. (48) yields
 $S_{Nn}(\Omega)$ $=$ $\omega_{0}^{2}|{{\mathfrak{B}}}|^{2}\left[1-\kappa\sin(2\Theta)\cos(2\theta)\right$.
@@
 $\epsilon$ is the incomplete parameter discussed Sec.
-VIII.
-Figure 7: The square root of the noise spectral density $\sqrt{S_{H}}/{{\mathfrak{R}}}$ with equal mass $(EQ)$ case where $m_{EM}=m_{ITM}$ and the offset is $\theta=\frac{\pi}{4}$.
-The behavior of this ratio is similar to Kimble’s signal-referred noise spectral density $\sqrt{S_{K}}$.
+VIII, while $\epsilon=0$ corresponds to the fact that the tuning point is the equilibrium point of the pendulum and $\epsilon=1$ corresponds to the tuning without taking into account of the displacement of the equilibrium point of the pendulum due to the classical radiation pressure from the laser.
+(See Sec.
+VIII.)
+Figure 7: The square root ratio of the noise spectral density $S_{H}^{(EQ)}/{{\mathfrak{R}}}^{2}$ and $S_{(K)}^{(K)}$ where $S_{(K)}^{(K)}$ is defined by Eq.
+(212).
+This ratio is approximately unity except in the high-frequency range $\sim 20$ kHz.
+At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$.
+We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$.
 From Eq. (26) with $\Omega\ll\omega_{0}$, Eqs. (213) and (214), the square of the absolute value of Eq. (219) is also given by
@@
 7 to clarify the difference between the derived noise-spectral density $S_{H}$ and the original Kimble’s noise spectral density.
-This ratio is approximately unity except for the high frequency range $\sim 20$kHz.
+This ratio is approximately unity except for the high-frequency range $\sim 20$kHz.
 At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$ defined by Eq.
@@
 Since the expectation value $|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$ given by Eq.
-(219) is the largest when $\theta=\pi/4$, we show Figures 6, 7 and 8 with $\theta=\pi/4$.
-Figure 8: The square root ratio of the noise spectral density $S_{H}^{(EQ)}/{{\mathfrak{R}}}^{2}$ and $S_{(K)}^{(K)}$ where $S_{(K)}^{(K)}$ is defined by Eq.
-(212).
-This ratio is approximately unity except for the high frequency range $\sim 20$kHz.
-At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$.
-We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$.
+(219) is the largest when $\theta=\pi/4$, we show Figures 6, 7, and 8 with $\theta=\pi/4$.
+Figure 8: The square root of the noise spectral density $\sqrt{S_{H}}/{{\mathfrak{R}}}$ with equal mass $(EQ)$ case where $m_{EM}=m_{ITM}$ and the offset is $\theta=\pi/4$.
+The behavior of this ratio is similar to Kimble’s signal-referred noise spectral density $\sqrt{S_{K}}$.
 The factor $\mathfrak{R}$}} in the signal referred noise spectral density $S_{H}(\Omega)$ comes from (214).
@@
 Due to this modification of the classical carrier field appears to the signal referred noise spectral density $S_{H}(\Omega)$, although it does not appear in the expectation value (219) of the output signal due to the reality condition of the gravitational-wave signals $H(\Omega)=H^{*}(-\Omega)$.
-Furthermore, we also note that the existence of the factor $\mathfrak{R}$}} leads to the signal-referred noise spectral density in which the shot noise is not reduced even if the injected power $I_{0}$ is increased.
+Furthermore, we also note that the existence of the factor $\mathfrak{R}$}} leads to the signal-referred noise spectral density in which the shot-noise contribution is not reduced even if the injected power $I_{0}$ is increased.
 This contradicts the consensus in the community of gravitational-wave experiments.
-Originally, the additional classical carrier field $\mathfrak{R}$}} comes from the solutions (172) and (174) to the Heisenberg equations as the terms proportional to $2\pi\delta(\Omega)$.
+Originally, the additional classical carrier field $\mathfrak{R}$}} comes from the solutions (172) and (LABEL:eq:Eq-of-motion-XYITM-com-tuned-wobifreq-sol-sum-approx) to the Heisenberg equations as the terms proportional to $2\pi\delta(\Omega)$.
 Furthermore, these terms come from the radiation pressure forces (162)–(165) as the classical constant force in the time domain.
@@
 In both classical and quantum mechanics of a forced harmonic oscillator, the constant force in the harmonic oscillator leads to changes in the equilibrium points of the oscillator.
-In actual experiments, the interferometers should be operated at these equilibrium points of mirrors.
+In actual experiments, the interferometers should be operated at the mirrors’ equilibrium positions.
 Therefore, to take into account the effects of the constant forces, we have to change the tuning condition (161) so that the laser is tuned at the equilibrium points of mirrors.
@@
 First, we consider the modification of the tuning condition (161).
-To do this, we note that the cavity propagation conditions (78) and (79) yields that the retarded effects in the electric field operators are determined by
+To do this, we note that the cavity propagation conditions (78) and (79) yield that the retarded effects in the electric field operators are determined by
 $\tau+\frac{1}{c}\hat{X}_{x}(t-\tau)$ $=$ $\frac{L+\hat{X}_{XEM}(t-\tau)-\hat{X}_{XITM}(t-\tau)}{c},$
@@
 $\quad\times\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega).$
+================ at 17:50 on 13th April in 2026
 Keep in our mind the existence of these constant terms ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$, the Heisenberg equations (141)–(144) of mirrors’ motion are given by
@@
 where $\hat{X}_{x}^{(fluc)}:=\hat{X}^{(fluc)}_{XEM}-\hat{X}_{XITM}^{(fluc)}$ and $\hat{X}_{y}^{(fluc)}:=\hat{X}^{(fluc)}_{YEM}-\hat{X}_{YITM}^{(fluc)}$.
-Here, $\tau^{(o)}$ is regarded as the operation point of the cavity length of the Fabri-Pérot interferometer.
+Here, $\tau^{(o)}$ is regarded as the operation point of the cavity length of the Fabry-Pérot interferometer.
 Equations (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom) indicate that we may replace
@@
 As in the cases of Eqs.
-(LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), $\tau_{x}^{(o)^{\prime}}$ and $\tau_{y}^{(o)^{\prime}}$ is regarded as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer.
+(LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), $\tau_{x}^{(o)^{\prime}}$ and $\tau_{y}^{(o)^{\prime}}$ is regarded as the operation points of the length between BS and ITMs in the Fabry-Pérot interferometer.
 Equation (255) and (256) indicate that we may replace
@@
 (234).
-This recover the conventional noise-spectral density except for the fact that we have to use the definitions (187) and (188) of the variables $h_{SQL}$ and $\kappa$ with the replacement $\tau\rightarrow\tau^{(o)}$, respectively.
+This recovers the conventional noise-spectral density except for the fact that we have to use the definitions (187) and (188) of the variables $h_{SQL}$ and $\kappa$ with the replacement $\tau\rightarrow\tau^{(o)}$, respectively.
 The difference between the resulting noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ and Kimble’s noise spectral density $S_{(K)}^{(K)}$ [10] are depicted in Fig.
@@
 If we achieve the above complete equilibrium tuning in experiments, we can realize Kimble’s noise spectral density [10] with the modifications of $\kappa$ and $h_{SQL}$ as an idealized case.
-However, in some experiments, we might not be able to accomplish the above complete equilibrium tuning for some reasons.
+However, in some experiments, we might not be able to achieve complete equilibrium tuning for various reasons.
 In this subsection, we consider the case where we cannot realize the above complete equilibrium tuning and how this incompleteness appears in the noise spectral densities through our consideration within this paper.
@@
 (240), (245), and (246), respectively.
-Here, $\epsilon$ is a phenomenological parameter within $[0,1]$.
-As in the case of the complete equilibrium tuning, we regard $\tau_{x}^{\epsilon(o)^{\prime}}$ and $\tau_{y}^{\epsilon(o)^{\prime}}$ as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer.
+Here, $\epsilon$ is a parameter of incompleteness of the tuning within $[0,1]$.
+We choose $\epsilon$ so that $\epsilon=0$ corresponds to the tuning point, which is the equilibrium point of the pendulum, and $\epsilon=1$ corresponds to the tuning without taking into account the displacement of the equilibrium point of the pendulum due to the classical radiation pressure from the laser.
+As in the case of the complete equilibrium tuning, we regard $\tau_{x}^{\epsilon(o)^{\prime}}$ and $\tau_{y}^{\epsilon(o)^{\prime}}$ as the operation points of the length between BS and ITMs in the Fabry-Pérot interferometer.
 Equations (270) and (272) indicate that we may replace
 $\tau_{x}^{\prime}\rightarrow\tau_{x}^{\epsilon(o)^{\prime}},$ (275)
-$\tau_{y}^{\prime}\rightarrow\tau_{y}^{\epsilon(o)^{\prime}},$ (275)
+$\tau_{y}^{\prime}\rightarrow\tau_{y}^{\epsilon(o)^{\prime}},$
 $\hat{X}_{XITM}(t)\rightarrow\hat{X}_{XITM}^{\epsilon(o)}(t),$ (277)
-$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{\epsilon(o)}(t)$ (277)
+$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{\epsilon(o)}(t)$
 in Eqs. (62), (63), (69), (70), and any equations derived from them.
@@
 $\hat{X}_{x}^{\epsilon(o)}:=\epsilon\left({{\mathscr{D}}}_{EM}-{{\mathscr{D}}}_{ITM}\right)+\hat{X}_{x}^{(fluc)}.$ (285)
-As in the case of the complete equilibrium tuning, we regard $\tau^{\epsilon(o)}$ as the operation point of the cavity length between ITMs and EMs in the Fabri-Pérot interferometer.
-Equations (278) and (LABEL:eq:L-incomplete-operation-point-y) indicate that we may replace
+As in the case of the complete equilibrium tuning, we regard $\tau^{\epsilon(o)}$ as the operation point of the cavity length between ITMs and EMs in the Fabry-Pérot interferometer.
+Equations (LABEL:eq:L-incomplete-operation-point-x) and (LABEL:eq:L-incomplete-operation-point-y) indicate that we may replace
 $\tau\rightarrow\tau^{\epsilon(o)},$ (286)
@@
 Figure 9: The schematic picture of the incomplete equilibrium tuning.
-The phenomenological parameter $\epsilon$ represents the deviations of the tuning point from the equilibrium point of the pendulum for the mirrors.
+The incomplete parameter $\epsilon$ represents the deviations of the tuning point from the equilibrium point of the pendulum for the mirrors.
 $\epsilon=0$ corresponds to the complete equilibrium tuning.
@@
 Even if we make the mirrors stay at the operation point so that $\epsilon\neq 0$, the mirrors feel their accelerations due to the classical radiation pressure and try to move with the frequency $\omega_{p}$ of the pendulum.
-In this case, these accelerations are measured through the feedback control system, and its signal is included in the feedback electric current.
-In this case, the noise in the feedback electric current may deviate from the noise at the photodetection.
+In this case, these accelerations are measured by the feedback control system, and their signals are included in the feedback electric current.
+In this case, the noise in the feedback electric current may differ from that at the photodetection.
 For this reason, we call the tuning conditions $\epsilon\neq 0$ as “incomplete equilibrium tuning.” The schematic picture of the incomplete equilibrium tuning is depicted in Fig.
 9.
-Figure 10: The square-root of the explicit signal referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=0)}$ for each $\epsilon$ is depicted in the range1Hz to 20 kHz with different values of the laser power $I_{0}$.
+Figure 10: The square root of the explicit signal referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=0)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$.
 $S_{(H)}^{(EQ)}$ is given by Eq.
@@
 However, if the incompleteness $\epsilon$ increases, the deviation from Kimble’s noise spectral density can be seen.
-After all, in the high-frequency region, the noise spectral density shows that the shot noise does not decrease due to the increase of the incident laser power $I_{0}$.
+After all, in the high-frequency region, the noise spectral density shows that the shot-noise contribution in the signal-referred noise spectral density does not decrease due to the increase of the incident laser power $I_{0}$.
 This is merely due to the contribution of the additional classical carrier field $\epsilon{{\mathfrak{R}}}$, which may be dominant in the incomplete equilibrium tuning.
@@
 [14].
-In the gravitational-wave community, it is well-known that the standard quantum limit can be violated through appropriate changes of the interferometer configurations.
+In the gravitational-wave community, it is well known that the standard quantum limit can be surpassed using squeezed laser states.
+This indicates that violating the “standard quantum limit” does not imply violating a physical principle.
 Therefore, we may regard the dip in Fig.
-10 as one of them.
-Figure 11: The square-root of the explicit signal-referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=\pi/4)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$.
+10 as not a violation of the physical principle.
+Figure 11: The square root of the explicit signal-referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=\pi/4)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$.
 $S_{(H)}^{(EQ)}$ is given by Eq.
@@
 As a result, the bottom figures in Fig.
-11 show the property that the shot noise is not reduced even if the input power $I_{0}$ is increased.
+11 show the property that the shot-noise contribution is not reduced even if the input power $I_{0}$ is increased.
 Incidentally, the bottom right panel in Fig.
@@
 From the input-output relation (189), if the term $\mathfrak{R}$}} in the classical carrier field, which is proportional to $2\pi\delta(\Omega)$ is negligible compared with $1$ in the same term, we can realize the ideal input-output relation.
-In the case of the incomplete equilibrium tuning, the phenomenological parameter $\epsilon$, which we introduced, appears as $\epsilon{{\mathfrak{R}}}$ in this classical carrier field of the input-output relation as mentioned above.
+In the case of the incomplete equilibrium tuning, the incomplete parameter $\epsilon$, which we introduced, appears as $\epsilon{{\mathfrak{R}}}$ in this classical carrier field of the input-output relation as mentioned above.
 Therefore, if we achieve the $\epsilon$ so that
-$\epsilon{{\mathfrak{R}}}\lesssim O(10^{-1}),$ (299)
-we can realize the idealized input-output relation which as shown in Ref. [8]. From the order of magnitude (218) of $\mathfrak{R}$}}, $\epsilon$ can be estimated as
-$\epsilon$ $\lesssim$ $O(10^{-1}){{\mathfrak{R}}}^{-1}$ (301)
-$\sim$ $O(10^{-1})\frac{mc^{2}\omega_{p}^{2}T^{2}}{64I_{0}\omega_{0}}$
-$\sim$ $3\times 10^{-4}\times\left(\frac{10^{2}W}{I_{0}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right)\left(\frac{m}{40\mbox{kg}}\right)$
+$\epsilon{{\mathfrak{R}}}\lesssim O(10^{-2}),$ (299)
+we can realize the idealized input-output relation which is shown in Ref. [10]. From the order of magnitude (218) of $\mathfrak{R}$}}, $\epsilon$ can be estimated as
+$\epsilon$ $\lesssim$ $O(10^{-2}){{\mathfrak{R}}}^{-1}$ (301)
+$\sim$ $O(10^{-2})\frac{mc^{2}\omega_{p}^{2}T^{2}}{64I_{0}\omega_{0}}$
+$\sim$ $3\times 10^{-5}\times\left(\frac{10^{2}W}{I_{0}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right)\left(\frac{m}{40\mbox{kg}}\right)$
 $\quad\quad\quad\quad\times\left(\frac{\omega_{p}}{2\pi\times 1\mbox{Hz}}\right)^{2}\left(\frac{T}{10^{-2}}\right)^{2}$
 for the equal mass case $m=m_{EM}=m_{ITM}$. In terms of the mirror displacements $\epsilon{{\mathscr{D}}}_{EM}$ and $\epsilon{{\mathscr{D}}}_{ITM}$, which are estimated in Eqs. (264)–(269) as
-$\epsilon{{\mathscr{D}}}_{EM}$ $\lesssim$ $+O(10^{-1})\frac{cT}{16\omega_{0}}$
-$\sim$ $+3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right)$
+$\epsilon{{\mathscr{D}}}_{EM}$ $\lesssim$ $+O(10^{-2})\frac{cT}{16\omega_{0}}$
+$\sim$ $+3\times 10^{-12}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right)$
 and
-$\epsilon{{\mathscr{D}}}_{ITM}$ $\lesssim$ $-O(10^{-1})\frac{cT}{16\omega_{0}}$
-$\sim$ $-3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right).$
+$\left|\epsilon{{\mathscr{D}}}_{ITM}\right|$ $\lesssim$ $O(10^{-2})\frac{cT}{16\omega_{0}}$
+$\sim$ $3\times 10^{-12}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right).$
 Thus, if we can control the mirrors’ positions whose deviations from the equilibrium points of the pendulum and the radiation pressure forces are less than $\epsilon{{\mathscr{D}}}_{EM}$ and $\epsilon{{\mathscr{D}}}_{ITM}$, we can realize the ideal Kimble noise spectral density with the modification of $\kappa$ and $h_{SQL}$.
@@
 IX Summary and Discussion
-The primary purpose of this paper was to develop the theoretical description of the quantum noise in gravitational-wave detectors in detail, which was motivated by the theoretical developments of the measurement theory in quantum field theories.
-In the mathematically rigorous measurement theory in quantum mechanics [12], the specification of the finally observed quantum operator is an essential issue in the theory.
+The primary purpose of this paper was to develop a detailed theoretical description of quantum noise in gravitational-wave detectors, motivated by advances in measurement theory in quantum field theory.
+In the mathematically rigorous measurement theory in quantum mechanics [14], the specification of the final observed quantum operator is an essential issue in the theory.
 We must extend this mathematical theory to quantum field theories to apply it to gravitational-wave detectors.
-Besides this extension of the quantum measurement theory to quantum field theories, the above specification of the finally observed quantum operator is crucially important.
+Besides this extension of quantum measurement theory to quantum field theories, the above specification of the final observed quantum operator is crucial.
 Actually, in quantum measurement theories, there is a famous problem of the von Neumann chain [19], i.e., where we should regard the measurement outcomes as the classical information in the sequence of the quantum measurements as emphasized in Sec.
 I.
-In this paper, we adopt the standing point that the von Neumann chain should be cut at the moment that the photodetectors convert the information in photon signals to the electric currents.
+In this paper, we adopt the view that the von Neumann chain should be cut at the point where photodetectors convert photon signals into electric currents.
 Note that actual gravitational-wave detectors employ a feedback system to control the stability of the detector.
 However, from our standpoint, we regard the feedback current as a classical electric current.
-Therefore, we do not have to care about the quantum properties of the feedback control system.
+Therefore, we do not need to consider the quantum properties of the feedback control system.
 If this premise is wrong, we have to discuss quantum feedback control systems [20] as emphasized in Sec.
 I, which is beyond the current scope of this paper.
-After the introduction of the notions which was used within this paper in Sec.
+After the introduction of the notions that were used within this paper in Sec.
 II, we first developed the description of the DC readout scheme in terms of quantum electrodynamics.
@@
 After the general arguments of the DC readout scheme in Sec.
-III, we carefully discussed the input-output relation of the Fabri-Pérot gravitational-wave detectors, because the details of the input-output relation are necessary to examine the properties of the DC readout scheme.
+III, we carefully discussed the input-output relation of the Fabry-Pérot gravitational-wave detectors, because the details of the input-output relation are necessary to examine the properties of the DC readout scheme.
 Therefore, in Secs.
-IV, V, and VI, we described the detailed theoretical description of the derivation of the input-output relation of the Fabri-Pérot interferometer.
+IV, V, and VI, we described the detailed theoretical description of the derivation of the input-output relation of the Fabry-Pérot interferometer.
 In Sec.
-IV, we derived the input-output relation through the propagation effects and the reflection effects due to the mirrors in the Fabri-Pérot interferometer.
+IV, we derived the input-output relation through the propagation effects and the reflection effects due to the mirrors in the Fabry-Pérot interferometer.
 In the previous works of gravitational-wave detectors, it has been considered only the end mirrors’ motions.
-However, in this paper, we also take into account the motions of the intermediate mirrors.
-We obtained the input-output relation without any specification of the mirror displacements.
+However, in this paper, we also account for the motions of the input test masses.
+We obtained the input-output relation without specifying the mirror displacements.
 The equations of motion of mirrors determine these mirror displacements.
-Since we discussed quantum motions of mirrors, we used the Heisenberg equations of motion in quantum mechanics as the equation of motion for mirrors.
+Since we discussed the quantum motion of mirrors, we used the Heisenberg equations of motion in quantum mechanics as the equations of motion for mirrors.
 Furthermore, we introduce the fundamental frequency $\omega_{p}$ of the mirrors’ pendulum, where $\omega_{p}$ is estimated as $\sim 1$Hz.
@@
 V, we briefly reviewed a quantum mechanical forced harmonic oscillator, which is essential to our arguments on the motions of mirrors in gravitational-wave detectors.
-Based on the understanding of this quantum forced harmonic oscillator, we explicitly derived the equations of motion for the mirrors in the Fabri-Përot interferometer in Sec.
+Based on the understanding of this quantum forced harmonic oscillator, we explicitly derived the equations of motion for the mirrors in the Fabry-Përot interferometer in Sec.
 V.2.
-Then, we pointed out that the evaluation of the radiation pressure effects is essential.
-Although the details are described in Appendix B, we review the outline of the derivation of the radiation pressure forces to the intermediate and end mirrors in Fabri-Pérot gravitational-wave detectors in Sec.
+Then, we noted that evaluating the effects of radiation pressure is essential.
+Although the details are described in Appendix B, we review the outline of the derivation of the radiation pressure forces to the input test masses and to the end test masses in Fabry-Pérot gravitational-wave detectors in Sec.
 V.3.
@@
 IV and the solutions to the Heisenberg equation derived in Sec.
-V.4, we obtain the final input-output relation of the Fabri-Pérot interferometer in Sec.
+V.4, we obtain the final input-output relation of the Fabry-Pérot interferometer in Sec.
 VI.
-Then, we reach the input-output relation for the Fabri-Pérot interferometer (189).
+Then, we reach the input-output relation for the Fabry-Pérot interferometer (189).
 Although the parameters $\beta$, $\kappa$, and $h_{SQL}$ in the input-output relation are different from these parameters in Ref.
@@
 4.
-The frequency dependence of these differences comes from the phase difference between the motion of EMs and the motion of ITMs.
+The frequency dependence of these differences arises from the phase difference between the motion of EMs and that of ITMs.
 Due to these differences, we reach the difference of the $S_{(K)}^{(EQ)}$ and $S_{(K)}^{(K)}$ as depicted in Fig.
@@
 In addition to the above difference between the input-output relation (189), the difference also exists in the classical carrier part as the parameter $\mathfrak{R}$}}.
-This classical carrier field with the modification $\mathfrak{R}$}} is used as the reference field to measure the gravitational-wave signal in the DC readout scheme.
+This classical carrier field, modified by $\mathfrak{R}$}}, is used as the reference field to measure the gravitational-wave signal in the DC readout scheme.
 Therefore, the classical carrier field is essential in the DC readout scheme.
@@
 [10].
-As a result of the additional classical carrier field $\mathfrak{R}$}}, we have to conclude that the shot noise in the high-frequency range does not decrease even if the incident laser power is increased.
-This contradicts to the description in Ref.
-[8] which is the common knowledge in the community of gravitational-wave experiments.
+As a result of the additional classical carrier field $\mathfrak{R}$}}, we have to conclude that the shot noise contribution to the high-frequency range of the signal-referred noise spectral density does not decrease even if the incident laser power is increased.
+This contradicts the description in Ref.
+[10], which is the common knowledge in the community of gravitational-wave experiments.
 In this situation, we reconsidered the tuning points (161) and (177) in Sec.
@@
 In the time domain, these terms, proportional to $2\pi\delta(\Omega)$, represent the constant forces in the radiation pressure forces.
-Here, we note that the equations of motion for mirrors are given in the form of a forced harmonic oscillator.
+Here, we note that the equations of motion for mirrors are expressed as a forced harmonic oscillator.
 In simple classical mechanics, or even in quantum mechanics, a constant force in a harmonic oscillator yields the deviation of the equilibrium point.
@@
 (266) and (269).
-We also considered the incomplete equilibrium tuning in which the elimination of the deviations of the mirror displacements from their equilibrium points is not complete through the introduction of the phenomenological parameter $\epsilon$.
+We also considered incomplete equilibrium tuning, in which deviations in the mirror displacements from their equilibrium points are partially eliminated through the introduction of the incomplete parameter $\epsilon$.
 We call this elimination the “incomplete equilibrium tuning.” If $\epsilon=0$, we can realize the complete equilibrium tuning discussed above.
@@
 (298).
-In both cases, if the incompleteness $\epsilon$ is sufficiently large, there are cases where the shot noise in the high-frequency region does not decrease even if the incident power is increased.
+In both cases, if the incompleteness $\epsilon$ is sufficiently large, there are cases where the shot-noise contribution to the high-frequency region in the signal-referred noise spectral density does not decrease even if the incident power is increased.
 We also evaluate the incompleteness $\epsilon$ when we can realize the behavior near the complete equilibrium tuning in Eqs.
@@
 [12, 13] due to the same reason.
-In other words, there are many rooms to be developed through an accurate understanding within this paper.
+In other words, there is ample room for development through an accurate understanding of the ingredients in this paper.
 We leave these rooms for future work.
@@
 We hope that our consideration within this paper will motivate us to develop the measurement theory in quantum field theories.
-Finally, we have to emphasize that the purpose of this paper is not to point out new techniques nor to develop new ideas for gravitational-wave detectors themselves.
-However, in this paper, we aim to explicitly show the derivation of the common knowledge in the community of the gravitational-wave detections to clarify their concepts from a more theoretically accurate point of view.
-Due to the more precise theoretical arguments on the noise spectral density, we could discuss the incomplete equilibrium tuning of the Frabri-Pérot interferometer as an imperfection of the interferometric gravitational-wave detectors.
+Finally, we emphasize that the purpose of this paper is not to propose new techniques nor to develop new ideas for gravitational-wave detectors themselves.
+However, in this paper, we aim to explicitly show the derivation of the common knowledge in the community of gravitational-wave detections, thereby clarifying their concepts from a more theoretically accurate perspective.
+Due to the more precise theoretical arguments on the noise spectral density, we could discuss the incomplete equilibrium tuning of the Frabry-Pérot interferometer as an imperfection of the interferometric gravitational-wave detectors.
 We hope the ingredients of this paper and this kind of research will be helpful for the further development of detector science in gravitational-wave detectors.
@@
 (33).
-As in the case of Glauber’s photon number, using the formulae
+As in the case of Glauber’s photon number, through the final input-output relation (189) which was derived in Sec.
+VI, we obtain the formulae
 $\langle\hat{b}_{n}^{\dagger}(\omega_{0}-\omega)\hat{b}_{n}^{\dagger}(\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega_{0}+\omega-\omega_{1}).$ (318)
@@
 $\langle\hat{b}_{n}(\omega_{0}+\omega)\hat{b}_{n}(\omega_{1})\rangle$ $\propto$ $2\pi\delta(\omega+\omega_{1}-\omega_{0}),$ (321)
-we reach the expression
+in the same way to obtain Eqs. (39)–(42) in Sec. II.2. Using these formulae, (318)–(321) we reach the expression
 $2\pi\delta(\Omega-\Omega^{\prime})S_{s_{P}}(\Omega)$ (322)
@@
 V.2.
-The crucial premise of this evaluation is that the radiation pressure forces on the mirrors are determined by the power of the laser, which corresponds to the pointing flux of the laser that affects the mirrors.
+The crucial premise of this evaluation is that the radiation pressure forces on the mirrors are determined by the laser’s power, which corresponds to the laser’s pointing flux and affects the mirrors.
 As described in Ref.
@@
 In any case, we have to note that even when we consider the power recycling technique and the signal recycling technique, the expressions of the radiation pressure forces (326), (327), (329), and (331) should be used.
-We also note that we have to evaluate the Fourier transformations of the mirror displacement operators $D_{d}^{\dagger}\hat{Z}_{x}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{y}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{XITM}D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{YITM}D_{d}$ instead of operators $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ when we evaluate the input-output relation (117).
-Here, the operator $D_{d}$ is the displacement operator associated with the coherent state for the optical quadrature $\hat{D}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ which is defined by Eq.
+We also note that we have to evaluate the Fourier transformations of the mirror displacement operators ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}{{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}{{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}{{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}{{\mathfrak{D}}}_{d}$ instead of operators $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ when we evaluate the input-output relation (117).
+Here, the operator ${{\mathfrak{D}}}_{d}$ is the displacement operator associated with the coherent state for the optical quadrature $\hat{D}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ which is defined by Eq.
 (101).
-For the quadrature $\hat{D}(\omega)$, $D_{d}^{\dagger}\hat{D}(\omega)D_{d}$ are given by Eq.
+For the quadrature $\hat{D}(\omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{D}(\omega){{\mathfrak{D}}}_{d}$ are given by Eq.
 (104).
-Trivially, the quadrature $\hat{A}(\omega)$ commutes with the displacement operator $D_{d}$.
-Keep in our mind the properties of the displacement operator $D_{d}$, we evaluate the radiation pressure forces (326), (327), (329), and (331) by the operations $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right.
+Trivially, the quadrature $\hat{A}(\omega)$ commutes with the displacement operator ${{\mathfrak{D}}}_{d}$.
+Keep in our mind the properties of the displacement operator ${{\mathfrak{D}}}_{d}$, we evaluate the radiation pressure forces (326), (327), (329), and (331) by the operations ${{\mathfrak{D}}}_{d}^{\dagger}$ from the left and ${{\mathfrak{D}}}_{d}$ from the right.
 Furthermore, we ignore the terms include $\hat{D}_{v}(\omega)\hat{D}_{v}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{A}(\omega^{\prime})$, and $\hat{A}(\omega)\hat{A}(\omega^{\prime})$.
@@
 Through these evaluations, the radiation pressure forces (326), (327), (329), and (331) are given by
-$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (332)
+$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXEM}(t){{\mathfrak{D}}}_{d}$ (332)
 $=$ $\frac{\hbar T}{2m_{EM}c}e^{+i\omega(\tau+\tau_{x}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
@@
 $+i\frac{\hbar T}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{2}\omega_{1}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 $+i\frac{\hbar T}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
@@
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 […]
-$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (333)
+$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYEM}(t){{\mathfrak{D}}}_{d}$ (333)
 $=$ $\frac{\hbar T}{2m_{EM}c}e^{+i\omega(\tau+\tau_{y}^{\prime})}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
@@
 $+i\frac{\hbar T}{m_{EM}c}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\frac{\omega_{2}}{c}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 $+i\frac{\hbar T}{m_{EM}c^{2}}e^{+i\omega\tau}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\omega_{2}\sqrt{|\omega_{1}\omega_{2}|}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
@@
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 […]
-$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$ (334)
+$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXITM}(t){{\mathfrak{D}}}_{d}$ (334)
 $=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{x}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
@@
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 $-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{x}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 […]
-$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$ (335)
+$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYITM}(t){{\mathfrak{D}}}_{d}$ (335)
 $=$ $-\sqrt{1-T}\frac{\hbar}{2m_{ITM}c}e^{+i\omega\tau_{y}^{\prime}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\sqrt{|\omega_{1}(\omega-\omega_{1})|}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}$
@@
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[e^{+2i(\omega-\omega_{1})\tau}+e^{+2i\omega_{1}\tau}-\sqrt{1-T}(1+e^{+2i\omega\tau})\right]$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 $-iT\sqrt{1-T}\frac{2\hbar}{m_{ITM}c^{2}}\int_{-\infty}^{+\infty}\frac{d\omega_{1}}{2\pi}\int_{-\infty}^{+\infty}\frac{d\omega_{2}}{2\pi}\sqrt{|\omega_{1}\omega_{2}|}\omega_{2}e^{+i(\omega_{2}+\omega-\omega_{1})\tau}e^{+i(\omega_{1}+\omega_{2})\tau_{y}^{\prime}}\left[1-\sqrt{1-T}e^{+2i\omega_{1}\tau}\right]^{-1}$
 $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{2}\tau}\right]^{-1}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
 […]
@@
 Then, we obtain
-$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (336)
+$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXEM}(t){{\mathfrak{D}}}_{d}$ (336)
 $=$ $+\frac{N^{2}T\hbar\omega_{0}}{m_{EM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}2\pi\delta(\Omega)$
@@
 $+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+i\Omega\tau}\left[1-e^{+2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
 $-4N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+2i\Omega\tau}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
-$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (337)
+$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYEM}(t){{\mathfrak{D}}}_{d}$ (337)
 $=$ $+\frac{N^{2}T\hbar\omega_{0}}{m_{EM}c}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}2\pi\delta(\Omega)$
@@
 $+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}e^{+i\Omega\tau}\left[1-e^{+2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
 $-4N^{2}T\sqrt{1-T}e^{+2i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{EM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (338)
-$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$ (338)
+$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXITM}(t){{\mathfrak{D}}}_{d}$
 $=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega)$
@@
 $-N^{2}T\sqrt{1-T}2\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
 $+4N^{2}T\sqrt{1-T}e^{+i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (339)
-$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$ (339)
+$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYITM}(t){{\mathfrak{D}}}_{d}$
 $=$ $-2N^{2}\sqrt{1-T}\frac{\hbar\omega_{0}}{m_{ITM}c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]2\pi\delta(\Omega)$
@@
 $-2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
 $+4N^{2}T\sqrt{1-T}e^{+i\Omega\tau}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{m_{ITM}c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
 […]
@@
 Furthermore, from the definition (LABEL:eq:radiation-pressure-Fourier-X-def) and (LABEL:eq:radiation-pressure-Fourier-Y-def) of ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$, we obtain
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (340)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$ (340)
 $=$ $\frac{N^{2}\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
@@
 $+\sin(2\omega_{0}\tau)\frac{2N^{2}T\sqrt{1-T}\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
 $-\sin(2\omega_{0}\tau)e^{+i\Omega\tau}\frac{4N^{2}T\sqrt{1-T}\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"},$
-$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ (341)
+${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ (341)
 $=$ $N^{2}\frac{\hbar\omega_{0}}{c}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[\frac{T}{m_{EM}}+\frac{2\sqrt{1-T}}{m_{ITM}}\left[\cos(2\omega_{0}\tau)-\sqrt{1-T}\right]\right]2\pi\delta(\Omega)$
@@
 $+2N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)\frac{\hbar\omega_{0}^{2}}{c^{2}}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
 $-4N^{2}T\sqrt{1-T}\sin(2\omega_{0}\tau)e^{+i\Omega\tau}\frac{\hbar\omega_{0}^{2}}{c^{2}}\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]\left[1-\sqrt{1-T}e^{+2i\omega_{0}\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{-2i\omega_{0}\tau}\right]^{-1}$
-$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
+$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
 […]
 $+\mbox{``rapid oscillation terms with the frequency $2\omega_{0}\pm\omega$"}.$
-In these expressions (338), (339), (340), and (341), we have the terms proportional to $D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z_{y}}(\Omega)D_{d}$ which includes the factor $\sin(2\omega_{0}\tau)$.
-These terms correspond to the optical spring effects of the Fabri-Pérot interferometer [32, 33].
+In these expressions (338), (339), (340), and (341), we have the terms proportional to ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z_{y}}(\Omega){{\mathfrak{D}}}_{d}$ which includes the factor $\sin(2\omega_{0}\tau)$.
+These terms correspond to the optical spring effects of the Fabry-Pérot interferometer [37, 38].
 However, these optical spring effects can be ignored through the tuning condition (161) due to the factor $\sin(2\omega_{0}\tau)$.